Jefferson Lab HALL A Logbook for March 11 - Full Listing

  • Main INDEX

    Thursday

  • 353971 03/31/11 23:58 israely Neutron screenshots
  • 353970 03/31/11 23:57 israely bigbite screenshots
  • 353969 03/31/11 23:56 israely Target screenshot
  • 353968 03/31/11 23:53 Sarah Swing shift run list (analysis)
  • 353967 03/31/11 23:50 Sarah Screen capture of neutron detector single rates during run 3197
  • 353966 03/31/11 23:50 Mitra S Target screens
  • 353965 03/31/11 23:49 Sarah Screen capture of scaler data screen and bigbite scintillator single rates during run 3197
  • 353964 03/31/11 23:45 a-onl Start_Run_3197,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production 4He 20 cm Target, 4 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353963 03/31/11 23:43 a-onl End_of_Run_3196
  • 353962 03/31/11 23:08 Ramesh spot check for shift check-list
  • 353961 03/31/11 22:53 eip RHRS momentum
  • 353960 03/31/11 22:50 Ramesh General Tools and BigBite magnet Control
  • 353959 03/31/11 22:38 Sarah Screen capture of neutron detector single rates during run 3196
  • 353958 03/31/11 22:37 Sarah Screen capture of scaler data screen and bigbite scintillator single rates during run 3196
  • 353957 03/31/11 22:31 Sarah CODA crash
  • 353956 03/31/11 22:29 a-onl Start_Run_3196,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production 4He 20 cm Target, 4 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353955 03/31/11 22:23 a-onl Start_Run_3195,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production 4He 20 cm Target, 4 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353954 03/31/11 22:22 a-onl End_of_Run_3194
  • 353953 03/31/11 21:13 Sarah Screen capture for neutron detector single rates during run 3194
  • 353952 03/31/11 21:11 Sarah Screen capture of scaler data screen and the bigbite scintillator single rates during run 3194
  • 353951 03/31/11 21:06 Mitra S Target screens
  • 353950 03/31/11 21:03 a-onl Start_Run_3194,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production 4He 20 cm Target, 4 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353949 03/31/11 21:02 a-onl End_of_Run_3193
  • 353948 03/31/11 20:35 auto_snap Radiation and Ozone snapshot
  • 353947 03/31/11 20:35 auto_snap Miscellaneous snapshot
  • 353946 03/31/11 20:35 auto_snap Hall A Tools snapshot
  • 353945 03/31/11 20:35 auto_snap Compton Polarimeter snapshot
  • 353944 03/31/11 20:35 auto_snap BigBox Magnet snapshot
  • 353943 03/31/11 20:35 auto_snap Beam Position and Current snapshot
  • 353942 03/31/11 19:56 Sarah Screen capture for neutron detector single rates during run 3193
  • 353941 03/31/11 19:54 Sarah Screen capture of the scaler data screen and the bigbite scintillator single rates during run 3193
  • 353940 03/31/11 19:48 a-onl Start_Run_3193,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production 4He 20 cm Target, 4 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353939 03/31/11 19:47 a-onl End_of_Run_3192
  • 353938 03/31/11 19:16 Ramesh Small history of Left Dipole NMR unlocking
  • 353937 03/31/11 19:09 Sarah Screen capture of the neutron detector single rates screen during run 3192
  • 353936 03/31/11 19:07 Sarah Screen capture for scaler screen and bigbite scintillator single rates for run 3192
  • 353935 03/31/11 18:35 Mitra S Target screens
  • 353934 03/31/11 18:31 a-onl Start_Run_3192,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production 4He 20 cm Target, 4 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353933 03/31/11 18:30 a-onl End_of_Run_3191
  • 353932 03/31/11 18:12 Ramesh Occasional Left Dipole alarms
  • 353931 03/31/11 17:53 V. Sulkosky/ Pmiss Comparison
  • 353930 03/31/11 17:26 Sarah Screen capture of the neutron detector single rates during run 3191
  • 353929 03/31/11 17:25 Sarah Screen capture of the scaler data screen and the bigbite scintillator single rates screen during run 3191
  • 353928 03/31/11 17:13 a-onl Start_Run_3191,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production 4He 20 cm Target, 4 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353927 03/31/11 17:11 a-onl End_of_Run_3190
  • 353926 03/31/11 17:07 yez LHRS:PRL2#15 still overflow. Will do more change during NO-production run
  • 353925 03/31/11 16:41 Sarah Screen capture of neutron detector single rates screen during run 3190
  • 353924 03/31/11 16:40 Sarah Screen capture of scaler data and bigbite scintillator single rates screens during run 3190
  • 353923 03/31/11 16:34 V. Sulkosky Deadtime Statistics for Run 3055 (8 uA)
  • 353922 03/31/11 16:30 V. Sulkosky Deadtime Statistics for Run 3068 (6 uA)
  • 353921 03/31/11 16:24 V. Sulkosky Deadtime Statistics for Run 3056 (2 uA)
  • 353920 03/31/11 16:22 Mitra S Target screens
  • 353919 03/31/11 16:20 V. Sulkosky Deadtime Statistics for Run 3044 (4 uA)
  • 353918 03/31/11 16:11 V. Sulkosky Deadtime from Current Scan
  • 353917 03/31/11 16:02 adaq Beam Time Accounting
  • 353916 03/31/11 16:01 gilman shift summary day 3/31
  • 353915 03/31/11 15:56 a-onl Start_Run_3190,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353914 03/31/11 15:55 a-onl End_of_Run_3189
  • 353913 03/31/11 15:24 gilman spot++ screens for shift checklist
  • 353912 03/31/11 15:18 gilman hall a tools for shift checklist
  • 353911 03/31/11 15:17 yez L-HRS: PRL2#15 HV reduced from -940V to -900V
  • 353910 03/31/11 15:17 a-onl Start_Run_3189,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353909 03/31/11 15:15 a-onl End_of_Run_3188
  • 353908 03/31/11 15:14 gilman setting up HRS right magnets
  • 353907 03/31/11 15:03 yez L-HRS: PRL2 #15 ADC is still overflow. Need to reduce more HV
  • 353906 03/31/11 14:45 a-onl Start_Run_3188,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353905 03/31/11 14:42 a-onl End_of_Run_3187
  • 353904 03/31/11 14:37 gilman raster for 1 uA on He
  • 353903 03/31/11 14:30 a-onl Start_Run_3187,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353902 03/31/11 14:18 a-onl End_of_Run_3186
  • 353901 03/31/11 14:17 gilman spot check with BeO
  • 353900 03/31/11 14:15 a-onl Start_Run_3186,Run_type=Spot++,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=6,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=spot++,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353899 03/31/11 14:13 a-onl End_of_Run_3185
  • 353898 03/31/11 14:12 gilman Q2,3 cycled on both HRS
  • 353897 03/31/11 14:12 gilman moved to BeO
  • 353896 03/31/11 13:50 Fansler Cryo
  • 353895 03/31/11 13:24 Fansler CRYO
  • 353894 03/31/11 13:20 dalton Bigbite magnet off
  • 353893 03/31/11 12:37 gilman Q1 right on cold return
  • 353892 03/31/11 12:34 auto_snap Radiation and Ozone snapshot
  • 353891 03/31/11 12:34 auto_snap Miscellaneous snapshot
  • 353890 03/31/11 12:34 auto_snap Hall A Tools snapshot
  • 353889 03/31/11 12:34 auto_snap Compton Polarimeter snapshot
  • 353888 03/31/11 12:34 auto_snap BigBox Magnet snapshot
  • 353887 03/31/11 12:34 auto_snap Beam Position and Current snapshot
  • 353886 03/31/11 12:00 Higinbotham Error Bars
  • 353885 03/31/11 11:19 yez ROC3 still uses ADC gates from back pannel
  • 353884 03/31/11 10:57 igorko carbon to deuterium ratio
  • 353883 03/31/11 10:45 yez L-HRS: Reduce HV of PRL2 #15counter
  • 353882 03/31/11 10:17 V. Sulkosky CODA Downloaded after 13-bit Mode Change
  • 353881 03/31/11 10:08 V. Sulkosky Cerenkov ADC Changed to 13bit Mode
  • 353880 03/31/11 10:02 a-onl Start_Run_3185,Run_type=Cosmics,target_type=20cm dummy,beam_e0=4.46,beam_raster=on,beam_current=6,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353879 03/31/11 09:57
  • 353878 03/31/11 09:49 igorko double in new setting
  • 353877 03/31/11 08:43 igorko X_bj
  • 353876 03/31/11 08:39 gilman Hall being surveyed now
  • 353875 03/31/11 08:14 gilman turning HRS_left magnets off
  • 353874 03/31/11 08:12 a-onl End_of_Run_3184
  • 353873 03/31/11 08:04 a-onl Start_Run_3184,Run_type=Production SRC,target_type=20cm dummy,beam_e0=4.46,beam_raster=on,beam_current=6,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production on 15cm Dummy target,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353872 03/31/11 08:03 a-onl End_of_Run_3183
  • 353871 03/31/11 08:03 adaq Beam Time Accounting
  • 353870 03/31/11 08:00 arie shift summary
  • 353869 03/31/11 07:51 israely Bigbite screenshot
  • 353868 03/31/11 07:51 israely Neutron screenshot
  • 353867 03/31/11 07:50 israely Target Screenshot
  • 353866 03/31/11 07:22 a-onl Start_Run_3183,Run_type=Production SRC,target_type=20cm dummy,beam_e0=4.46,beam_raster=on,beam_current=6,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production on 15cm Dummy target,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353865 03/31/11 07:16 a-onl End_of_Run_3182
  • 353864 03/31/11 07:04 Fansler RQ1
  • 353863 03/31/11 06:37 a-onl Start_Run_3182,Run_type=Production SRC,target_type=20cm dummy,beam_e0=4.46,beam_raster=on,beam_current=6,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production on 15cm Dummy target,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353862 03/31/11 06:27 a-onl End_of_Run_3181
  • 353861 03/31/11 05:56 israely Neutron screenshot
  • 353860 03/31/11 05:44 a-onl Start_Run_3181,Run_type=Production SRC,target_type=20cm dummy,beam_e0=4.46,beam_raster=on,beam_current=6,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production on 15cm Dummy target,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353859 03/31/11 05:42 a-onl End_of_Run_3180
  • 353858 03/31/11 05:20 Higinbotham Dummy Target Restriction Wording Wrong In MCC List
  • 353857 03/31/11 05:17 israely BBS12001 S13 L10.2 is activating the alarm.
  • 353856 03/31/11 05:04 a-onl Start_Run_3180,Run_type=Production SRC,target_type=20cm dummy,beam_e0=4.46,beam_raster=on,beam_current=7,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production on 15cm Dummy target,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353855 03/31/11 05:02 a-onl End_of_Run_3179
  • 353854 03/31/11 05:01 israely Neutron screenshot
  • 353853 03/31/11 05:00 israely bigbite screenshots
  • 353852 03/31/11 04:55 a-onl Start_Run_3179,Run_type=Production SRC,target_type=20cm dummy,beam_e0=4.46,beam_raster=on,beam_current=10,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production on 15cm Dummy target,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353851 03/31/11 04:54 israely MCC didn't allow 4cm dummy 10uA so we moved to 15cm dummy
  • 353850 03/31/11 04:46 israely Moved to 10 cm dummy target
  • 353849 03/31/11 04:38 a-onl End_of_Run_3178
  • 353848 03/31/11 04:34 auto_snap Radiation and Ozone snapshot
  • 353847 03/31/11 04:34 auto_snap Miscellaneous snapshot
  • 353846 03/31/11 04:34 auto_snap Hall A Tools snapshot
  • 353845 03/31/11 04:34 auto_snap Compton Polarimeter snapshot
  • 353844 03/31/11 04:34 auto_snap BigBox Magnet snapshot
  • 353843 03/31/11 04:34 auto_snap Beam Position and Current snapshot
  • 353842 03/31/11 04:31 arie tools
  • 353841 03/31/11 04:26 arie raster_L
  • 353840 03/31/11 04:19 israely Target screenshot
  • 353839 03/31/11 03:18 a-onl Start_Run_3178,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1.5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production on 10cm LD target, 1.5 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353838 03/31/11 03:17 a-onl End_of_Run_3177
  • 353837 03/31/11 02:18 a-onl Start_Run_3177,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1.5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production on 10cm LD target, 1.5 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353836 03/31/11 02:18 a-onl End_of_Run_3176
  • 353835 03/31/11 01:38 a-onl Start_Run_3176,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1.5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production on 10cm LD target, 1.5 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353834 03/31/11 01:37 a-onl End_of_Run_3175
  • 353833 03/31/11 00:38 a-onl Start_Run_3175,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1.5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production on 10cm LD target, 1.5 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353832 03/31/11 00:36 a-onl End_of_Run_3174
  • 353831 03/31/11 00:24 israely N6 TDC
  • 353830 03/31/11 00:10 adaq Beam Time Accounting
  • 353829 03/31/11 00:10 gilman swing shift summary
  • 353828 03/31/11 00:02 israely Target screenshot
  • 353827 03/31/11 00:01 israely neutron screenshot
  • Wednesday

  • 353826 03/30/11 23:59 israely Coda screenshot
  • 353825 03/30/11 23:58 israely bigbite scintillator screenshot
  • 353824 03/30/11 23:56 Sarah Swing shift run list (analysis)
  • 353823 03/30/11 23:54 Mitra S Target screens
  • 353822 03/30/11 23:34 a-onl Start_Run_3174,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1.5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production on 10cm LD target, 1.5 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353821 03/30/11 23:33 a-onl End_of_Run_3173
  • 353820 03/30/11 23:04 lselvy No problems seem to exist
  • 353819 03/30/11 22:31 a-onl Start_Run_3173,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1.5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production on 10cm LD target, 1.5 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353818 03/30/11 22:30 a-onl End_of_Run_3172
  • 353817 03/30/11 21:55 Sarah Screen capture of the neutron detector single rates for run 3172
  • 353816 03/30/11 21:55 Sarah Screen capture of scaler data screen and bigbite scintillator single rates (run 3172)
  • 353815 03/30/11 21:50 Sarah Strange histograms in Run 3172
  • 353814 03/30/11 21:30 gilman spot++ for shift checklist
  • 353813 03/30/11 21:21 a-onl Start_Run_3172,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1.5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production on 10cm LD target, 1.5 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353812 03/30/11 21:19 a-onl End_of_Run_3171
  • 353811 03/30/11 21:16 a-onl Start_Run_3171,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production on 10cm LD target, 3 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353810 03/30/11 21:15 gilman Hall A Tools for shift checklist
  • 353809 03/30/11 21:14 Mitra S Moved target to 10cm LD2
  • 353808 03/30/11 21:05 a-onl End_of_Run_3170
  • 353807 03/30/11 20:53 Sarah Screen capture of the neutron detector single rates during run 3170
  • 353806 03/30/11 20:53 Mitras Target screens
  • 353805 03/30/11 20:53 Sarah Screen capture of the scaler data screen and the bigbite scintillator single rates during run 3170
  • 353804 03/30/11 20:34 auto_snap Radiation and Ozone snapshot
  • 353803 03/30/11 20:34 auto_snap Miscellaneous snapshot
  • 353802 03/30/11 20:34 auto_snap Hall A Tools snapshot
  • 353801 03/30/11 20:34 auto_snap Compton Polarimeter snapshot
  • 353800 03/30/11 20:34 auto_snap BigBox Magnet snapshot
  • 353799 03/30/11 20:33 auto_snap Beam Position and Current snapshot
  • 353798 03/30/11 20:10 igorko HV GUI problem
  • 353797 03/30/11 20:05 Sarah The HV is now 2025 for bar 4, right side, layer 6
  • 353796 03/30/11 20:02 a-onl Start_Run_3170,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=40,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production on slanted carbon, 40 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353795 03/30/11 19:51 gilman 40 uA
  • 353794 03/30/11 19:44 a-onl End_of_Run_3169
  • 353793 03/30/11 19:44 Sarah Nope, it was at 1900 V
  • 353792 03/30/11 19:00 Sarah Screen capture of the neutron detector single rates (run 3169)
  • 353791 03/30/11 18:59 Sarah Screen capture of the scaler data screen and the bigbite scintillator single rates (run 3169)
  • 353790 03/30/11 18:53 Mitras Target screens
  • 353789 03/30/11 18:48 gilman called MCC and asked for 20 uA
  • 353788 03/30/11 18:45 Sarah N6 plane bar 4 right HV has been set back to original value
  • 353787 03/30/11 18:42 a-onl Start_Run_3169,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=20,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production on slanted carbon, 20 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353786 03/30/11 18:40 a-onl End_of_Run_3168
  • 353785 03/30/11 17:43 a-onl Start_Run_3168,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production on slanted carbon, 30 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353784 03/30/11 17:42 a-onl End_of_Run_3167
  • 353783 03/30/11 17:09 Sarah Screen capture of neutron detector single rates (run 3167)
  • 353782 03/30/11 17:08 Sarah Screen capture of scaler data and bigbite scintillator single rates (Run 3167).
  • 353781 03/30/11 16:53 Sarah N6 plane bar 4 right HV has been reduced
  • 353780 03/30/11 16:42 gilman disabled right Q1 cryo alarms
  • 353779 03/30/11 16:27 a-onl Start_Run_3167,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production on slanted carbon, 30 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353778 03/30/11 16:25 Mitra Shabes Target moved to slanted carbon
  • 353777 03/30/11 16:23 mitra shabes Target moved to 5mm carbon
  • 353776 03/30/11 16:21 a-onl End_of_Run_3166
  • 353775 03/30/11 16:18 a-onl Start_Run_3166,Run_type=Check out,target_type=Thick Carbon,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Ramping current to check rates,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353774 03/30/11 16:12 a-onl End_of_Run_3165
  • 353773 03/30/11 16:04 dalton Runplan tonight
  • 353772 03/30/11 16:03 dalton Runplan tonight
  • 353771 03/30/11 16:03 adaq Beam Time Accounting
  • 353770 03/30/11 16:01 Ole Hansen Day Shift Summary
  • 353769 03/30/11 16:01 Jack (Techs) Some optocouplers replaced on LQ2 and LQ3 power supplies
  • 353768 03/30/11 15:57 a-onl Start_Run_3165,Run_type=Check out,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Check out the detectors,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353767 03/30/11 15:54 israely Spot_L
  • 353766 03/30/11 15:51 a-onl End_of_Run_3164
  • 353765 03/30/11 15:43 Fansler RQ1
  • 353764 03/30/11 15:41 a-onl Start_Run_3164,Run_type=Spot++,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=0,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Spot++,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353763 03/30/11 15:36 Ole Hansen Ready for beam
  • 353762 03/30/11 15:34 a-onl End_of_Run_3163
  • 353761 03/30/11 15:32 yez target moved to Slanted Carbon
  • 353760 03/30/11 15:23 Ole Hansen Left NMR
  • 353759 03/30/11 15:18 yez target screen
  • 353758 03/30/11 15:13 orchen BB and HAND angles
  • 353757 03/30/11 15:04 israely his is correction to the angle of bigbit
  • 353756 03/30/11 15:03 Ole Hansen BB & HAND ready
  • 353755 03/30/11 14:59 a-onl Start_Run_3163,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.201,rightarm_theta=24.5,rightarm_sieve=out,comment_text=cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353754 03/30/11 14:54 a-onl End_of_Run_3162
  • 353753 03/30/11 14:53 a-onl End_of_Run_3161
  • 353752 03/30/11 14:51 folts Neutron detector move
  • 353751 03/30/11 14:50 a-onl Start_Run_3161,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=24.5,rightarm_sieve=out,comment_text=cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353750 03/30/11 14:49 a-onl End_of_Run_3160
  • 353749 03/30/11 13:57 a-onl Start_Run_3160,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=24.5,rightarm_sieve=out,comment_text=cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353748 03/30/11 12:35 a-onl End_of_Run_3159
  • 353747 03/30/11 12:33 auto_snap Radiation and Ozone snapshot
  • 353746 03/30/11 12:33 auto_snap Miscellaneous snapshot
  • 353745 03/30/11 12:33 auto_snap Hall A Tools snapshot
  • 353744 03/30/11 12:33 auto_snap Compton Polarimeter snapshot
  • 353743 03/30/11 12:33 auto_snap BigBox Magnet snapshot
  • 353742 03/30/11 12:33 auto_snap Beam Position and Current snapshot
  • 353741 03/30/11 12:25 a-onl Start_Run_3159,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=cosmics for BB,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353740 03/30/11 12:21 V. Sulkosky BigBite Angle now 92 degrees
  • 353739 03/30/11 12:19 yez Target Screen Snapshot
  • 353738 03/30/11 12:18 Ole Hansen BigBite to 93 degrees
  • 353737 03/30/11 12:03 igorko hit wires for BB
  • 353736 03/30/11 12:01 a-onl End_of_Run_3158
  • 353735 03/30/11 11:57 a-onl Start_Run_3158,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=cosmics for BB,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353734 03/30/11 11:56 a-onl End_of_Run_3157
  • 353733 03/30/11 11:53 a-onl Start_Run_3157,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=cosmics for BB,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353732 03/30/11 11:52 Fansler VESDA
  • 353731 03/30/11 11:41 V. Sulkosky Interesting Messages from EB2
  • 353730 03/30/11 11:05 nmuangma fixed missing wire sections
  • 353729 03/30/11 11:00 igorko HAND HV
  • 353728 03/30/11 10:52 a-onl End_of_Run_3155
  • 353727 03/30/11 10:51 a-onl Start_Run_3155,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=rates check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353726 03/30/11 10:50 Ole Hansen Work on quads
  • 353725 03/30/11 10:35 a-onl End_of_Run_3153
  • 353724 03/30/11 10:32 a-onl Start_Run_3153,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=BB missing wire checking,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353723 03/30/11 10:30 V. Sulkosky Right Dipole Settle Time
  • 353722 03/30/11 10:21 Ole Hansen RHRS dipole ramp-down
  • 353721 03/30/11 10:15 a-onl End_of_Run_3151
  • 353720 03/30/11 10:14 Ole Hansen RHRS ramping down??
  • 353719 03/30/11 10:01 Ole Hansen Magnet status
  • 353718 03/30/11 09:58 Ole Hansen RHRS in position
  • 353717 03/30/11 09:34 Ole Hansen Access requested
  • 353716 03/30/11 09:07 Ole Hansen Quad trips
  • 353715 03/30/11 09:05 Ole Hansen RHRS motion problems
  • 353714 03/30/11 09:05 yez Target Heaters' power go higher but still look OK
  • 353713 03/30/11 08:46 Ole Hansen Kinematics change
  • 353712 03/30/11 08:45 igorko missing wires
  • 353711 03/30/11 08:42 a-onl End_of_Run_3151
  • 353710 03/30/11 08:38 a-onl Start_Run_3151,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353709 03/30/11 08:37 a-onl End_of_Run_3150
  • 353708 03/30/11 08:12 israely The golder root is not the same in the bigbite
  • 353707 03/30/11 08:04 adaq Beam Time Accounting
  • 353706 03/30/11 08:04 israely Target screenshot
  • 353705 03/30/11 08:03 Gilman owl shift summary March 30
  • 353704 03/30/11 08:03 israely Neutron Detector screenshot
  • 353703 03/30/11 08:02 israely Bigbite scintillator
  • 353702 03/30/11 07:28 a-onl Start_Run_3150,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353701 03/30/11 07:27 a-onl End_of_Run_3149
  • 353700 03/30/11 06:18 a-onl Start_Run_3149,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353699 03/30/11 06:08 gilman mcc running qe measurment
  • 353698 03/30/11 06:07 a-onl End_of_Run_3148
  • 353697 03/30/11 06:05 gilman HV trip on bb
  • 353696 03/30/11 06:05 a-onl Start_Run_3148,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353695 03/30/11 06:04 a-onl End_of_Run_3147
  • 353694 03/30/11 05:54 a-onl Start_Run_3147,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353693 03/30/11 05:53 a-onl End_of_Run_3146
  • 353692 03/30/11 05:28 arie neutron singles rates
  • 353691 03/30/11 05:27 arie scalers bigbite singles
  • 353690 03/30/11 04:59 gilman spot++ for shift checklist
  • 353689 03/30/11 04:56 gilman Hall A Tolls for shift checklist
  • 353688 03/30/11 04:43 a-onl Start_Run_3146,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353687 03/30/11 04:42 a-onl End_of_Run_3145
  • 353686 03/30/11 04:33 auto_snap Radiation and Ozone snapshot
  • 353685 03/30/11 04:33 auto_snap Miscellaneous snapshot
  • 353684 03/30/11 04:33 auto_snap Hall A Tools snapshot
  • 353683 03/30/11 04:33 auto_snap Compton Polarimeter snapshot
  • 353682 03/30/11 04:33 auto_snap BigBox Magnet snapshot
  • 353681 03/30/11 04:33 auto_snap Beam Position and Current snapshot
  • 353680 03/30/11 03:33 a-onl Start_Run_3145,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353679 03/30/11 03:32 a-onl End_of_Run_3144
  • 353678 03/30/11 02:34 arie neutron rates
  • 353677 03/30/11 02:31 arie scalers bigbite
  • 353676 03/30/11 02:24 eip hv off for hand/bb
  • 353675 03/30/11 02:19 a-onl Start_Run_3144,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353674 03/30/11 02:18 a-onl End_of_Run_3143
  • 353673 03/30/11 01:08 a-onl Start_Run_3143,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353672 03/30/11 01:06 a-onl End_of_Run_3142
  • 353671 03/30/11 00:12 wenbiao target
  • 353670 03/30/11 00:01 adaq Beam Time Accounting
  • 353669 03/30/11 00:00 Sarah Swing shift run list (analysis)
  • Tuesday

  • 353668 03/29/11 23:57 a-onl Start_Run_3142,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353667 03/29/11 23:56 a-onl End_of_Run_3141
  • 353666 03/29/11 23:51 Carlos Swing shift summary
  • 353665 03/29/11 23:36 Sarah Screen capture of neutron detector single rates (run 3141)
  • 353664 03/29/11 23:35 Sarah Screen capture of scalers (Run 3141)
  • 353663 03/29/11 23:27 V. Sulkosky Updated DBB database file
  • 353662 03/29/11 23:00 zwzhao target screen
  • 353661 03/29/11 22:58 Sarah Noisy histogram for N5 bar 3
  • 353660 03/29/11 22:40 Sarah HV trip
  • 353659 03/29/11 22:32 a-onl Start_Run_3141,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353658 03/29/11 22:21 Carlos Beam permit
  • 353657 03/29/11 22:17 a-onl End_of_Run_3140
  • 353656 03/29/11 22:14 a-onl Start_Run_3140,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=JUNK,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353655 03/29/11 22:13 a-onl End_of_Run_3139
  • 353654 03/29/11 22:11 a-onl Start_Run_3139,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=JUNK,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353653 03/29/11 22:11 a-onl End_of_Run_3138
  • 353652 03/29/11 22:02 a-onl Start_Run_3138,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=JUNK,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353651 03/29/11 21:47 Carlos Escorted access requested to look into BB
  • 353650 03/29/11 21:46 a-onl End_of_Run_3137
  • 353649 03/29/11 21:45 Sarah Weird histogram in replay of run 3137
  • 353648 03/29/11 21:31 orchen And this time with the lots attuched
  • 353647 03/29/11 21:22 Sarah Screen Capture of Neutron Detector Single Rates (Run 3137)
  • 353646 03/29/11 21:21 orchen & Igo HAND veto reduction with improved veto timing
  • 353645 03/29/11 21:21 Sarah Screen Capture of Scalers (Run 3137)
  • 353644 03/29/11 21:16 a-onl Start_Run_3137,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353643 03/29/11 21:09 a-onl End_of_Run_3136
  • 353642 03/29/11 20:36 adaq hall a tools
  • 353641 03/29/11 20:35 Carlos spot+
  • 353640 03/29/11 20:33 auto_snap Radiation and Ozone snapshot
  • 353639 03/29/11 20:33 auto_snap Miscellaneous snapshot
  • 353638 03/29/11 20:33 auto_snap Hall A Tools snapshot
  • 353637 03/29/11 20:33 auto_snap Compton Polarimeter snapshot
  • 353636 03/29/11 20:33 auto_snap BigBox Magnet snapshot
  • 353635 03/29/11 20:32 auto_snap Beam Position and Current snapshot
  • 353634 03/29/11 20:29 Sarah Screen Capture of Neutron Detector Singles Rates (run 3136)
  • 353633 03/29/11 20:26 Sarah Screen Capture of Scalers (run 3136)
  • 353632 03/29/11 20:01 a-onl Start_Run_3136,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353631 03/29/11 20:00 a-onl End_of_Run_3135
  • 353630 03/29/11 19:28 V. Sulkosky Re: Target windows for LHRS and RHRS (T5)
  • 353629 03/29/11 18:57 V. Sulkosky Target windows for LHRS and RHRS
  • 353628 03/29/11 18:55 Sarah Screen capture of neutron detector single rates
  • 353627 03/29/11 18:54 Sarah Screen capture of scalers
  • 353626 03/29/11 18:42 Sarah HV trip
  • 353625 03/29/11 18:34 a-onl Start_Run_3135,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353624 03/29/11 18:33 a-onl End_of_Run_3134
  • 353623 03/29/11 18:32 Sarah Run 3132 is Junk
  • 353622 03/29/11 18:09 a-onl Start_Run_3134,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Cosmics, no beam,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353621 03/29/11 18:07 a-onl End_of_Run_3133
  • 353620 03/29/11 18:06 Sarah N5 Plane Right Bar 3 HV Tripped Off
  • 353619 03/29/11 17:44 Sarah Ended run 3132 and restarted CODA
  • 353618 03/29/11 17:42 a-onl Start_Run_3133,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353617 03/29/11 17:37 a-onl End_of_Run_3132
  • 353616 03/29/11 17:34 V. Sulkosky bbps1:2008 HV OFF
  • 353615 03/29/11 17:32 Sarah HV trip; ended run 3131 early
  • 353614 03/29/11 17:31 a-onl Start_Run_3132,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353613 03/29/11 17:28 a-onl End_of_Run_3131
  • 353612 03/29/11 17:16 a-onl Start_Run_3131,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353611 03/29/11 17:15 a-onl End_of_Run_3130
  • 353610 03/29/11 16:47 lselvy HAND and BB HV alarm
  • 353609 03/29/11 16:04 gilman shift summary March 29 day
  • 353608 03/29/11 16:04 adaq Beam Time Accounting
  • 353607 03/29/11 15:51 X. Zheng Target snap shot
  • 353606 03/29/11 15:48 a-onl Start_Run_3130,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353605 03/29/11 15:38 a-onl End_of_Run_3129
  • 353604 03/29/11 14:59 a-onl Start_Run_3129,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353603 03/29/11 14:58 a-onl End_of_Run_3128
  • 353602 03/29/11 14:26 gilman Hall A tools for shift checklist
  • 353601 03/29/11 14:10 gilman spot check for shift checklist
  • 353600 03/29/11 13:57 gilman beam will be down about 3:30
  • 353599 03/29/11 13:31 a-onl Start_Run_3128,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353598 03/29/11 13:30 a-onl End_of_Run_3127
  • 353597 03/29/11 12:32 auto_snap Radiation and Ozone snapshot
  • 353596 03/29/11 12:32 auto_snap Miscellaneous snapshot
  • 353595 03/29/11 12:32 auto_snap Hall A Tools snapshot
  • 353594 03/29/11 12:32 auto_snap Compton Polarimeter snapshot
  • 353593 03/29/11 12:32 auto_snap BigBox Magnet snapshot
  • 353592 03/29/11 12:32 auto_snap Beam Position and Current snapshot
  • 353591 03/29/11 12:07 a-onl Start_Run_3127,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353590 03/29/11 12:06 a-onl End_of_Run_3126
  • 353589 03/29/11 11:46 yez 1st and 3rd figure is for run3125, 2nd one is for run3005,
  • 353588 03/29/11 11:42 yez LHRS VDC has a little bit low efficiency
  • 353587 03/29/11 11:34 orchen BB triple coincedence - 1.4C, 46-26 events
  • 353586 03/29/11 10:39 a-onl Start_Run_3126,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353585 03/29/11 10:39 a-onl End_of_Run_3125
  • 353584 03/29/11 09:15 Norum L-10 trip
  • 353583 03/29/11 09:15 a-onl Start_Run_3125,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353582 03/29/11 09:13 gilman sending beam back
  • 353581 03/29/11 09:04 gilman MCC stopping beam
  • 353580 03/29/11 09:04 a-onl End_of_Run_3124
  • 353579 03/29/11 08:22 a-onl Start_Run_3124,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353578 03/29/11 08:14 a-onl End_of_Run_3123
  • 353577 03/29/11 08:05 adaq Beam Time Accounting
  • 353576 03/29/11 07:56 arie shift summary
  • 353575 03/29/11 07:38 israely Target screenshot
  • 353574 03/29/11 07:23 arie general tools
  • 353573 03/29/11 07:20 arie spot++
  • 353572 03/29/11 06:48 a-onl Start_Run_3123,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353571 03/29/11 06:44 a-onl End_of_Run_3122
  • 353570 03/29/11 06:01 a-onl Start_Run_3122,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353569 03/29/11 06:00 a-onl End_of_Run_3121
  • 353568 03/29/11 05:47 Hisham Target status
  • 353567 03/29/11 04:37 a-onl Start_Run_3121,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353566 03/29/11 04:36 a-onl End_of_Run_3120
  • 353565 03/29/11 04:32 auto_snap Radiation and Ozone snapshot
  • 353564 03/29/11 04:32 auto_snap Miscellaneous snapshot
  • 353563 03/29/11 04:32 auto_snap Hall A Tools snapshot
  • 353562 03/29/11 04:32 auto_snap Compton Polarimeter snapshot
  • 353561 03/29/11 04:32 auto_snap BigBox Magnet snapshot
  • 353560 03/29/11 04:32 auto_snap Beam Position and Current snapshot
  • 353559 03/29/11 03:17 igorko N6 - 1,2 L
  • 353558 03/29/11 03:10 a-onl Start_Run_3120,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353557 03/29/11 03:05 a-onl End_of_Run_3119
  • 353556 03/29/11 03:01 israely We are back to production from run 3119
  • 353555 03/29/11 02:56 Hisham Target moved back to He4
  • 353554 03/29/11 02:51 a-onl Start_Run_3119,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production 4He 20 cm Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353553 03/29/11 02:44 arie spot_r
  • 353552 03/29/11 02:42 Hisham BeO target for BR
  • 353551 03/29/11 02:41 igorko Pmiss so far
  • 353550 03/29/11 02:39 a-onl End_of_Run_3118
  • 353549 03/29/11 02:38 a-onl Start_Run_3118,Run_type=Spot++,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Beam Recovery,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353548 03/29/11 01:41 eip total charge
  • 353547 03/29/11 01:38 igorko HAND N6-1L and N6 -2L problem
  • 353546 03/29/11 01:12 israely No Beam
  • 353545 03/29/11 01:11 igorko triple so far
  • 353544 03/29/11 00:44 israely We currently have problem with the beam.
  • 353543 03/29/11 00:40 a-onl End_of_Run_3117
  • 353542 03/29/11 00:37 a-onl Start_Run_3117,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production SRC,target_type=4He 20cm,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353541 03/29/11 00:34 a-onl End_of_Run_3116
  • 353540 03/29/11 00:24 israely Prescales screenshots
  • 353539 03/29/11 00:15 israely Neutron detector screenshot
  • 353538 03/29/11 00:13 israely Bigbite scintillator screenshot
  • 353537 03/29/11 00:09 a-onl Start_Run_3116,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production SRC,target_type=4He 20cm,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353536 03/29/11 00:08 a-onl End_of_Run_3115
  • 353535 03/29/11 00:07 Hisham Target status
  • 353534 03/29/11 00:02 adaq Beam Time Accounting
  • 353533 03/29/11 00:02 igorko difference between runs
  • Monday

  • 353532 03/28/11 23:57 Carlos Munoz Swing shift summary
  • 353531 03/28/11 22:57 adaq Hall A tools
  • 353530 03/28/11 22:56 adaq spot++
  • 353529 03/28/11 22:39 a-onl Start_Run_3115,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production SRC,target_type=4He 20cm,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353528 03/28/11 22:38 a-onl End_of_Run_3114
  • 353527 03/28/11 21:12 Carlos bbps1:2001 S13 L10.2 tripped again
  • 353526 03/28/11 20:58 Carlos bbps1:2001 S13 L10.2 tripped
  • 353525 03/28/11 20:57 a-onl Start_Run_3114,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production SRC,target_type=4He 20cm,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353524 03/28/11 20:56 a-onl End_of_Run_3113
  • 353523 03/28/11 20:35 a-onl Start_Run_3113,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production SRC,target_type=4He 20cm,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353522 03/28/11 20:34 a-onl End_of_Run_3112
  • 353521 03/28/11 20:32 auto_snap Radiation and Ozone snapshot
  • 353520 03/28/11 20:32 auto_snap Miscellaneous snapshot
  • 353519 03/28/11 20:32 auto_snap Hall A Tools snapshot
  • 353518 03/28/11 20:32 auto_snap Compton Polarimeter snapshot
  • 353517 03/28/11 20:32 auto_snap BigBox Magnet snapshot
  • 353516 03/28/11 20:31 auto_snap Beam Position and Current snapshot
  • 353515 03/28/11 18:55 a-onl Start_Run_3112,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production SRC,target_type=4He 20cm,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353514 03/28/11 18:48 a-onl End_of_Run_3111
  • 353513 03/28/11 18:19 JinGe target screen
  • 353512 03/28/11 17:17 a-onl Start_Run_3111,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production SRC,target_type=4He 20cm,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353511 03/28/11 17:16 a-onl End_of_Run_3110
  • 353510 03/28/11 16:05 Sarah Day shift run list (analysis)
  • 353509 03/28/11 16:01 adaq Beam Time Accounting
  • 353508 03/28/11 16:00 dalton Day Shift Summary
  • 353507 03/28/11 16:00 dalton Day Shift Summary
  • 353506 03/28/11 15:51 a-onl Start_Run_3110,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production SRC,target_type=4He 20cm,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353505 03/28/11 15:50 a-onl End_of_Run_3109
  • 353504 03/28/11 15:34 yez RHRS: 0 current of HV slot#15.#4, which is Lumi004
  • 353503 03/28/11 15:21 yez RHRS: All HVs are on now.
  • 353502 03/28/11 14:45 dalton Right arm HVs off
  • 353501 03/28/11 14:27 a-onl Start_Run_3109,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production SRC,target_type=4He 20cm,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353500 03/28/11 14:26 a-onl End_of_Run_3108
  • 353499 03/28/11 14:13 dalton Right arm VDC gas flow unstable
  • 353498 03/28/11 14:00 Sarah Spot_R plots for run 3108 (for shift checklist)
  • 353497 03/28/11 13:00 a-onl Start_Run_3108,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production SRC,target_type=4He 20cm,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353496 03/28/11 12:59 a-onl End_of_Run_3107
  • 353495 03/28/11 12:31 auto_snap Radiation and Ozone snapshot
  • 353494 03/28/11 12:31 auto_snap Miscellaneous snapshot
  • 353493 03/28/11 12:31 auto_snap Hall A Tools snapshot
  • 353492 03/28/11 12:31 auto_snap Compton Polarimeter snapshot
  • 353491 03/28/11 12:31 auto_snap BigBox Magnet snapshot
  • 353490 03/28/11 12:31 auto_snap Beam Position and Current snapshot
  • 353489 03/28/11 12:16 Sarah Weird histogram in replay of run 3107
  • 353488 03/28/11 11:46 dalton Another L10.2 trip
  • 353487 03/28/11 11:38 dalton T3 rate on 20 cm 4He
  • 353486 03/28/11 11:37 a-onl Start_Run_3107,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production SRC,target_type=4He 20cm ,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353485 03/28/11 11:29 a-onl End_of_Run_3106
  • 353484 03/28/11 11:21 J. Zhang N5 Bar3 hv tripped
  • 353483 03/28/11 11:15 a-onl Start_Run_3106,Run_type=Production SRC,target_type=20cm dummy,beam_e0=4.46,beam_raster=on,beam_current=6,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Dummy target run, 15 cm aluminum dummy target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353482 03/28/11 11:14 a-onl End_of_Run_3105
  • 353481 03/28/11 11:14 Sarah Plane 5, right bar 3 tripped off
  • 353480 03/28/11 11:12 igorko N5-3R trip
  • 353479 03/28/11 10:42 dalton T3 rate on dummy
  • 353478 03/28/11 10:34 Sarah Spot Plots for Run 3105
  • 353477 03/28/11 10:34 dalton Beam is back
  • 353476 03/28/11 10:30 a-onl Start_Run_3105,Run_type=Production SRC,target_type=20cm dummy,beam_e0=4.46,beam_raster=on,beam_current=3,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Dummy target run, 15 cm aluminum dummy target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353475 03/28/11 10:28 a-onl End_of_Run_3104
  • 353474 03/28/11 10:16 yez RHRS s2 ADC are normal
  • 353473 03/28/11 10:02 a-onl Start_Run_3104,Run_type=Cosmics,target_type=20cm dummy,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=cosmics (no beam), 15 cm aluminum dummy target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353472 03/28/11 09:53 a-onl End_of_Run_3103
  • 353471 03/28/11 09:49 Sarah R-arm S2 ADCs
  • 353470 03/28/11 09:08 wenbiao target status screen
  • 353469 03/28/11 09:07 adaq Beam Time Accounting
  • 353468 03/28/11 08:49 dalton Left arm dipole alarm
  • 353467 03/28/11 08:39 Sarah Snapshot of the scaler screen
  • 353466 03/28/11 08:36 a-onl Start_Run_3103,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353465 03/28/11 08:35 dalton Friday trip with picture
  • 353464 03/28/11 08:32 dalton RHRS Q2 cycled
  • 353463 03/28/11 08:28 dalton Right dipole trip on Friday
  • 353462 03/28/11 08:26 Sarah Stopped run 3102 to cycle right Q2.
  • 353461 03/28/11 08:25 a-onl End_of_Run_3102
  • 353460 03/28/11 08:18 Sarah Run 3101 is junk
  • 353459 03/28/11 08:17 dalton Right arm Q2 was down for 20 mins
  • 353458 03/28/11 08:17 a-onl Start_Run_3102,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He ,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353457 03/28/11 08:16 a-onl End_of_Run_3101
  • 353456 03/28/11 08:15 Sarah/Dalton Right Arm Quadrupole 2 tripped
  • 353455 03/28/11 08:15 dalton Right arm Q2 coming back up
  • 353454 03/28/11 08:14 a-onl Start_Run_3101,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He JUNK RUN!!!!! ,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353453 03/28/11 08:13 a-onl End_of_Run_3100
  • 353452 03/28/11 08:02 israely coda screenshot
  • 353451 03/28/11 08:02 arie shift summary
  • 353450 03/28/11 07:55 wenbiao target
  • 353449 03/28/11 07:25 Reimer Target Status Screens
  • 353448 03/28/11 07:21 arie general tools
  • 353447 03/28/11 07:13 arie spot++
  • 353446 03/28/11 07:00 a-onl Start_Run_3100,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353445 03/28/11 06:58 a-onl End_of_Run_3099
  • 353444 03/28/11 05:36 a-onl Start_Run_3099,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353443 03/28/11 05:35 a-onl End_of_Run_3098
  • 353442 03/28/11 04:31 auto_snap Radiation and Ozone snapshot
  • 353441 03/28/11 04:31 auto_snap Miscellaneous snapshot
  • 353440 03/28/11 04:31 auto_snap Hall A Tools snapshot
  • 353439 03/28/11 04:31 auto_snap Compton Polarimeter snapshot
  • 353438 03/28/11 04:31 auto_snap BigBox Magnet snapshot
  • 353437 03/28/11 04:31 auto_snap Beam Position and Current snapshot
  • 353436 03/28/11 04:13 a-onl Start_Run_3098,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353435 03/28/11 04:12 a-onl End_of_Run_3097
  • 353434 03/28/11 02:50 a-onl Start_Run_3097,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353433 03/28/11 02:49 a-onl End_of_Run_3096
  • 353432 03/28/11 01:24 a-onl Start_Run_3096,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353431 03/28/11 01:20 a-onl Start_Run_3095,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353430 03/28/11 01:20 a-onl End_of_Run_3094
  • 353429 03/28/11 00:59 igorko Pmiss
  • 353428 03/28/11 00:31 igorko currents cont.
  • 353427 03/28/11 00:23 igorko triple coincidence
  • 353426 03/28/11 00:02 adaq Beam Time Accounting
  • Sunday

  • 353425 03/27/11 23:58 LeRose Shift Summary
  • 353424 03/27/11 23:58 a-onl Start_Run_3094,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353423 03/27/11 23:56 a-onl End_of_Run_3093
  • 353422 03/27/11 23:45 hisham Target status
  • 353421 03/27/11 22:33 a-onl Start_Run_3093,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353420 03/27/11 22:33 a-onl End_of_Run_3092
  • 353419 03/27/11 21:10 igorko different currents
  • 353418 03/27/11 21:09 a-onl Start_Run_3092,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353417 03/27/11 21:09 a-onl End_of_Run_3091
  • 353416 03/27/11 20:31 auto_snap Radiation and Ozone snapshot
  • 353415 03/27/11 20:31 auto_snap Miscellaneous snapshot
  • 353414 03/27/11 20:31 auto_snap Hall A Tools snapshot
  • 353413 03/27/11 20:31 auto_snap Compton Polarimeter snapshot
  • 353412 03/27/11 20:31 auto_snap BigBox Magnet snapshot
  • 353411 03/27/11 20:30 auto_snap Beam Position and Current snapshot
  • 353410 03/27/11 20:10 Hisham Target status
  • 353409 03/27/11 19:43 a-onl Start_Run_3091,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353408 03/27/11 19:43 a-onl End_of_Run_3090
  • 353407 03/27/11 19:02 igorko ROC sync
  • 353406 03/27/11 18:50 LeRose Hall A Tools
  • 353405 03/27/11 18:46 LeRose Spot++
  • 353404 03/27/11 18:19 a-onl Start_Run_3090,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353403 03/27/11 18:19 a-onl End_of_Run_3089
  • 353402 03/27/11 16:56 a-onl Start_Run_3089,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353401 03/27/11 16:55 a-onl End_of_Run_3088
  • 353400 03/27/11 16:15 Hisham Target status
  • 353399 03/27/11 16:10 adaq Beam Time Accounting
  • 353398 03/27/11 16:09 Boeglin Shift Summary
  • 353397 03/27/11 15:41 reimer target screen shots
  • 353396 03/27/11 15:36 israely Hall A screenshots
  • 353395 03/27/11 15:30 a-onl Start_Run_3088,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353394 03/27/11 15:29 a-onl End_of_Run_3087
  • 353393 03/27/11 14:06 a-onl Start_Run_3087,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353392 03/27/11 14:06 a-onl End_of_Run_3086
  • 353391 03/27/11 13:32 igorko HAND triple so far
  • 353390 03/27/11 12:40 a-onl Start_Run_3086,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353389 03/27/11 12:34 a-onl Start_Run_3085,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353388 03/27/11 12:33 a-onl End_of_Run_3084
  • 353387 03/27/11 12:30 auto_snap Radiation and Ozone snapshot
  • 353386 03/27/11 12:30 auto_snap Miscellaneous snapshot
  • 353385 03/27/11 12:30 auto_snap Hall A Tools snapshot
  • 353384 03/27/11 12:30 auto_snap Compton Polarimeter snapshot
  • 353383 03/27/11 12:30 auto_snap BigBox Magnet snapshot
  • 353382 03/27/11 12:30 auto_snap Beam Position and Current snapshot
  • 353381 03/27/11 11:54 Boeglin Hall A Tools snapshot
  • 353380 03/27/11 11:52 Boeglin Spot++
  • 353379 03/27/11 11:10 a-onl Start_Run_3084,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353378 03/27/11 11:09 a-onl End_of_Run_3083
  • 353377 03/27/11 09:43 a-onl Start_Run_3083,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353376 03/27/11 09:42 a-onl End_of_Run_3082
  • 353375 03/27/11 08:25 arie neutron singles
  • 353374 03/27/11 08:24 arie scalers
  • 353373 03/27/11 08:13 a-onl Start_Run_3082,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353372 03/27/11 08:12 a-onl End_of_Run_3081
  • 353371 03/27/11 08:01 adaq Beam Time Accounting
  • 353370 03/27/11 08:01 ipom Owl shift summary
  • 353369 03/27/11 07:25 ipom spot++
  • 353368 03/27/11 06:47 a-onl Start_Run_3081,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353367 03/27/11 06:46 a-onl End_of_Run_3080
  • 353366 03/27/11 06:02 D.Wang target status
  • 353365 03/27/11 05:20 arie neutrons rates
  • 353364 03/27/11 05:18 arie scalers
  • 353363 03/27/11 05:17 a-onl Start_Run_3080,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353362 03/27/11 05:16 a-onl End_of_Run_3079
  • 353361 03/27/11 04:30 auto_snap Radiation and Ozone snapshot
  • 353360 03/27/11 04:30 auto_snap Miscellaneous snapshot
  • 353359 03/27/11 04:30 auto_snap Hall A Tools snapshot
  • 353358 03/27/11 04:30 auto_snap Compton Polarimeter snapshot
  • 353357 03/27/11 04:30 auto_snap BigBox Magnet snapshot
  • 353356 03/27/11 04:30 auto_snap Beam Position and Current snapshot
  • 353355 03/27/11 04:29 D.Wang target status
  • 353354 03/27/11 04:21 a-onl Start_Run_3079,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353353 03/27/11 04:20 a-onl End_of_Run_3078
  • 353352 03/27/11 02:58 a-onl Start_Run_3078,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353351 03/27/11 02:57 a-onl End_of_Run_3077
  • 353350 03/27/11 02:44 arie beck neutron rates
  • 353349 03/27/11 02:42 arie beck scalers
  • 353348 03/27/11 02:28 D.Wang target status
  • 353347 03/27/11 01:34 a-onl Start_Run_3077,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353346 03/27/11 01:33 a-onl End_of_Run_3076
  • 353345 03/27/11 01:33 eip optimal beam current ?
  • 353344 03/27/11 01:30 ipom I misstyped: coda died on run #3075
  • 353343 03/27/11 01:16 igorko e,e'N_recoil - different currents
  • 353342 03/27/11 01:15 igorko e,e'N_recoil - different currents
  • 353341 03/27/11 00:53 ipom coda died on Run 3076
  • 353340 03/27/11 00:50 jixie Make sure histograms are in the same binning
  • 353339 03/27/11 00:29 Jixie Zhang To Shift Leader and 3rd person: Please check the End-of-Run Log to make sure the DAQ work properly
  • 353338 03/27/11 00:09 a-onl Start_Run_3076,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353337 03/27/11 00:05 a-onl Start_Run_3075,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353336 03/27/11 00:03 adaq Beam Time Accounting
  • 353335 03/27/11 00:02 D.Wang target status
  • Saturday

  • 353334 03/26/11 23:56 a-onl End_of_Run_3074
  • 353333 03/26/11 23:56 Carlos Munoz Swing shift summary
  • 353332 03/26/11 23:04 igorko update
  • 353331 03/26/11 22:50 M Defurne screeshots
  • 353330 03/26/11 22:50 Carlos Hall A tools
  • 353329 03/26/11 22:48 igorko (e,e'N_recoil) for 5 uA
  • 353328 03/26/11 22:48 adaq spot++
  • 353327 03/26/11 22:34 a-onl Start_Run_3074,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353326 03/26/11 22:33 a-onl End_of_Run_3073
  • 353325 03/26/11 21:09 a-onl Start_Run_3073,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353324 03/26/11 21:08 a-onl End_of_Run_3072
  • 353323 03/26/11 20:30 auto_snap Radiation and Ozone snapshot
  • 353322 03/26/11 20:30 auto_snap Miscellaneous snapshot
  • 353321 03/26/11 20:30 auto_snap Hall A Tools snapshot
  • 353320 03/26/11 20:30 auto_snap Compton Polarimeter snapshot
  • 353319 03/26/11 20:30 auto_snap BigBox Magnet snapshot
  • 353318 03/26/11 20:29 auto_snap Beam Position and Current snapshot
  • 353317 03/26/11 20:04 M Defurne scrennshots
  • 353316 03/26/11 19:42 a-onl Start_Run_3072,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353315 03/26/11 19:40 a-onl End_of_Run_3071
  • 353314 03/26/11 19:39 nmuangma bigibite calibration continue...
  • 353313 03/26/11 18:20 a-onl Start_Run_3071,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353312 03/26/11 18:19 a-onl End_of_Run_3070
  • 353311 03/26/11 17:56 M Defurne Screnshots- Heat Power Ok
  • 353309 03/26/11 16:53 a-onl End_of_Run_3069
  • 353308 03/26/11 16:02 adaq Beam Time Accounting
  • 353307 03/26/11 16:01 Boeglin Shift Summary
  • 353306 03/26/11 15:34 a-onl Start_Run_3069,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He ,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353305 03/26/11 15:32 a-onl End_of_Run_3068
  • 353304 03/26/11 14:59 boeglin spot++
  • 353303 03/26/11 14:31 Boeglin End of run info problem
  • 353302 03/26/11 14:12 a-onl Start_Run_3068,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He increased current to 5uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353301 03/26/11 14:09 a-onl End_of_Run_3067
  • 353300 03/26/11 12:46 a-onl Start_Run_3067,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353299 03/26/11 12:45 a-onl End_of_Run_3066
  • 353298 03/26/11 12:44 orchen Different bining for BB and link to (e,e'p_recoil) for time comparison
  • 353297 03/26/11 12:44 a-onl Start_Run_3066,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353296 03/26/11 12:40 orchen & Igo (e,e'N_recoil) for 2, 4, and 8 uA
  • 353295 03/26/11 12:38 a-onl End_of_Run_3065
  • 353294 03/26/11 12:34 reimer Target
  • 353293 03/26/11 12:29 auto_snap Radiation and Ozone snapshot
  • 353292 03/26/11 12:29 auto_snap Miscellaneous snapshot
  • 353291 03/26/11 12:29 auto_snap Hall A Tools snapshot
  • 353290 03/26/11 12:29 auto_snap Compton Polarimeter snapshot
  • 353289 03/26/11 12:29 auto_snap BigBox Magnet snapshot
  • 353288 03/26/11 12:29 auto_snap Beam Position and Current snapshot
  • 353287 03/26/11 12:04 reimer target cernox
  • 353286 03/26/11 12:01 Reimer target cernox alarm
  • 353285 03/26/11 11:41 yez L-HRS: Cer Number of after HVs change
  • 353284 03/26/11 11:39 orchen Analysis results for 4He(e,e'pp) and 4He(e,e'pn) for 625MeV/c kinematics
  • 353283 03/26/11 11:35 a-onl Start_Run_3065,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353282 03/26/11 11:34 a-onl End_of_Run_3064
  • 353281 03/26/11 11:07 wenbiao neutron detector single rates
  • 353280 03/26/11 11:01 wenbiao scalers
  • 353279 03/26/11 10:53 yez L-HRS: Cer SPE before and after HVs change
  • 353278 03/26/11 10:10 a-onl Start_Run_3064,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353277 03/26/11 10:08 a-onl End_of_Run_3063
  • 353276 03/26/11 08:44 a-onl Start_Run_3063,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353275 03/26/11 08:12 a-onl End_of_Run_3062
  • 353274 03/26/11 08:10 ipom Owl shift summary
  • 353273 03/26/11 08:05 adaq Beam Time Accounting
  • 353272 03/26/11 07:53 D.Wang target status
  • 353271 03/26/11 07:17 a-onl Start_Run_3062,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353270 03/26/11 07:12 a-onl End_of_Run_3061
  • 353269 03/26/11 06:58 arie beck neutron single rate
  • 353268 03/26/11 06:56 arie beck scalers
  • 353267 03/26/11 06:48 ipom CODA died, run 3060 is junk
  • 353266 03/26/11 06:45 a-onl Start_Run_3061,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353265 03/26/11 06:38 a-onl Start_Run_3060,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353264 03/26/11 06:37 a-onl End_of_Run_3059
  • 353263 03/26/11 06:29 D.Wang target status
  • 353262 03/26/11 05:18 arie beck nuetron single
  • 353261 03/26/11 05:17 arie beck scalers
  • 353260 03/26/11 05:06 a-onl Start_Run_3059,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353259 03/26/11 05:04 a-onl End_of_Run_3058
  • 353258 03/26/11 04:29 auto_snap Radiation and Ozone snapshot
  • 353257 03/26/11 04:29 auto_snap Miscellaneous snapshot
  • 353256 03/26/11 04:29 auto_snap Hall A Tools snapshot
  • 353255 03/26/11 04:29 auto_snap Compton Polarimeter snapshot
  • 353254 03/26/11 04:29 auto_snap BigBox Magnet snapshot
  • 353253 03/26/11 04:29 auto_snap Beam Position and Current snapshot
  • 353252 03/26/11 04:18 D.Wang target status
  • 353251 03/26/11 04:11 nmuangma 2nd t0 calibration: modify in db_BB.mwdc.dat
  • 353250 03/26/11 04:04 arie beck neutron singles rate
  • 353249 03/26/11 04:02 arie beck scaler
  • 353248 03/26/11 03:40 a-onl Start_Run_3058,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353247 03/26/11 03:39 a-onl End_of_Run_3056
  • 353246 03/26/11 03:37 a-onl End_of_Run_3056
  • 353245 03/26/11 02:55 arie beck neutron detector single rates
  • 353244 03/26/11 02:54 arie beck scalers screenshot
  • 353243 03/26/11 02:39 a-onl Start_Run_3056,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=2,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353242 03/26/11 02:37 a-onl End_of_Run_3055
  • 353241 03/26/11 01:57 D.Wang target status
  • 353240 03/26/11 01:35 ipom going for 8uA
  • 353239 03/26/11 01:32 a-onl Start_Run_3055,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=8,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He ,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353238 03/26/11 01:31 a-onl End_of_Run_3054
  • 353237 03/26/11 01:31 eip adc change
  • 353236 03/26/11 01:30 a-onl Start_Run_3054,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=8,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He - Test run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353235 03/26/11 01:27 nmuangma result of wirechamber calibration with T0
  • 353234 03/26/11 01:26 a-onl End_of_Run_3053
  • 353233 03/26/11 01:00 yez L-HRS: Cer ADCs are still overfload after I reduced the HVs
  • 353232 03/26/11 00:16 eip double coinc rate
  • 353231 03/26/11 00:09 yez L-HRS: Reduce HVs of Cerenkov and create a new data base ~/20110326
  • 353230 03/26/11 00:08 adaq Beam Time Accounting
  • 353229 03/26/11 00:06 a-onl Start_Run_3053,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4.5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353228 03/26/11 00:05 a-onl End_of_Run_3052
  • Friday

  • 353227 03/25/11 23:58 D.Wang target status
  • 353226 03/25/11 23:54 Carlos Munoz Swing shift summary
  • 353225 03/25/11 23:39 igorko HAND scalers
  • 353224 03/25/11 23:38 igorko scalers
  • 353223 03/25/11 23:24 igorko double coinc check
  • 353222 03/25/11 23:02 a-onl Start_Run_3052,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4.5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353221 03/25/11 23:01 a-onl End_of_Run_3051
  • 353220 03/25/11 22:36 igorko Singles Rates
  • 353219 03/25/11 22:30 Carlos Hall A tools
  • 353218 03/25/11 22:29 Carlos Spot ++
  • 353217 03/25/11 22:28 Hisham Target status
  • 353216 03/25/11 21:44 a-onl Start_Run_3051,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4.5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353215 03/25/11 21:44 Carlos RQ2 back
  • 353214 03/25/11 21:33 Carlos Right Q2
  • 353213 03/25/11 21:26 a-onl End_of_Run_3050
  • 353212 03/25/11 21:15 narbe Run Numbers Not Matching On CODA
  • 353211 03/25/11 21:12 a-onl Start_Run_3050,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4.5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353210 03/25/11 21:10 a-onl End_of_Run_3049
  • 353209 03/25/11 20:29 auto_snap Radiation and Ozone snapshot
  • 353208 03/25/11 20:29 auto_snap Miscellaneous snapshot
  • 353207 03/25/11 20:29 auto_snap Hall A Tools snapshot
  • 353206 03/25/11 20:29 auto_snap Compton Polarimeter snapshot
  • 353205 03/25/11 20:29 auto_snap BigBox Magnet snapshot
  • 353204 03/25/11 20:28 auto_snap Beam Position and Current snapshot
  • 353203 03/25/11 19:54 a-onl Start_Run_3049,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4.5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353202 03/25/11 19:38 a-onl End_of_Run_3048
  • 353201 03/25/11 19:38 Hisham Target Status
  • 353200 03/25/11 19:30 Jixie Zhang script to sum BCM U1 charge using halog files
  • 353199 03/25/11 18:42 Carlos Beam current increased to 4.5 uA
  • 353198 03/25/11 18:42 a-onl Start_Run_3048,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4.5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353197 03/25/11 18:38 a-onl End_of_Run_3047
  • 353196 03/25/11 17:51 israely neutron detector screenshots
  • 353195 03/25/11 17:51 israely Bigbite scintilator & dcaler screenshot
  • 353194 03/25/11 17:21 a-onl Start_Run_3047,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353193 03/25/11 17:20 a-onl End_of_Run_3046
  • 353192 03/25/11 16:04 Hisham Target status
  • 353191 03/25/11 16:02 adaq Beam Time Accounting
  • 353190 03/25/11 16:00 Boeglin Shift Summary
  • 353189 03/25/11 15:58 a-onl Start_Run_3046,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353188 03/25/11 15:57 a-onl End_of_Run_3045
  • 353187 03/25/11 15:42 yez map of Raster Beam is corrected.
  • 353186 03/25/11 15:37 yez cable map of Raster Beam is still wrong
  • 353185 03/25/11 15:05 dalton Target is stable
  • 353184 03/25/11 14:37 a-onl Start_Run_3045,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353183 03/25/11 14:36 a-onl End_of_Run_3044
  • 353182 03/25/11 13:13 a-onl Start_Run_3044,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353181 03/25/11 13:12 Boeglin Hall A Tools
  • 353180 03/25/11 13:12 a-onl End_of_Run_3043
  • 353179 03/25/11 13:11 adaq spot++
  • 353178 03/25/11 12:44 dalton Target is stable
  • 353177 03/25/11 12:28 auto_snap Radiation and Ozone snapshot
  • 353176 03/25/11 12:28 auto_snap Miscellaneous snapshot
  • 353175 03/25/11 12:28 auto_snap Hall A Tools snapshot
  • 353174 03/25/11 12:28 auto_snap Compton Polarimeter snapshot
  • 353173 03/25/11 12:28 auto_snap BigBox Magnet snapshot
  • 353172 03/25/11 12:28 auto_snap Beam Position and Current snapshot
  • 353171 03/25/11 12:21 igorko beam current
  • 353170 03/25/11 12:17 eip beam current
  • 353169 03/25/11 11:53 Punjabi spot check
  • 353168 03/25/11 11:52 a-onl Start_Run_3043,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=production SRC on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353167 03/25/11 11:50 a-onl End_of_Run_3042
  • 353166 03/25/11 11:49 a-onl Start_Run_3042,Run_type=Spot++,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=20cm 4He, to check spot,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353165 03/25/11 11:46 a-onl Start_Run_3041,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=20cm 4He, to check spot,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353164 03/25/11 11:20 dalton Target recovered
  • 353163 03/25/11 11:11 Boeglin Going to Beam Permit
  • 353162 03/25/11 10:49 yez New Data Base 20110324/ has been created for HVs change of GC
  • 353161 03/25/11 10:13 Boeglin Hall A going to Controlled access
  • 353160 03/25/11 10:05 israely There is no beam and It will take arround 3h to fix the target
  • 353159 03/25/11 09:46 yez L-HRS: Cerenkov ADCs vs PRL after reducing HVs
  • 353158 03/25/11 09:43 israely Lost the he4 Target.
  • 353157 03/25/11 09:43 dalton Loop 1 target event
  • 353156 03/25/11 09:37 a-onl End_of_Run_3040
  • 353155 03/25/11 09:02 dalton Target Status
  • 353154 03/25/11 08:59 israely Target screenshot
  • 353153 03/25/11 08:58 israely neutron detector screenshots
  • 353152 03/25/11 08:57 israely Bigbite scintilator & scaler screenshot
  • 353151 03/25/11 08:57 israely Coda screenshots
  • 353150 03/25/11 08:30 a-onl Start_Run_3040,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production Run on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353149 03/25/11 08:29 a-onl End_of_Run_3039
  • 353148 03/25/11 08:06 a-onl Start_Run_3039,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production Run on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353147 03/25/11 08:05 adaq Beam Time Accounting
  • 353146 03/25/11 08:04 a-onl End_of_Run_3038
  • 353145 03/25/11 07:50 ipom Owl shift summary
  • 353144 03/25/11 07:32 D.Wang target status
  • 353143 03/25/11 06:37 a-onl Start_Run_3038,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production Run on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353142 03/25/11 06:35 a-onl End_of_Run_3037
  • 353141 03/25/11 06:10 orchen missing momentum distributions for 500MeV/c and 625MeV/c kinematics
  • 353140 03/25/11 05:56 D.Wang target status
  • 353139 03/25/11 05:10 a-onl Start_Run_3037,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production Run on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353138 03/25/11 05:09 a-onl End_of_Run_3036
  • 353137 03/25/11 04:28 auto_snap Radiation and Ozone snapshot
  • 353136 03/25/11 04:28 auto_snap Miscellaneous snapshot
  • 353135 03/25/11 04:28 auto_snap Hall A Tools snapshot
  • 353134 03/25/11 04:28 auto_snap Compton Polarimeter snapshot
  • 353133 03/25/11 04:28 auto_snap BigBox Magnet snapshot
  • 353132 03/25/11 04:28 auto_snap Beam Position and Current snapshot
  • 353131 03/25/11 03:55 D.Wang target status
  • 353130 03/25/11 03:49 a-onl Start_Run_3036,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production Run on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353129 03/25/11 03:47 a-onl End_of_Run_3035
  • 353128 03/25/11 02:26 a-onl Start_Run_3035,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production Run on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353127 03/25/11 02:24 a-onl End_of_Run_3034
  • 353126 03/25/11 02:10 D.Wang target status
  • 353125 03/25/11 01:03 a-onl Start_Run_3034,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production Run on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353124 03/25/11 01:02 a-onl End_of_Run_3033
  • 353123 03/25/11 00:30 israely coda screenshot
  • 353122 03/25/11 00:30 igorko & orc e,e'p for 625 MeV/c
  • 353121 03/25/11 00:29 igorko & orc e,e'p for 625 MeV/c
  • 353120 03/25/11 00:04 D. Wang target status
  • 353119 03/25/11 00:04 adaq Beam Time Accounting
  • 353118 03/25/11 00:01 israely Coda screenshot
  • Thursday

  • 353117 03/24/11 23:58 Boeglin Shift Summary
  • 353116 03/24/11 23:41 a-onl Start_Run_3033,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production Run on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353115 03/24/11 23:40 a-onl End_of_Run_3032
  • 353114 03/24/11 22:59 Carlos Muno Target screens
  • 353113 03/24/11 22:19 igorko double coincidence for 625 MeV/c
  • 353112 03/24/11 22:19 a-onl Start_Run_3032,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production Run on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353111 03/24/11 22:16 a-onl End_of_Run_3031
  • 353110 03/24/11 21:56 a-onl Start_Run_3031,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production Run on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353109 03/24/11 21:54 a-onl End_of_Run_3030
  • 353108 03/24/11 20:38 Boeglin Hall A Tools screen
  • 353107 03/24/11 20:36 adaq spot_R for run 3030
  • 353106 03/24/11 20:36 a-onl Start_Run_3030,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production Run on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353105 03/24/11 20:33 adaq strange spot_R
  • 353104 03/24/11 20:30 a-onl End_of_Run_3029
  • 353103 03/24/11 20:28 auto_snap Radiation and Ozone snapshot
  • 353102 03/24/11 20:28 auto_snap Miscellaneous snapshot
  • 353101 03/24/11 20:28 auto_snap Hall A Tools snapshot
  • 353100 03/24/11 20:28 auto_snap Compton Polarimeter snapshot
  • 353099 03/24/11 20:28 auto_snap BigBox Magnet snapshot
  • 353098 03/24/11 20:28 auto_snap Beam Position and Current snapshot
  • 353097 03/24/11 19:59 a-onl Start_Run_3029,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production Run on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353096 03/24/11 19:58 a-onl End_of_Run_3028
  • 353095 03/24/11 18:38 a-onl Start_Run_3028,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production Run on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353094 03/24/11 18:30 yez A correction for my last halog
  • 353093 03/24/11 18:27 a-onl End_of_Run_3027
  • 353092 03/24/11 18:19 yez L-HRS: Cerenkov PID Efficiency before and after HVs change
  • 353091 03/24/11 17:56 a-onl Start_Run_3027,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production Run on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353090 03/24/11 17:51 a-onl End_of_Run_3026
  • 353089 03/24/11 17:46 yez This is the db before I changed the HVs this morning. I have not calibrated new db for current HV setting
  • 353088 03/24/11 17:35 nmuangma should't you make a new date for the data base?
  • 353087 03/24/11 17:17 yez L-HRS: Cerenkov ADCs are still overfload after reducing HVs
  • 353086 03/24/11 16:29 a-onl Start_Run_3026,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production Run on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353085 03/24/11 16:22 yez Data Base of LHRS Cer, PRL1&2 are updated in 20100310/
  • 353084 03/24/11 16:05 Solvignon Day Shift Summary
  • 353083 03/24/11 16:05 adaq Beam Time Accounting
  • 353082 03/24/11 15:57 Solvignon Correction: LQ2 turned itself off
  • 353081 03/24/11 15:48 Solvignon LQ2 was not a trip but somebody turned it off.
  • 353080 03/24/11 15:46 Fansler LQ2
  • 353079 03/24/11 15:35 Solvignon Left Q2 tripped
  • 353078 03/24/11 15:32 a-onl End_of_Run_3025
  • 353077 03/24/11 15:26 a-onl Start_Run_3025,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Production Run on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353076 03/24/11 15:22 Holmstrom Target Move to He
  • 353075 03/24/11 15:20 Solvignon HRSs ready for data taking
  • 353074 03/24/11 15:15 A. Camsonne Don't forget to update the database in analyzer for Gas Cerenkov
  • 353073 03/24/11 15:13 Solvignon Spot++
  • 353072 03/24/11 15:13 a-onl End_of_Run_3024
  • 353071 03/24/11 15:12 a-onl Start_Run_3024,Run_type=Spot++,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.30825,rightarm_theta=29.0,rightarm_sieve=out,comment_text=Spot check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353070 03/24/11 15:09 A. Camsonne Right arm momemtum setting
  • 353069 03/24/11 15:04 Holmstrom Move to BeO
  • 353068 03/24/11 14:34 Solvignon Hall A in beam permit
  • 353067 03/24/11 12:27 auto_snap Radiation and Ozone snapshot
  • 353066 03/24/11 12:27 auto_snap Miscellaneous snapshot
  • 353065 03/24/11 12:27 auto_snap Hall A Tools snapshot
  • 353064 03/24/11 12:27 auto_snap Compton Polarimeter snapshot
  • 353063 03/24/11 12:27 auto_snap BigBox Magnet snapshot
  • 353062 03/24/11 12:27 auto_snap Beam Position and Current snapshot
  • 353061 03/24/11 12:11 Solvignon Ramping up the dipoles
  • 353060 03/24/11 12:05 Jack (Techs) LQ2_LQ3 DAC swapped out
  • 353059 03/24/11 11:25 yez LHRS: Cerenkov HVs adjusted
  • 353058 03/24/11 10:43 yez L-HRS: Cerenkov ADC overfload. Need to recude the HVs
  • 353057 03/24/11 08:55 Holmstrom Move Target Empty
  • 353056 03/24/11 08:51 orchen correction.
  • 353055 03/24/11 08:45 a-onl End_of_Run_3022
  • 353054 03/24/11 08:25 a-onl Start_Run_3022,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353053 03/24/11 08:24 a-onl End_of_Run_3021
  • 353052 03/24/11 08:05 igorko & orc BB TOF peak with smaller bining
  • 353051 03/24/11 08:02 adaq Beam Time Accounting
  • 353050 03/24/11 07:59 ipom Owl shift summary
  • 353049 03/24/11 07:11 siqbal target screen shots
  • 353048 03/24/11 07:10 Higinbotham BigBite Events
  • 353047 03/24/11 07:01 a-onl Start_Run_3021,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353046 03/24/11 07:00 a-onl End_of_Run_3020
  • 353045 03/24/11 05:36 a-onl Start_Run_3020,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353044 03/24/11 05:34 a-onl End_of_Run_3019
  • 353043 03/24/11 05:17 siqbal target screen shots
  • 353042 03/24/11 04:51 a-onl Start_Run_3019,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353041 03/24/11 04:50 a-onl End_of_Run_3018
  • 353040 03/24/11 04:45 orchen removed - didn't change the picture
  • 353039 03/24/11 04:42 orchen HAND TOF peak with a Pmiss dirantional cut
  • 353038 03/24/11 04:29 rshneor can't use BB.tr.tof cut, when plotting the same variable.
  • 353037 03/24/11 04:27 auto_snap Radiation and Ozone snapshot
  • 353036 03/24/11 04:27 auto_snap Miscellaneous snapshot
  • 353035 03/24/11 04:27 auto_snap Hall A Tools snapshot
  • 353034 03/24/11 04:27 auto_snap Compton Polarimeter snapshot
  • 353033 03/24/11 04:27 auto_snap BigBox Magnet snapshot
  • 353032 03/24/11 04:27 auto_snap Beam Position and Current snapshot
  • 353031 03/24/11 04:22 orchen the cuts used
  • 353030 03/24/11 04:13 orchen summery of the triple data we have so far.
  • 353029 03/24/11 03:27 a-onl Start_Run_3018,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353028 03/24/11 03:26 a-onl End_of_Run_3017
  • 353027 03/24/11 03:20 siqbal target screen shots
  • 353026 03/24/11 02:05 a-onl Start_Run_3017,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353025 03/24/11 02:04 a-onl End_of_Run_3016
  • 353024 03/24/11 01:56 nmuangma db_BB.mwdc.dat change: T0 recalibrated and drift_time change from (-900,900) to (0,200) corrected to the wide of the time peak
  • 353023 03/24/11 01:55 ipom spot++ for shift checklist
  • 353022 03/24/11 01:28 siqbal target screen shots
  • 353021 03/24/11 00:43 a-onl Start_Run_3016,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353020 03/24/11 00:43 a-onl End_of_Run_3015
  • 353019 03/24/11 00:10 adaq Beam Time Accounting
  • 353018 03/24/11 00:09 Ole Hansen Swing Shift Summary
  • 353017 03/24/11 00:08 Ole Hansen Charge summing script
  • Wednesday

  • 353016 03/23/11 23:20 H. Albataine Target snapshot
  • 353015 03/23/11 23:19 a-onl Start_Run_3015,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353014 03/23/11 23:18 a-onl End_of_Run_3014
  • 353013 03/23/11 21:52 a-onl Start_Run_3014,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353012 03/23/11 21:45 a-onl Start_Run_3013,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353011 03/23/11 21:44 a-onl End_of_Run_3012
  • 353010 03/23/11 20:58 V. Sulkosky Re: db_run script updated
  • 353009 03/23/11 20:27 auto_snap Radiation and Ozone snapshot
  • 353008 03/23/11 20:27 auto_snap Miscellaneous snapshot
  • 353007 03/23/11 20:27 auto_snap Hall A Tools snapshot
  • 353006 03/23/11 20:27 auto_snap Compton Polarimeter snapshot
  • 353005 03/23/11 20:27 auto_snap BigBox Magnet snapshot
  • 353004 03/23/11 20:27 auto_snap Beam Position and Current snapshot
  • 353003 03/23/11 20:23 a-onl Start_Run_3012,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353002 03/23/11 20:22 a-onl End_of_Run_3011
  • 353001 03/23/11 18:57 a-onl Start_Run_3011,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 353000 03/23/11 18:57 igorko db_run
  • 352999 03/23/11 18:51 Ole Hansen Right Q2 back
  • 352998 03/23/11 18:47 yez L-HRS: Some issues with the PID of Cerenkov and PRL1&2
  • 352997 03/23/11 18:43 Ole Hansen Right Q2 low cryo levels
  • 352996 03/23/11 18:39 a-onl End_of_Run_3010
  • 352995 03/23/11 18:39 Ole Hansen Right Q2 trip
  • 352994 03/23/11 17:24 a-onl Start_Run_3010,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352993 03/23/11 17:23 a-onl End_of_Run_3009
  • 352992 03/23/11 16:47 a-onl Start_Run_3009,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352991 03/23/11 16:31 H. Albataine Target snapshot
  • 352990 03/23/11 16:02 adaq Beam Time Accounting
  • 352989 03/23/11 16:00 Solvignon Day Shift Summary
  • 352988 03/23/11 15:51 a-onl End_of_Run_3008
  • 352987 03/23/11 15:11 igorko small correction
  • 352986 03/23/11 14:43 a-onl Start_Run_3008,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352985 03/23/11 14:42 a-onl End_of_Run_3007
  • 352984 03/23/11 14:36 igorko BB triple coincidence
  • 352983 03/23/11 14:09 Hanretty Target screens
  • 352982 03/23/11 13:13 a-onl Start_Run_3007,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352981 03/23/11 13:13 a-onl End_of_Run_3006
  • 352980 03/23/11 12:27 auto_snap Radiation and Ozone snapshot
  • 352979 03/23/11 12:26 auto_snap Miscellaneous snapshot
  • 352978 03/23/11 12:26 auto_snap Hall A Tools snapshot
  • 352977 03/23/11 12:26 auto_snap Compton Polarimeter snapshot
  • 352976 03/23/11 12:26 auto_snap BigBox Magnet snapshot
  • 352975 03/23/11 12:26 auto_snap Beam Position and Current snapshot
  • 352974 03/23/11 11:50 a-onl Start_Run_3006,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352973 03/23/11 11:49 a-onl End_of_Run_3005
  • 352972 03/23/11 11:31 Hanretty Target screens
  • 352971 03/23/11 11:09 Solvignon Spot++ for shift checklist
  • 352970 03/23/11 11:09 Solvignon Hall A tool for shift checklist
  • 352969 03/23/11 10:27 a-onl Start_Run_3005,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352968 03/23/11 10:26 a-onl End_of_Run_3004
  • 352967 03/23/11 09:06 posik run 3003 junk CODA restart
  • 352966 03/23/11 09:05 a-onl Start_Run_3004,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352965 03/23/11 09:00 Hanretty Target screens
  • 352964 03/23/11 09:00 a-onl Start_Run_3003,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352963 03/23/11 08:59 a-onl End_of_Run_3002
  • 352962 03/23/11 08:02 adaq Beam Time Accounting
  • 352961 03/23/11 07:57 ipom Owl shift summary
  • 352960 03/23/11 07:38 a-onl Start_Run_3002,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352959 03/23/11 07:37 a-onl End_of_Run_3001
  • 352958 03/23/11 07:24 siqbal target screen shots
  • 352957 03/23/11 06:15 a-onl Start_Run_3001,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352956 03/23/11 06:11 a-onl End_of_Run_3000
  • 352955 03/23/11 04:49 a-onl Start_Run_3000,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell, back to 4uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352954 03/23/11 04:47 a-onl End_of_Run_2999
  • 352953 03/23/11 04:26 auto_snap Radiation and Ozone snapshot
  • 352952 03/23/11 04:26 auto_snap Miscellaneous snapshot
  • 352951 03/23/11 04:26 auto_snap Hall A Tools snapshot
  • 352950 03/23/11 04:26 auto_snap Compton Polarimeter snapshot
  • 352949 03/23/11 04:26 auto_snap BigBox Magnet snapshot
  • 352948 03/23/11 04:26 auto_snap Beam Position and Current snapshot
  • 352947 03/23/11 04:08 siqbal target screen shot
  • 352946 03/23/11 03:50 ipom Test run with 2uA
  • 352945 03/23/11 03:47 a-onl Start_Run_2999,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=2,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell, test run on 2uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352944 03/23/11 03:45 a-onl End_of_Run_2998
  • 352943 03/23/11 02:23 a-onl Start_Run_2998,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352942 03/23/11 02:23 a-onl End_of_Run_2997
  • 352941 03/23/11 01:37 siqbal target screen shots
  • 352940 03/23/11 01:01 a-onl Start_Run_2997,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352939 03/23/11 01:00 a-onl End_of_Run_2996
  • 352938 03/23/11 00:08 adaq Beam Time Accounting
  • 352937 03/23/11 00:07 Boeglin shift summary
  • Tuesday

  • 352936 03/22/11 23:37 a-onl Start_Run_2996,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352935 03/22/11 23:36 a-onl End_of_Run_2995
  • 352934 03/22/11 23:24 H. Albataine Target snapshot
  • 352933 03/22/11 22:15 a-onl Start_Run_2995,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352932 03/22/11 22:14 a-onl End_of_Run_2994
  • 352931 03/22/11 21:54 igorko C / LD2 ratio
  • 352930 03/22/11 21:44 igorko & orc HAND lib change
  • 352929 03/22/11 21:12 Boeglin Hall A Tools Screen
  • 352928 03/22/11 21:10 adaq spot++ 2994
  • 352927 03/22/11 21:01 israely Target screenshot
  • 352926 03/22/11 21:00 israely neutron detector screenshots
  • 352925 03/22/11 20:59 israely Bigbite scintilator & scaler screenshot
  • 352924 03/22/11 20:57 israely Coda screenshot
  • 352923 03/22/11 20:56 israely We are back and taking data.
  • 352922 03/22/11 20:53 a-onl Start_Run_2994,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352921 03/22/11 20:37 a-onl End_of_Run_2993
  • 352920 03/22/11 20:36 a-onl Start_Run_2993,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352919 03/22/11 20:26 auto_snap Radiation and Ozone snapshot
  • 352918 03/22/11 20:26 auto_snap Miscellaneous snapshot
  • 352917 03/22/11 20:26 auto_snap Hall A Tools snapshot
  • 352916 03/22/11 20:26 auto_snap Compton Polarimeter snapshot
  • 352915 03/22/11 20:26 auto_snap BigBox Magnet snapshot
  • 352914 03/22/11 20:25 auto_snap Beam Position and Current snapshot
  • 352913 03/22/11 20:11 a-onl End_of_Run_2992
  • 352912 03/22/11 20:11 a-onl Start_Run_2992,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352911 03/22/11 20:09 a-onl End_of_Run_2991
  • 352910 03/22/11 20:09 israely problem stills accures after restarting coda
  • 352909 03/22/11 20:07 a-onl Start_Run_2991,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352908 03/22/11 19:59 israely Bigbite is rebooting
  • 352907 03/22/11 19:57 a-onl Start_Run_2990,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352906 03/22/11 19:57 a-onl End_of_Run_2989
  • 352905 03/22/11 18:32 a-onl Start_Run_2989,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352904 03/22/11 18:31 a-onl End_of_Run_2988
  • 352903 03/22/11 17:32 israely Target screenshot
  • 352902 03/22/11 17:10 a-onl Start_Run_2988,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352901 03/22/11 17:09 a-onl End_of_Run_2987
  • 352900 03/22/11 16:08 adaq Beam Time Accounting
  • 352899 03/22/11 15:58 israely Target screenshots
  • 352898 03/22/11 15:57 israely neutron detector screenshots
  • 352897 03/22/11 15:56 israely Bigbite and Scalers screenshot
  • 352896 03/22/11 15:48 a-onl Start_Run_2987,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352895 03/22/11 15:48 X. Zheng Target snap shots
  • 352894 03/22/11 15:46 a-onl End_of_Run_2986
  • 352893 03/22/11 15:46 Norum Run summary
  • 352892 03/22/11 15:44 israely Coda screenshot
  • 352891 03/22/11 15:18 orchen mixup in the run numbers
  • 352890 03/22/11 14:57 jpchen TO: In emergency situation when target warmed up, you need to reduce fan speed.
  • 352889 03/22/11 14:43 jpchen Fan speed are at 48 Hz for both loops, do not change them
  • 352888 03/22/11 14:25 a-onl Start_Run_2986,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352887 03/22/11 14:25 a-onl End_of_Run_2985
  • 352886 03/22/11 13:19 X. Zheng Target snap shots
  • 352885 03/22/11 13:10 posik 4He 4uA rates
  • 352884 03/22/11 13:04 a-onl Start_Run_2985,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm cell,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352883 03/22/11 13:04 Norum Production running
  • 352882 03/22/11 13:03 a-onl End_of_Run_2984
  • 352881 03/22/11 12:58 a-onl Start_Run_2984,Run_type=Spot++,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Slanted Carbon.,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352880 03/22/11 12:56 X. Zheng Target moved to He4 loop 1 20cm
  • 352879 03/22/11 12:46 a-onl End_of_Run_2983
  • 352878 03/22/11 12:39 posik spot++ 2x2
  • 352877 03/22/11 12:35 a-onl Start_Run_2983,Run_type=Spot++,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Slanted Carbon.,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352876 03/22/11 12:32 a-onl End_of_Run_2982
  • 352875 03/22/11 12:25 auto_snap Radiation and Ozone snapshot
  • 352874 03/22/11 12:25 auto_snap Miscellaneous snapshot
  • 352873 03/22/11 12:25 auto_snap Hall A Tools snapshot
  • 352872 03/22/11 12:25 auto_snap Compton Polarimeter snapshot
  • 352871 03/22/11 12:25 auto_snap BigBox Magnet snapshot
  • 352870 03/22/11 12:25 auto_snap Beam Position and Current snapshot
  • 352869 03/22/11 12:08 igorko additional plot
  • 352868 03/22/11 11:52 igorko Carbon and deutron target
  • 352867 03/22/11 11:48 jpchen TO needs to be careful when cahnge fan speed
  • 352866 03/22/11 11:40 jpchen Helium filled to full by Dave. Target ready for beam.
  • 352865 03/22/11 09:45 a-onl Start_Run_2982,Run_type=Cosmics,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Slanted Carbon.,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352864 03/22/11 09:27 posik 2978-2979 LHRS VDC Plane/RHRS Cer sum
  • 352863 03/22/11 09:20 X. Zheng Target moved to empty
  • 352862 03/22/11 09:17 posik run 2978 ROC1 vs ROC2 histo not found
  • 352861 03/22/11 09:17 Norum Beam stopped
  • 352860 03/22/11 09:15 a-onl End_of_Run_2981
  • 352859 03/22/11 08:59 posik run 2979 ROC10 restart
  • 352858 03/22/11 08:59 a-onl Start_Run_2981,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Slanted Carbon.,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352857 03/22/11 08:58 Norum Tools
  • 352856 03/22/11 08:57 Norum Methane alarm
  • 352855 03/22/11 08:42 a-onl End_of_Run_2979
  • 352854 03/22/11 08:26 posik run 2978 coda got hung up
  • 352853 03/22/11 08:19 a-onl Start_Run_2979,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Slanted Carbon.,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352852 03/22/11 08:04 adaq Beam Time Accounting
  • 352851 03/22/11 07:57 ipom Owl shift summary
  • 352849 03/22/11 07:47 ipom BB HV power of not from GUI warning
  • 352848 03/22/11 07:47 a-onl Start_Run_2978,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Slanted Carbon.,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352847 03/22/11 07:46 siqbal target screens
  • 352846 03/22/11 07:45 a-onl End_of_Run_2977
  • 352845 03/22/11 06:55 a-onl Start_Run_2977,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Slanted Carbon. ,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352844 03/22/11 06:49 ipom 2 hours of LD2 done,going back to slanted C
  • 352843 03/22/11 06:47 a-onl End_of_Run_2976
  • 352842 03/22/11 05:54 a-onl Start_Run_2976,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=20cm LD target, ,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352841 03/22/11 05:53 a-onl End_of_Run_2975
  • 352840 03/22/11 05:49 orchen 750MeV/c protons energy deposit
  • 352839 03/22/11 05:45 siqbal target screen shots
  • 352838 03/22/11 05:23 orchen BB 12C(e,e'p_recoil) with a -0.1V treshold
  • 352837 03/22/11 05:09 orchen HV values for BB and HAND are set fine
  • 352836 03/22/11 05:07 orchen another small modification to the treshold
  • 352835 03/22/11 04:56 orchen Final BB tresholds
  • 352834 03/22/11 04:55 ipom BB HV on chamber #1 went off.
  • 352833 03/22/11 04:55 a-onl Start_Run_2975,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=20cm LD target, prescale factor is adjusted,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352832 03/22/11 04:53 a-onl End_of_Run_2974
  • 352831 03/22/11 04:33 orchen BB treshold change and "bad" runs
  • 352830 03/22/11 04:25 a-onl Start_Run_2974,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=20cm LD target, prescale factor is adjusted,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352829 03/22/11 04:25 auto_snap Radiation and Ozone snapshot
  • 352828 03/22/11 04:25 auto_snap Miscellaneous snapshot
  • 352827 03/22/11 04:25 auto_snap Hall A Tools snapshot
  • 352826 03/22/11 04:25 auto_snap Compton Polarimeter snapshot
  • 352825 03/22/11 04:25 auto_snap BigBox Magnet snapshot
  • 352824 03/22/11 04:25 auto_snap Beam Position and Current snapshot
  • 352823 03/22/11 04:24 a-onl End_of_Run_2973
  • 352822 03/22/11 04:24 ipom BB HV
  • 352821 03/22/11 04:16 a-onl Start_Run_2973,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=20cm LD target, prescale factor is adjusted,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352820 03/22/11 04:16 ipom BB HV are off.
  • 352819 03/22/11 04:15 a-onl End_of_Run_2972
  • 352818 03/22/11 04:08 ipom Rates on LD2
  • 352817 03/22/11 04:06 a-onl Start_Run_2972,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=20cm LD target, prescale factor is adjusted,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352816 03/22/11 04:05 a-onl End_of_Run_2971
  • 352815 03/22/11 03:56 a-onl Start_Run_2971,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=20cm LD target, prescale factor is adjusted,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352814 03/22/11 03:55 a-onl End_of_Run_2970
  • 352813 03/22/11 03:54 a-onl Start_Run_2970,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=20cm LD target ,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352812 03/22/11 03:44 ipom Done 2 hrs slanted C target. Going to LD2
  • 352811 03/22/11 03:42 a-onl End_of_Run_2969
  • 352810 03/22/11 03:15 a-onl Start_Run_2969,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Slanted Carbon after DAQ fixed,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352809 03/22/11 03:14 a-onl End_of_Run_2968
  • 352808 03/22/11 02:14 orchen New BB tresholds
  • 352807 03/22/11 02:12 a-onl Start_Run_2968,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Slanted Carbon after DAQ fixed,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352806 03/22/11 02:10 a-onl End_of_Run_2967
  • 352805 03/22/11 01:40 ipom rates look ok now
  • 352804 03/22/11 01:36 a-onl Start_Run_2967,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Slanted Carbon after DAQ fixed,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352803 03/22/11 01:23 siqbal target screens
  • 352802 03/22/11 01:18 a-onl End_of_Run_2966
  • 352801 03/22/11 01:17 A. Camsonne Bad connection at output of NIM to ECL level translator
  • 352800 03/22/11 01:15 a-onl Start_Run_2966,Run_type=Other,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Test DAQ,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352799 03/22/11 01:09 a-onl End_of_Run_2965
  • 352798 03/22/11 01:04 a-onl Start_Run_2965,Run_type=Other,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Test DAQ T3 only,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352797 03/22/11 01:03 a-onl End_of_Run_2964
  • 352796 03/22/11 01:03 a-onl Start_Run_2964,Run_type=Other,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Test DAQ T3,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352795 03/22/11 00:51 a-onl Start_Run_2963,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Slanted Carbon - after BBTS1 power cycling,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352794 03/22/11 00:46 A. Camsonne Problem with T3
  • 352793 03/22/11 00:45 a-onl End_of_Run_2962
  • 352792 03/22/11 00:44 a-onl Start_Run_2962,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Slanted Carbon - Only T1 ps1=1 ps3=10 all others are 0,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352791 03/22/11 00:43 a-onl End_of_Run_2961
  • 352790 03/22/11 00:43 a-onl Start_Run_2961,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Slanted Carbon - Only T1 ps1=1 all others are 0,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352789 03/22/11 00:42 a-onl End_of_Run_2960
  • 352788 03/22/11 00:39 a-onl Start_Run_2960,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Slanted Carbon - Only T3 ps3=1,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352787 03/22/11 00:38 a-onl End_of_Run_2959
  • 352786 03/22/11 00:38 a-onl Start_Run_2959,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Slanted Carbon - Test all triggers ps to 0,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352785 03/22/11 00:37 a-onl End_of_Run_2958
  • 352784 03/22/11 00:36 a-onl Start_Run_2958,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Slanted Carbon - Test ps1=0,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352783 03/22/11 00:35 a-onl End_of_Run_2957
  • 352782 03/22/11 00:33 a-onl Start_Run_2957,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Slanted Carbon - After CODA restart,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352781 03/22/11 00:30 a-onl End_of_Run_2956
  • 352780 03/22/11 00:29 a-onl Start_Run_2956,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Slanted Carbon - After BBTS1 reboot,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352779 03/22/11 00:27 A. Camsonne Rebooting bbts1
  • 352778 03/22/11 00:25 a-onl End_of_Run_2955
  • 352777 03/22/11 00:22 a-onl Start_Run_2955,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Slanted Carbon,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352776 03/22/11 00:21 a-onl End_of_Run_2954
  • 352775 03/22/11 00:14 a-onl Start_Run_2954,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Slanted Carbon,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352774 03/22/11 00:11 israely Target screenshot
  • 352773 03/22/11 00:06 adaq Beam Time Accounting
  • Monday

  • 352772 03/21/11 23:55 A. Camsonne Left Q1 magnet quenched
  • 352771 03/21/11 23:53 JinGe fan of the target turned down
  • 352770 03/21/11 23:48 Norum. LHe LHRS Q1 quench
  • 352769 03/21/11 23:48 JinGe target is moved to slanted carbon
  • 352768 03/21/11 23:39 JinGe target is now moved to Carbon
  • 352767 03/21/11 23:27 JinGe target is moved to 4He 20cm
  • 352766 03/21/11 23:18 israely neutron detector screenshots
  • 352765 03/21/11 23:17 a-onl End_of_Run_2953
  • 352764 03/21/11 23:16 a-onl Start_Run_2953,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=He4 - 20cm Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352763 03/21/11 23:14 a-onl End_of_Run_2952
  • 352762 03/21/11 23:12 israely spot++
  • 352761 03/21/11 23:10 Norum Hall A tools
  • 352760 03/21/11 23:09 a-onl Start_Run_2952,Run_type=Spot++,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=spot++,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352759 03/21/11 22:21 JinGe target is moved to BeO
  • 352758 03/21/11 21:32 Norum Beam update
  • 352757 03/21/11 20:51 JinGe beam expected to be back at 930 pm
  • 352756 03/21/11 20:25 auto_snap Radiation and Ozone snapshot
  • 352755 03/21/11 20:25 auto_snap Miscellaneous snapshot
  • 352754 03/21/11 20:25 auto_snap Hall A Tools snapshot
  • 352753 03/21/11 20:25 auto_snap Compton Polarimeter snapshot
  • 352752 03/21/11 20:25 auto_snap BigBox Magnet snapshot
  • 352751 03/21/11 20:25 auto_snap Beam Position and Current snapshot
  • 352750 03/21/11 18:54 orchen beam update
  • 352749 03/21/11 18:49 igorko E plane thresholds
  • 352748 03/21/11 18:48 JinGe call from chief crew: beam will be back in 2.5 hours if things go smoothly
  • 352747 03/21/11 18:20 orchen BB tresholds (TOO LOW)
  • 352746 03/21/11 17:58 JinGe HallA tools alarm
  • 352745 03/21/11 17:35 JinGe haCryo_supply
  • 352744 03/21/11 17:24 JinGe target snapshots
  • 352743 03/21/11 17:23 JinGe haCryo_supply below low limit sometimes
  • 352742 03/21/11 16:57 A. Camsonne Target filled about half way for tonight - Pay attention to fan on Loop 1 tripping off
  • 352741 03/21/11 16:24 A. Camsonne Target will be refilled soon
  • 352740 03/21/11 16:17 A. Camsonne He4 target vented
  • 352739 03/21/11 16:11 J. Zhang day shift summary
  • 352738 03/21/11 16:10 J. Zhang Helium Target Vented
  • 352737 03/21/11 15:58 a-onl Start_Run_2951,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=cosmics PS1 changed to 1,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352736 03/21/11 15:57 a-onl End_of_Run_2951
  • 352735 03/21/11 15:56 A. Camsonne MCC update
  • 352734 03/21/11 15:51 Jack (Techs) Left Dipole NMR cabling
  • 352733 03/21/11 15:46 A. Camsonne Electronics board put in Hall A Radcon staging cabinet
  • 352732 03/21/11 15:43 J. Zhang Cycle the HRS magnets
  • 352731 03/21/11 15:40 Jack (Techs) LQ1 power supply water hoses
  • 352730 03/21/11 15:37 Jack (Techs) Donut transformers on the magnet power supply inputs
  • 352729 03/21/11 15:28 J. Zhang target snapshot
  • 352728 03/21/11 14:35 a-onl End_of_Run_2950
  • 352727 03/21/11 14:11 ipom Batch replay script
  • 352726 03/21/11 12:24 auto_snap Radiation and Ozone snapshot
  • 352725 03/21/11 12:24 auto_snap Miscellaneous snapshot
  • 352724 03/21/11 12:24 auto_snap Hall A Tools snapshot
  • 352723 03/21/11 12:24 auto_snap Compton Polarimeter snapshot
  • 352722 03/21/11 12:24 auto_snap BigBox Magnet snapshot
  • 352721 03/21/11 12:24 auto_snap Beam Position and Current snapshot
  • 352720 03/21/11 11:30 Fansler Cryo
  • 352719 03/21/11 10:54 a-onl Start_Run_2950,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=cosmics PS1 changed to 1,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352718 03/21/11 09:14 Fansler Cryo
  • 352717 03/21/11 09:13 J. Zhang Loop2 temp alarm
  • 352716 03/21/11 09:06 J. Zhang Loop2 temperature alram
  • 352715 03/21/11 08:19 Jixie Zhang move to Empty Target
  • 352714 03/21/11 08:17 A. Camsonne Accelerator update : no beam for now ARC power supply still needs to be fixed
  • 352713 03/21/11 07:28 siqbal target screens
  • 352712 03/21/11 04:31 siqbal target screens
  • 352711 03/21/11 04:24 auto_snap Radiation and Ozone snapshot
  • 352710 03/21/11 04:24 auto_snap Miscellaneous snapshot
  • 352709 03/21/11 04:24 auto_snap Hall A Tools snapshot
  • 352708 03/21/11 04:24 auto_snap Compton Polarimeter snapshot
  • 352707 03/21/11 04:24 auto_snap BigBox Magnet snapshot
  • 352706 03/21/11 04:24 auto_snap Beam Position and Current snapshot
  • 352705 03/21/11 02:29 siqbal target screens
  • 352704 03/21/11 00:16 adaq Beam Time Accounting
  • 352703 03/21/11 00:06 a-onl End_of_Run_2949
  • 352702 03/21/11 00:02 adaq Beam Time Accounting
  • Sunday

  • 352701 03/20/11 23:51 M Defurne screenshots
  • 352700 03/20/11 23:05 LeRose Shift Summary
  • 352699 03/20/11 23:04 LeRose No Beam Owl shift on standby
  • 352698 03/20/11 20:24 auto_snap Radiation and Ozone snapshot
  • 352697 03/20/11 20:24 auto_snap Miscellaneous snapshot
  • 352696 03/20/11 20:24 auto_snap Hall A Tools snapshot
  • 352695 03/20/11 20:24 auto_snap Compton Polarimeter snapshot
  • 352694 03/20/11 20:24 auto_snap BigBox Magnet snapshot
  • 352693 03/20/11 20:24 auto_snap Beam Position and Current snapshot
  • 352692 03/20/11 18:14 M Defurne screenshots
  • 352691 03/20/11 18:11 LeRose Left NMR
  • 352690 03/20/11 17:37 LeRose Left NMR O-scope pictures
  • 352689 03/20/11 17:22 LeRose Left Dipole NMR Lock
  • 352688 03/20/11 16:54 igorko BB triple coincidence
  • 352687 03/20/11 16:15 Jack (Techs) 480VAC phase issues for the magnet power supplies
  • 352686 03/20/11 16:01 adaq Beam Time Accounting
  • 352685 03/20/11 16:00 Jack (Techs) Left Dipole NMR lock
  • 352684 03/20/11 15:58 eschulte Day Shift Log
  • 352683 03/20/11 15:22 a-onl Start_Run_2949,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=cosmics PS1 changed to 1,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352682 03/20/11 15:21 a-onl End_of_Run_2948
  • 352681 03/20/11 15:09 eschulte MCC Update: No ETA
  • 352680 03/20/11 12:32 igorko target
  • 352679 03/20/11 12:23 auto_snap Radiation and Ozone snapshot
  • 352678 03/20/11 12:23 auto_snap Miscellaneous snapshot
  • 352677 03/20/11 12:23 auto_snap Hall A Tools snapshot
  • 352676 03/20/11 12:23 auto_snap Compton Polarimeter snapshot
  • 352675 03/20/11 12:23 auto_snap BigBox Magnet snapshot
  • 352674 03/20/11 12:23 auto_snap Beam Position and Current snapshot
  • 352673 03/20/11 10:43 eschulte HRS Magnets (NMR issues)
  • 352672 03/20/11 10:39 igorko target
  • 352671 03/20/11 09:11 eschulte Hall in Controlled Access
  • 352670 03/20/11 09:08 A. Camsonne Beam status update - No beam before swing shift - Going to controlled access
  • 352669 03/20/11 08:41 xqian Target Screen
  • 352668 03/20/11 08:39 eip s/b
  • 352667 03/20/11 08:19 xqian some thinkings about S/B
  • 352666 03/20/11 08:19 xqian some thinkings about S/B
  • 352665 03/20/11 08:18 orchen Priliminary 4He(e,e'pn) analysis in X>1.2 and Q2=2
  • 352664 03/20/11 08:04 adaq Beam Time Accounting
  • 352663 03/20/11 07:58 ipom Owl shift summary
  • 352662 03/20/11 07:56 ipom No beam until swing shift
  • 352661 03/20/11 06:51 H. Albataine Target snapshot
  • 352660 03/20/11 05:56 ipom No beam
  • 352659 03/20/11 04:43 a-onl Start_Run_2948,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=cosmics,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352658 03/20/11 04:28 orchen Low rate run summery and conclusions
  • 352657 03/20/11 04:23 auto_snap Radiation and Ozone snapshot
  • 352656 03/20/11 04:23 auto_snap Miscellaneous snapshot
  • 352655 03/20/11 04:23 auto_snap Hall A Tools snapshot
  • 352654 03/20/11 04:23 auto_snap Compton Polarimeter snapshot
  • 352653 03/20/11 04:23 auto_snap BigBox Magnet snapshot
  • 352652 03/20/11 04:23 auto_snap Beam Position and Current snapshot
  • 352651 03/20/11 02:18 orchen HAND (e,e'N_recoil) peak for run 2942 (2.5uA)
  • 352650 03/20/11 02:15 nmuangma T0 setting
  • 352649 03/20/11 02:08 ipom No beam
  • 352648 03/20/11 02:06 orchen BB (e,e'p_recoil) from run 2942 (2.5uA)
  • 352647 03/20/11 00:37 H. Albataine Target snapshot
  • 352646 03/20/11 00:21 ipom MCC says that they got a fire alarm and some sprinkels went off in a power supply room
  • 352645 03/20/11 00:05 adaq Beam Time Accounting
  • Saturday

  • 352644 03/19/11 23:48 mhuang Swing Shift Summary
  • 352643 03/19/11 23:45 eip compare to low current
  • 352642 03/19/11 23:42 eip please do the same to the 2.5 micro run
  • 352641 03/19/11 23:38 brads roc8 ped suppression off -- is that intentional?
  • 352640 03/19/11 22:40 igorko HAND TOF
  • 352639 03/19/11 22:19 a-onl End_of_Run_2947
  • 352638 03/19/11 21:52 pzhu run 2947's neutron
  • 352637 03/19/11 21:51 pzhu run 2947's scaler,bigbite
  • 352636 03/19/11 21:51 pzhu run 2947's RC1,SawDTmon
  • 352635 03/19/11 21:50 a-onl Start_Run_2947,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352634 03/19/11 21:39 a-onl End_of_Run_2946
  • 352633 03/19/11 21:31 mhuang Hall A tools screen
  • 352632 03/19/11 21:30 adaq spot++ for 2946
  • 352631 03/19/11 21:22 pzhu run 2946's neutron
  • 352630 03/19/11 21:21 pzhu run 2946's bigbite and scaler
  • 352629 03/19/11 21:21 pzhu run 2946's RC1,SawDTmon
  • 352628 03/19/11 21:19 a-onl Start_Run_2946,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352627 03/19/11 20:23 auto_snap Radiation and Ozone snapshot
  • 352626 03/19/11 20:23 auto_snap Miscellaneous snapshot
  • 352625 03/19/11 20:23 auto_snap Hall A Tools snapshot
  • 352624 03/19/11 20:23 auto_snap Compton Polarimeter snapshot
  • 352623 03/19/11 20:23 auto_snap BigBox Magnet snapshot
  • 352622 03/19/11 20:23 auto_snap Beam Position and Current snapshot
  • 352621 03/19/11 20:19 pzhu no beam
  • 352620 03/19/11 18:58 a-onl End_of_Run_2945
  • 352619 03/19/11 18:06 pzhu run 2945's neutron
  • 352618 03/19/11 18:05 pzhu run 2945's scaler and bigbite
  • 352617 03/19/11 18:04 pzhu run 2945's RC1 and SawDTmon
  • 352616 03/19/11 18:04 a-onl Start_Run_2945,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352615 03/19/11 18:03 a-onl End_of_Run_2944
  • 352614 03/19/11 17:44 A. Camsonne Retimnf
  • 352613 03/19/11 17:11 orchen Bigbite (e,e'p_recoil) status
  • 352612 03/19/11 16:50 pzhu run 2944's neutron
  • 352611 03/19/11 16:49 pzhu run 2944's bigbite,scaler
  • 352610 03/19/11 16:48 pzhu run 2944's RC1 and SawDTmon
  • 352609 03/19/11 16:48 a-onl Start_Run_2944,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352608 03/19/11 16:44 A. Camsonne Beam down - stuck in the injector for now
  • 352607 03/19/11 16:41 a-onl End_of_Run_2943
  • 352606 03/19/11 16:23 israely Target screens
  • 352605 03/19/11 16:03 pzhu run 2943's neutron
  • 352604 03/19/11 16:02 pzhu run 2943's scaler,bigbite
  • 352603 03/19/11 16:02 adaq Beam Time Accounting
  • 352602 03/19/11 16:01 pzhu run 2943's RC1,SawDTmon
  • 352601 03/19/11 16:01 eschulte Day Shift Log
  • 352600 03/19/11 16:00 a-onl Start_Run_2943,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352599 03/19/11 16:00 eschulte Changed Beam Current to 5uA
  • 352598 03/19/11 16:00 eschulte Day Shift Run LIst
  • 352597 03/19/11 15:59 a-onl End_of_Run_2942
  • 352596 03/19/11 15:48 eschulte HRSL NMR Lock
  • 352595 03/19/11 15:24 eschulte Beam's back!
  • 352594 03/19/11 15:19 eschulte Extended Beam Down Run 2942
  • 352593 03/19/11 14:39 eschulte Run 2942 BigBite Wire Chambers
  • 352592 03/19/11 13:54 pzhu run 2942's RC1 and SawDTmon
  • 352591 03/19/11 13:53 pzhu run 2942's neutron
  • 352590 03/19/11 13:53 pzhu run 2942's scaler and bigbite
  • 352589 03/19/11 13:51 a-onl Start_Run_2942,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=2.5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352588 03/19/11 13:51 eschulte Reducing current to 2.5 uA
  • 352587 03/19/11 13:50 igorko & orc S/N check
  • 352586 03/19/11 13:49 a-onl End_of_Run_2941
  • 352585 03/19/11 13:39 Higinbotham T3 Rate During 12C(e,e'pN) Experiment (March 2005)
  • 352584 03/19/11 13:38 pzhu 2941's neutron
  • 352583 03/19/11 13:37 pzhu 2941's scaler and bigbite
  • 352582 03/19/11 13:36 pzhu run 2941's RC1,SawDTmon
  • 352581 03/19/11 13:35 a-onl Start_Run_2941,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352580 03/19/11 13:34 a-onl End_of_Run_2940
  • 352579 03/19/11 13:20 eschulte Extended Beam Down Run 2940
  • 352578 03/19/11 13:07 pzhu run 2940's neutron
  • 352577 03/19/11 13:06 pzhu run 2940's bigbite, scaler
  • 352576 03/19/11 13:05 pzhu run 2940's RC1,SawDTmon
  • 352575 03/19/11 12:30 a-onl Start_Run_2940,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352574 03/19/11 12:29 a-onl End_of_Run_2939
  • 352573 03/19/11 12:22 auto_snap Radiation and Ozone snapshot
  • 352572 03/19/11 12:22 auto_snap Miscellaneous snapshot
  • 352571 03/19/11 12:22 auto_snap Hall A Tools snapshot
  • 352570 03/19/11 12:22 auto_snap Compton Polarimeter snapshot
  • 352569 03/19/11 12:22 auto_snap BigBox Magnet snapshot
  • 352568 03/19/11 12:22 auto_snap Beam Position and Current snapshot
  • 352567 03/19/11 12:00 orchen BB (e,e'p_recoil) TOF peak with some basic PID cuts
  • 352566 03/19/11 11:30 adaq CODA and Deadtime monitor
  • 352565 03/19/11 11:25 chench screenshot for neutron detectors sigle rate between run2939
  • 352564 03/19/11 11:24 chench screen shot for BB scintillator single rates and trigger rates
  • 352563 03/19/11 11:21 chench Spot_R after beam back run2938
  • 352562 03/19/11 11:20 yawei Target Screens.
  • 352561 03/19/11 11:20 a-onl Start_Run_2939,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352560 03/19/11 11:18 eschulte Beam Back
  • 352559 03/19/11 11:18 a-onl End_of_Run_2938
  • 352558 03/19/11 11:16 chench rates screenshot after beam back
  • 352557 03/19/11 11:16 a-onl Start_Run_2938,Run_type=Spot++,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=short beam recovery,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352556 03/19/11 10:59 a-onl Start_Run_2937,Run_type=Other,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=test DAQ SRC additional delay on bigbite- 20 ns readded on T3 before TS,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352555 03/19/11 10:50 A. Camsonne Network switch at BigBite weldment was power cycled
  • 352554 03/19/11 10:15 Higinbotham Trying To Check Pacticle Type Of Peak
  • 352553 03/19/11 09:52 A. Camsonne Network down on BigBite weldment - most likely switch needs to be power cycled or is dead
  • 352552 03/19/11 09:52 eschulte BigBite: Lost Communication
  • 352551 03/19/11 09:24 a-onl Start_Run_2936,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=production SRC additional delay on bigbite,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352550 03/19/11 09:24 eschulte Half Wave Plate
  • 352549 03/19/11 09:21 a-onl End_of_Run_2935
  • 352548 03/19/11 09:15 A. Camsonne Yes you can delete whatever is on the /adaql10/work1
  • 352547 03/19/11 09:08 adaq coda and rates
  • 352546 03/19/11 08:39 orchen Is there data from ole experiments on adaql10
  • 352545 03/19/11 08:31 adaq online GUI for run2935
  • 352544 03/19/11 08:03 adaq Beam Time Accounting
  • 352543 03/19/11 08:01 chench Screenshot of Neutron Deterctor single rates
  • 352542 03/19/11 08:00 chench Screenshot of BB scintillator single rates
  • 352541 03/19/11 07:57 a-onl Start_Run_2935,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=production SRC additional delay on bigbite,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352540 03/19/11 07:56 a-onl End_of_Run_2934
  • 352539 03/19/11 07:45 israely Owl shift Summary
  • 352538 03/19/11 07:14 israely Target screenshot
  • 352537 03/19/11 07:13 israely Hall A screenshot
  • 352536 03/19/11 07:03 siqbal bigbite and scalars
  • 352535 03/19/11 07:01 siqbal neutron detector screen shots
  • 352534 03/19/11 06:44 a-onl Start_Run_2934,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=production SRC additional delay on bigbite,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352533 03/19/11 06:43 a-onl End_of_Run_2933
  • 352532 03/19/11 05:22 a-onl Start_Run_2933,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=production SRC additional delay on bigbite,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352531 03/19/11 05:22 a-onl End_of_Run_2932
  • 352530 03/19/11 05:00 igorko DISK SPACE
  • 352529 03/19/11 04:50 israely Coda screenshot
  • 352528 03/19/11 04:41 igorko BB double coin
  • 352527 03/19/11 04:22 auto_snap Radiation and Ozone snapshot
  • 352526 03/19/11 04:22 auto_snap Miscellaneous snapshot
  • 352525 03/19/11 04:22 auto_snap Hall A Tools snapshot
  • 352524 03/19/11 04:22 auto_snap Compton Polarimeter snapshot
  • 352523 03/19/11 04:22 auto_snap BigBox Magnet snapshot
  • 352522 03/19/11 04:22 auto_snap Beam Position and Current snapshot
  • 352521 03/19/11 04:21 igorko T6 and T7 info
  • 352520 03/19/11 04:11 israely Beam Compensation screenshot
  • 352519 03/19/11 04:10 a-onl Start_Run_2932,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=production SRC additional delay on bigbite,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352518 03/19/11 04:09 a-onl End_of_Run_2931
  • 352517 03/19/11 03:38 israely Lecroy HV screenshot
  • 352516 03/19/11 03:36 israely Hall A screenshots
  • 352515 03/19/11 03:34 israely FSD screenshot
  • 352514 03/19/11 03:01 israely Hall A screenshot
  • 352513 03/19/11 02:52 A. Camsonne Rates on He4 at 1 uA Ok
  • 352512 03/19/11 02:52 a-onl Start_Run_2931,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=production SRC additional delay on bigbite,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352511 03/19/11 02:50 siqbal scalars
  • 352510 03/19/11 02:47 siqbal scalars
  • 352509 03/19/11 02:47 israely Target screenshot
  • 352508 03/19/11 02:37 A. Camsonne Beam and raster visually checked on BeO
  • 352507 03/19/11 02:36 a-onl End_of_Run_2930
  • 352506 03/19/11 02:35 israely spot++
  • 352505 03/19/11 02:33 a-onl Start_Run_2930,Run_type=Spot++,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=spot ++,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352504 03/19/11 02:30 A. Camsonne Cabling completed starting beam recovery
  • 352503 03/19/11 02:24 igorko BB delay cables
  • 352502 03/19/11 01:39 israely Igor and alex are still connecting the cables
  • 352501 03/19/11 01:36 A. Camsonne Delay removed only on T3 input at trigger supervisor
  • 352500 03/19/11 01:34 A. Camsonne Additionnal channels in TDCs
  • 352499 03/19/11 00:53 a-onl Start_Run_2929,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=test run - check new signals in 1190, F1 and fastbus,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352498 03/19/11 00:37 israely The beam is changed to unpolarized beam
  • 352497 03/19/11 00:20 a-onl End_of_Run_2928
  • 352496 03/19/11 00:02 adaq Beam Time Accounting
  • 352495 03/19/11 00:01 eschulte Swing Shift Log
  • Friday

  • 352494 03/18/11 23:56 eip 20 ns from t3
  • 352493 03/18/11 23:35 eschulte Swing Shift Run Tables
  • 352492 03/18/11 23:34 orchen BigBite IS NOT IN TIME - working on fixing it.
  • 352491 03/18/11 23:12 Xin Qian SRC Timing Problem
  • 352490 03/18/11 21:53 a-onl Start_Run_2928,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=test run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352489 03/18/11 21:51 eschulte To be clear...
  • 352488 03/18/11 21:51 a-onl End_of_Run_2927
  • 352487 03/18/11 21:50 eschulte Actually a test run (2927)
  • 352486 03/18/11 21:15 a-onl Start_Run_2927,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=test run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352485 03/18/11 21:12 eschulte Hall in Controlled Access
  • 352484 03/18/11 21:12 a-onl End_of_Run_2926
  • 352483 03/18/11 20:36 a-onl End_of_Run_2925
  • 352482 03/18/11 20:22 auto_snap Radiation and Ozone snapshot
  • 352481 03/18/11 20:22 auto_snap Miscellaneous snapshot
  • 352480 03/18/11 20:22 auto_snap Hall A Tools snapshot
  • 352479 03/18/11 20:22 auto_snap Compton Polarimeter snapshot
  • 352478 03/18/11 20:22 auto_snap BigBox Magnet snapshot
  • 352477 03/18/11 20:21 auto_snap Beam Position and Current snapshot
  • 352476 03/18/11 20:13 Punjabi spot++ for run 2925
  • 352475 03/18/11 19:36 a-onl Start_Run_2925,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production Running,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352474 03/18/11 19:35 a-onl End_of_Run_2924
  • 352473 03/18/11 18:35 Jack (Techs) Left Dipole NMR lock
  • 352472 03/18/11 18:27 Jack (Techs) EPICS alarm fields not set
  • 352471 03/18/11 18:22 a-onl Start_Run_2924,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production Running,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352470 03/18/11 18:22 brads Removed 20ns from T3 signal at the TS input starting for run 2922
  • 352469 03/18/11 18:21 a-onl End_of_Run_2923
  • 352468 03/18/11 18:20 a-onl Start_Run_2923,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production Running,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352467 03/18/11 18:19 eschulte Back to Production at 5uA
  • 352466 03/18/11 18:18 a-onl End_of_Run_2922
  • 352465 03/18/11 18:16 eschulte Beam Restored
  • 352464 03/18/11 18:15 a-onl Start_Run_2922,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production Running,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352463 03/18/11 17:35 eschulte Escorted Access Requested
  • 352462 03/18/11 16:49 yawei Target Screens
  • 352461 03/18/11 16:41 eschulte Q2 investigation
  • 352460 03/18/11 16:01 adaq Beam Time Accounting
  • 352459 03/18/11 15:59 ellie Day Shift Summary
  • 352458 03/18/11 15:54 a-onl End_of_Run_2921
  • 352457 03/18/11 15:53 ellie Beam down for injector work
  • 352456 03/18/11 14:55 a-onl Start_Run_2921,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production Running,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352455 03/18/11 14:50 lselvy Permissions for scripts
  • 352454 03/18/11 14:33 ellie Beam Back!
  • 352453 03/18/11 14:27 ellie Beam Down - MCC Lost a Bunch of Magnets
  • 352452 03/18/11 14:26 a-onl End_of_Run_2918
  • 352451 03/18/11 14:22 a-onl Start_Run_2918,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production Running,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352450 03/18/11 14:19 a-onl End_of_Run_2917
  • 352449 03/18/11 14:08 xqian target screens
  • 352448 03/18/11 13:07 a-onl Start_Run_2917,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production Running,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352447 03/18/11 13:06 a-onl End_of_Run_2916
  • 352446 03/18/11 13:06 A. Camsonne Energy back to previous value after Hall C is back
  • 352445 03/18/11 12:53 ellie BB & HAND Start of Run Script Problems
  • 352444 03/18/11 12:25 ellie Hall A Tools Screen Capture
  • 352443 03/18/11 12:23 ellie Spot++ for Run 2916
  • 352442 03/18/11 12:21 auto_snap Radiation and Ozone snapshot
  • 352441 03/18/11 12:21 auto_snap Miscellaneous snapshot
  • 352440 03/18/11 12:21 auto_snap Hall A Tools snapshot
  • 352439 03/18/11 12:21 auto_snap Compton Polarimeter snapshot
  • 352438 03/18/11 12:21 auto_snap BigBox Magnet snapshot
  • 352437 03/18/11 12:21 auto_snap Beam Position and Current snapshot
  • 352436 03/18/11 11:54 a-onl Start_Run_2916,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production Running,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352435 03/18/11 11:53 a-onl End_of_Run_2915
  • 352434 03/18/11 11:51 ellie Glasses Found in Counting House
  • 352433 03/18/11 11:35 A. Camsonne Energy shifted since yesterday
  • 352432 03/18/11 10:43 ellie BHWP Moved to OUT
  • 352431 03/18/11 10:42 a-onl Start_Run_2915,Run_type=Other,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=TEST FOR START SCRIPT,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352430 03/18/11 10:38 a-onl End_of_Run_2914
  • 352429 03/18/11 10:35 xqian Target Screen
  • 352428 03/18/11 10:33 lselvy Run Script Test
  • 352427 03/18/11 10:26 a-onl Start_Run_2914,Run_type=Other,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=TEST FOR START SCRIPT,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352426 03/18/11 09:44 a-onl Start_Run_2914,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production 5 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352425 03/18/11 09:44 a-onl End_of_Run_2913
  • 352424 03/18/11 09:39 lselvy Errors explained
  • 352423 03/18/11 08:47 V. Sulkosky Bubbler Camera
  • 352422 03/18/11 08:34 narbe End of Run 2912 Called For Reset
  • 352421 03/18/11 08:32 a-onl Start_Run_2913,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production 5 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352420 03/18/11 08:29 a-onl End_of_Run_2912
  • 352419 03/18/11 08:29 a-onl End_of_Run_2912
  • 352418 03/18/11 08:14 israely Can anyone tell me how to view the gas chamers in the camera?
  • 352417 03/18/11 08:01 adaq Beam Time Accounting
  • 352416 03/18/11 07:58 danez Owl Shift Summary
  • 352415 03/18/11 07:49 israely target screenshot
  • 352414 03/18/11 07:48 israely Bigbite scintilator & dcaler screenshot
  • 352413 03/18/11 07:47 israely HAND Screenshot(2912)
  • 352412 03/18/11 07:18 siqbal target screens
  • 352411 03/18/11 07:15 a-onl Start_Run_2912,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production 5 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352410 03/18/11 07:14 a-onl End_of_Run_2911
  • 352409 03/18/11 06:18 Higinbotham Acceptance causes positiive y-target peaking of Al walls
  • 352408 03/18/11 06:09 danez Tools Screenshot
  • 352407 03/18/11 06:08 danez Spot++ Screenshot (2911)
  • 352406 03/18/11 06:00 a-onl Start_Run_2911,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production 5 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352405 03/18/11 05:59 a-onl End_of_Run_2910
  • 352404 03/18/11 04:42 rbziel HAND Screenshot(2910)
  • 352403 03/18/11 04:41 rbziel Bigbite Scintillator and Scaler Screenshots (2910)
  • 352402 03/18/11 04:38 a-onl Start_Run_2910,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production 5 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352401 03/18/11 04:37 a-onl End_of_Run_2909
  • 352400 03/18/11 04:21 auto_snap Radiation and Ozone snapshot
  • 352399 03/18/11 04:21 auto_snap Miscellaneous snapshot
  • 352398 03/18/11 04:21 auto_snap Hall A Tools snapshot
  • 352397 03/18/11 04:21 auto_snap Compton Polarimeter snapshot
  • 352396 03/18/11 04:21 auto_snap BigBox Magnet snapshot
  • 352395 03/18/11 04:21 auto_snap Beam Position and Current snapshot
  • 352394 03/18/11 03:24 a-onl Start_Run_2909,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production 5 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352393 03/18/11 03:16 a-onl End_of_Run_2908
  • 352392 03/18/11 02:47 siqbal target screens
  • 352391 03/18/11 02:08 rbziel HAND Screenshot (2908)
  • 352390 03/18/11 02:07 rbziel Bigbite Scintillator and Scaler Screenshot(2908)
  • 352389 03/18/11 02:05 a-onl Start_Run_2908,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production 5 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352388 03/18/11 02:04 a-onl End_of_Run_2907
  • 352387 03/18/11 00:53 a-onl Start_Run_2907,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production 5 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352386 03/18/11 00:52 yez Wrong cratemap?
  • 352385 03/18/11 00:51 a-onl End_of_Run_2906
  • 352384 03/18/11 00:18 siqbal target screen
  • 352383 03/18/11 00:01 adaq Beam Time Accounting
  • 352382 03/18/11 00:01 adaq Swing Shift Summary
  • Thursday

  • 352381 03/17/11 23:43 pzhu run 2906 of neutron
  • 352380 03/17/11 23:42 pzhu run 2906 of scaler,bigbite
  • 352379 03/17/11 23:41 pzhu run 2906 of RC1,Saw
  • 352378 03/17/11 23:40 a-onl Start_Run_2906,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production 5 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352377 03/17/11 23:39 a-onl End_of_Run_2905
  • 352376 03/17/11 22:56 eip z traget
  • 352375 03/17/11 22:34 pzhu 2905 run of neutron
  • 352374 03/17/11 22:33 pzhu 2905 run of scaler and bigbite
  • 352373 03/17/11 22:32 mhuang Neutron detector left side HV trip
  • 352372 03/17/11 22:32 pzhu 2905run of RC1,Saw
  • 352371 03/17/11 22:22 pzhu run 2904 not good because left side of neutron detector problem
  • 352370 03/17/11 22:21 a-onl Start_Run_2905,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production 5 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352369 03/17/11 22:20 a-onl End_of_Run_2904
  • 352368 03/17/11 22:02 A. Camsonne Turned regulation off on Left dipole
  • 352367 03/17/11 21:52 pzhu 2904 run of Neutron detctor screenshot
  • 352366 03/17/11 21:51 pzhu 2904 run of scaler and bigbite
  • 352365 03/17/11 21:50 pzhu 2904 run of RC1 and SawDTmon
  • 352364 03/17/11 21:47 a-onl Start_Run_2904,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production 5 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352363 03/17/11 21:38 lselvy Further Further Sanity Check
  • 352362 03/17/11 21:32 a-onl Start_Run_2903,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production 5 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352361 03/17/11 21:32 a-onl End_of_Run_2902
  • 352360 03/17/11 21:29 igorko analysis
  • 352359 03/17/11 21:08 lselvy Further Sanity Check
  • 352358 03/17/11 21:05 pzhu Left arm NMR not stable
  • 352357 03/17/11 20:45 lselvy Possible reason for strange peak in 1190
  • 352356 03/17/11 20:40 a-onl Start_Run_2902,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production 5 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352355 03/17/11 20:39 igorko vertex reconstruction
  • 352354 03/17/11 20:39 a-onl End_of_Run_2901
  • 352353 03/17/11 20:21 auto_snap Radiation and Ozone snapshot
  • 352352 03/17/11 20:21 auto_snap Miscellaneous snapshot
  • 352351 03/17/11 20:21 auto_snap Hall A Tools snapshot
  • 352350 03/17/11 20:21 auto_snap Compton Polarimeter snapshot
  • 352349 03/17/11 20:21 auto_snap BigBox Magnet snapshot
  • 352348 03/17/11 20:20 auto_snap Beam Position and Current snapshot
  • 352347 03/17/11 20:13 lselvy Strange data structure
  • 352346 03/17/11 19:38 pzhu RC1,prescale,sawdtmon for run 2901
  • 352345 03/17/11 19:36 pzhu bigbite and scaler for run 2901
  • 352344 03/17/11 19:35 a-onl Start_Run_2901,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production 5 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352343 03/17/11 19:35 xqian Target Screen
  • 352342 03/17/11 19:35 a-onl End_of_Run_2900
  • 352341 03/17/11 19:15 mhuang Hall A tools screen during checklist
  • 352340 03/17/11 19:12 Higinbotham Offsite Test
  • 352339 03/17/11 19:11 pzhu spot++ for checklist
  • 352338 03/17/11 19:07 A. Camsonne Test HALOG 19:07
  • 352337 03/17/11 19:06 A. Camsonne HALOG possibly fixed
  • 352336 03/17/11 18:58 A. Camsonne Test HALOG
  • 352335 03/17/11 18:53 A. Camsonne HALOG in bad shape
  • 352334 03/17/11 18:21 A. Camsonne Actually it is 2899 previous run was a spot++
  • 352333 03/17/11 18:19 A. Camsonne Actually it is 2899 previous run was a spot++
  • 352332 03/17/11 18:17 a-onl Start_Run_2900,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production 5 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352331 03/17/11 18:16 a-onl End_of_Run_2899
  • 352330 03/17/11 18:04 israely target screenshot
  • 352329 03/17/11 18:03 israely Neutron detctor screenshot
  • 352328 03/17/11 18:02 israely Bigbite scintilator & dcaler screenshot
  • 352327 03/17/11 18:00 pzhu spot++
  • 352326 03/17/11 17:54 pzhu RC1, prescale window for run 2899, 5uA, 2x2raster
  • 352325 03/17/11 17:52 mhuang Back to production run
  • 352324 03/17/11 17:51 pzhu scaler production 5uA
  • 352323 03/17/11 17:50 A. Camsonne First production run after DAQ change 2898
  • 352322 03/17/11 17:48 a-onl Start_Run_2899,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production 5 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352321 03/17/11 17:47 a-onl End_of_Run_2898
  • 352320 03/17/11 17:44 a-onl Start_Run_2898,Run_type=Spot++,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3825,rightarm_theta=33.5,rightarm_sieve=out,comment_text=spot++ run 1uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352319 03/17/11 17:43 a-onl End_of_Run_2897
  • 352318 03/17/11 17:33 orchen Fixed timing problems in F1 and 1190 TDCs
  • 352317 03/17/11 17:30 Higinbotham [CHANGE] BigBite DAQ Modified After Run 2896
  • 352316 03/17/11 17:29 mhuang Beam back
  • 352315 03/17/11 16:54 a-onl Start_Run_2897,Run_type=Other,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=13,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=electronics checks,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352314 03/17/11 16:35 a-onl End_of_Run_2896
  • 352313 03/17/11 16:33 israely Bigbite scintilator & dcaler screenshot
  • 352312 03/17/11 16:33 israely Bigbite scintilator & dcaler screenshot
  • 352311 03/17/11 16:32 israely Neutron detctor screenshot
  • 352310 03/17/11 16:31 israely target screenshot
  • 352309 03/17/11 16:30 pzhu L-arm VDC and R-arm cerenkov
  • 352308 03/17/11 16:04 a-onl Start_Run_2896,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=13,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=20cm He 8uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352307 03/17/11 16:04 adaq Beam Time Accounting
  • 352306 03/17/11 16:04 Solvignon Day shift summary
  • 352305 03/17/11 16:03 a-onl End_of_Run_2895
  • 352304 03/17/11 15:56 A. Camsonne Moving files e07006_all_1* to /adaql7/work1/e07006/adaql7
  • 352303 03/17/11 15:41 israely target screenshot
  • 352302 03/17/11 15:40 israely Neutron detctor screenshot
  • 352301 03/17/11 15:39 israely Coda screenshot
  • 352300 03/17/11 15:05 a-onl Start_Run_2895,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=13,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=20cm He 8uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352299 03/17/11 14:56 a-onl End_of_Run_2894
  • 352298 03/17/11 14:56 Solvignon Spot++
  • 352297 03/17/11 14:55 gilman target charts as we turn beam back on
  • 352296 03/17/11 14:54 a-onl Start_Run_2894,Run_type=Spot++,target_type=20cm dummy,beam_e0=4.46,beam_raster=on,beam_current=13,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=15cm dummy 5uA spot ++_,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352295 03/17/11 14:49 a-onl Start_Run_2893,Run_type=Spot++,target_type=20cm dummy,beam_e0=4.46,beam_raster=on,beam_current=13,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=15cm dummy 5uA spot ++_ ,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352294 03/17/11 14:46 Solvignon Back to production data
  • 352293 03/17/11 14:10 orchen correction - disabled by hardware, not by software
  • 352292 03/17/11 13:59 orchen T6 and T7 triggers disabled by software
  • 352291 03/17/11 13:57 orchen HAND HV trips fixed (IMPORTANT HARDWARE CHANGE!!!)
  • 352290 03/17/11 13:00 brads re: F1s do require explicit ref time subtraction
  • 352289 03/17/11 12:58 brads re: F1 reference time fun
  • 352288 03/17/11 12:20 auto_snap Radiation and Ozone snapshot
  • 352287 03/17/11 12:20 auto_snap Miscellaneous snapshot
  • 352286 03/17/11 12:20 auto_snap Hall A Tools snapshot
  • 352285 03/17/11 12:20 auto_snap Compton Polarimeter snapshot
  • 352284 03/17/11 12:20 auto_snap BigBox Magnet snapshot
  • 352283 03/17/11 12:20 auto_snap Beam Position and Current snapshot
  • 352282 03/17/11 12:01 saw F1 reference time fun
  • 352281 03/17/11 11:55 Solvignon Left Q2 at zero. No alarm
  • 352280 03/17/11 11:55 a-onl End_of_Run_2892
  • 352279 03/17/11 11:51 a-onl Start_Run_2892,Run_type=Production SRC,target_type=20cm dummy,beam_e0=4.46,beam_raster=on,beam_current=13,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=15cm dummy 5uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352278 03/17/11 11:49 gilman target status
  • 352277 03/17/11 11:40 a-onl End_of_Run_2891
  • 352276 03/17/11 11:12 Solvignon Hall A tools during checklist
  • 352275 03/17/11 11:10 Solvignon Spot++ during checklist
  • 352274 03/17/11 10:59 israely Target screenshot
  • 352273 03/17/11 10:58 saw Notes on the F1
  • 352272 03/17/11 10:58 a-onl Start_Run_2891,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=13,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=4He 20cm 1uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352271 03/17/11 10:58 israely Bigbite scintilator & dcaler screenshot
  • 352270 03/17/11 10:57 israely Neutron detector screenshot
  • 352269 03/17/11 10:55 gilman target status
  • 352268 03/17/11 10:52 saw F1 window changed from 600 to 1000 ns
  • 352267 03/17/11 10:43 Kalyan Re: Timing for current run
  • 352266 03/17/11 10:41 brads Followup to halog 352225 (BB concerns)
  • 352265 03/17/11 10:38 Solvignon Spot++
  • 352264 03/17/11 10:38 a-onl End_of_Run_2890
  • 352263 03/17/11 10:36 a-onl Start_Run_2890,Run_type=Spot++,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=13,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=spot check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352262 03/17/11 10:31 solvignon Target work done. Tuned beam
  • 352261 03/17/11 10:18 israely Starting beam recovery with BeO
  • 352260 03/17/11 10:16 fansler- RQ2 N2
  • 352259 03/17/11 10:16 israely FSD is triped
  • 352258 03/17/11 10:14 israely Target charts screenshot
  • 352257 03/17/11 10:13 israely Target screenshot
  • 352256 03/17/11 10:13 gilman high voltages changed
  • 352255 03/17/11 08:22 Solvignon HRS magnets at 0 for target re-pressurization
  • 352254 03/17/11 08:07 israely target screenshot
  • 352253 03/17/11 08:06 israely Bigbite scintilator & dcaler screenshot
  • 352252 03/17/11 08:05 israely Neutron detector screenshot
  • 352251 03/17/11 08:02 adaq Beam Time Accounting
  • 352250 03/17/11 07:58 danez Owl Shift Summary
  • 352249 03/17/11 07:37 danez HAND HV slot 5 Update 2
  • 352248 03/17/11 07:35 danez Controlled Access
  • 352247 03/17/11 07:02 a-onl End_of_Run_2889
  • 352246 03/17/11 06:36 danez Tools Screenshot
  • 352245 03/17/11 06:34 danez Spot++ Screenshot
  • 352244 03/17/11 06:18 siqbal neutron detector rates
  • 352243 03/17/11 05:56 a-onl Start_Run_2889,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=13,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production run on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352242 03/17/11 05:55 a-onl End_of_Run_2888
  • 352241 03/17/11 05:51 Higinbotham Cut on the protons
  • 352240 03/17/11 05:35 orchen following doug's instructionas
  • 352239 03/17/11 04:38 Higinbotham Check for Protons First
  • 352238 03/17/11 04:31 a-onl Start_Run_2888,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=13,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production run on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352237 03/17/11 04:30 a-onl End_of_Run_2887
  • 352236 03/17/11 04:15 orchen Change of target mass in reaplay scripts
  • 352235 03/17/11 03:42 orchen lookes like no change in the physical layout of the HV cards for HAND
  • 352234 03/17/11 03:20 aniol target moved to 20cm He
  • 352233 03/17/11 03:19 a-onl Start_Run_2887,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=13,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production run on 20cm 4He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352232 03/17/11 03:13 a-onl End_of_Run_2886
  • 352231 03/17/11 02:41 aniol target screens
  • 352230 03/17/11 02:39 aniol target screens, LD2
  • 352229 03/17/11 02:38 a-onl Start_Run_2886,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3.5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production run on 10cm LD2,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352228 03/17/11 02:27 aniol target moved to LD2, 10cm
  • 352227 03/17/11 02:22 a-onl End_of_Run_2885
  • 352226 03/17/11 02:14 igorko bigbite
  • 352225 03/17/11 02:11
  • 352224 03/17/11 01:54 danez HAND HV slot 5 Update
  • 352223 03/17/11 01:30 danez HAND HV slot 5
  • 352222 03/17/11 01:20 igorko HAND voltages
  • 352221 03/17/11 01:17 a-onl Start_Run_2885,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=13,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production run on 20cm He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352220 03/17/11 01:13 israely target screenshot
  • 352219 03/17/11 01:13 israely coda screenshot
  • 352218 03/17/11 01:12 israely Bigbite scintilator & dcaler screenshot
  • 352217 03/17/11 01:12 israely Neutron detector screenshot
  • 352216 03/17/11 01:04 israely Beam shutdown
  • 352215 03/17/11 00:59 a-onl End_of_Run_2884
  • 352214 03/17/11 00:37 eip hv dispaly
  • 352213 03/17/11 00:33 lselvy gui issues
  • 352212 03/17/11 00:26 adaq Beam Time Accounting
  • 352211 03/17/11 00:16 Ole Hansen Swing Shift Summary
  • Wednesday

  • 352210 03/16/11 23:48 narbe EB2 Crashed
  • 352209 03/16/11 23:47 Ole Hansen CODA back up
  • 352208 03/16/11 23:47 a-onl Start_Run_2884,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=13,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production run on 20cm He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352207 03/16/11 23:44 Ole Hansen CODA crash
  • 352206 03/16/11 23:39 danez HAND HV Layer 1 - PMTs off?
  • 352205 03/16/11 23:34 israely target screenshot
  • 352204 03/16/11 23:33 israely Neutron Detector Screenshot
  • 352203 03/16/11 23:33 israely Bigbite scintilator & dcaler screenshot
  • 352202 03/16/11 23:32 israely coda screenshot
  • 352201 03/16/11 22:57 narbe Correct Beam Current for 2879-2883
  • 352200 03/16/11 22:46 eip current ?
  • 352199 03/16/11 22:41 a-onl Start_Run_2883,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production run on 20cm He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352198 03/16/11 22:41 Ole Hansen HV trip
  • 352197 03/16/11 22:40 a-onl End_of_Run_2882
  • 352196 03/16/11 22:06 xqian Target Screen Shot
  • 352195 03/16/11 22:05 a-onl Start_Run_2882,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production run on 20cm He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352194 03/16/11 22:04 Ole Hansen HV trip
  • 352193 03/16/11 22:04 a-onl End_of_Run_2881
  • 352192 03/16/11 21:30 Ole Hansen HV crate replaced. Requesting beam
  • 352191 03/16/11 21:30 a-onl Start_Run_2881,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production run on 20cm He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352190 03/16/11 20:28 israely Bigbite scintilator & dcaler screenshot
  • 352189 03/16/11 20:27 israely Neutron Detector Screenshot
  • 352188 03/16/11 20:25 israely Target screenshot
  • 352187 03/16/11 20:24 israely Target screenshot
  • 352186 03/16/11 20:23 a-onl End_of_Run_2880
  • 352185 03/16/11 20:20 A. Camsonne Lost HV crate Bad 24 V supply
  • 352184 03/16/11 20:18 Ole Hansen HAND power supply failure
  • 352183 03/16/11 20:15 narbe F1-TDC ROC5 Error.
  • 352182 03/16/11 20:14 a-onl Start_Run_2880,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production run on 20cm He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352181 03/16/11 20:10 Ole Hansen ROC5 errors
  • 352180 03/16/11 20:09 a-onl End_of_Run_2879
  • 352179 03/16/11 20:00 lselvy Questions about anomalous peak
  • 352178 03/16/11 19:12 Ole Hansen Start production
  • 352177 03/16/11 19:12 a-onl Start_Run_2879,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production run on 20cm He,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352176 03/16/11 19:11 xqian Target Change
  • 352175 03/16/11 19:08 a-onl End_of_Run_2878
  • 352174 03/16/11 19:07 narbe Raster Check
  • 352173 03/16/11 19:01 a-onl Start_Run_2878,Run_type=Spot++,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Raster Check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352172 03/16/11 19:00 a-onl End_of_Run_2877
  • 352171 03/16/11 18:59 a-onl Start_Run_2877,Run_type=Spot++,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Raster Check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352170 03/16/11 18:51 Ole Hansen Setting up for production
  • 352169 03/16/11 18:46 a-onl End_of_Run_2876
  • 352168 03/16/11 18:24 a-onl Start_Run_2876,Run_type=Calibration,target_type=Optics: multi 1mm C foils,beam_e0=4.46,beam_raster=off,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=sieve,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=optics target - sieve in on Left HRS,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352167 03/16/11 18:23 a-onl End_of_Run_2875
  • 352166 03/16/11 18:17 igorko ADCs gate
  • 352165 03/16/11 18:16 igorko F1 and 1190 unclear pattern
  • 352164 03/16/11 17:52 brads re: BB adc vs. tdc
  • 352163 03/16/11 17:44 narbe Restarted CODA due to BBTS1
  • 352162 03/16/11 17:43 a-onl Start_Run_2875,Run_type=Calibration,target_type=Optics: multi 1mm C foils,beam_e0=4.46,beam_raster=off,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=sieve,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=optics target - sieve in on Left HRS,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352161 03/16/11 17:37 a-onl Start_Run_2874,Run_type=Calibration,target_type=Optics: multi 1mm C foils,beam_e0=4.46,beam_raster=off,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=sieve,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=optics target - sieve in on Left HRS,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352160 03/16/11 17:37 a-onl End_of_Run_2873
  • 352159 03/16/11 17:17 igorko BigBite issue
  • 352158 03/16/11 16:56 a-onl Start_Run_2873,Run_type=Calibration,target_type=Optics: multi 1mm C foils,beam_e0=4.46,beam_raster=off,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=sieve,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=optics target - sieve in on Left HRS,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352157 03/16/11 16:54 a-onl End_of_Run_2872
  • 352156 03/16/11 16:13 a-onl Start_Run_2872,Run_type=Calibration,target_type=Optics: multi 1mm C foils,beam_e0=4.46,beam_raster=off,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=sieve,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=optics target - sieve in on Left HRS,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352155 03/16/11 16:12 a-onl End_of_Run_2871
  • 352154 03/16/11 16:02 adaq Beam Time Accounting
  • 352153 03/16/11 16:00 Solvignon Day shift summary
  • 352152 03/16/11 15:28 a-onl Start_Run_2871,Run_type=Calibration,target_type=Optics: multi 1mm C foils,beam_e0=4.46,beam_raster=off,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=sieve,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=optics target - sieve in on Left HRS,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352151 03/16/11 15:25 A. Camsonne Left HRS Sieve in after run 2870
  • 352150 03/16/11 15:19 saw, igorko ROC5 F1 setup restored to 600ns window
  • 352149 03/16/11 15:18 A. Camsonne ROC5 rebooted after run 2870
  • 352148 03/16/11 15:16 a-onl End_of_Run_2870
  • 352147 03/16/11 15:07 a-onl Start_Run_2870,Run_type=Calibration,target_type=Optics: multi 1mm C foils,beam_e0=4.46,beam_raster=off,beam_current=10,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=optics target - F1 code modified by Steve Wood,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352146 03/16/11 15:06 A. Camsonne ROC5 rebooted after run 2869
  • 352145 03/16/11 15:05 saw, igorko F1 TDC Window changed to 1000ns temporarily
  • 352144 03/16/11 15:04 a-onl End_of_Run_2869
  • 352143 03/16/11 14:32 a-onl Start_Run_2869,Run_type=Calibration,target_type=Optics: multi 1mm C foils,beam_e0=4.46,beam_raster=off,beam_current=10,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=optics target ,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352142 03/16/11 14:26 a-onl End_of_Run_2868
  • 352141 03/16/11 13:57 Jack Left VDCs high voltage
  • 352140 03/16/11 13:55 Solvignon Hall A tools for day shift checklist
  • 352139 03/16/11 13:52 a-onl Start_Run_2868,Run_type=Calibration,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=BeO target - Pointing run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352138 03/16/11 13:49 a-onl End_of_Run_2867
  • 352137 03/16/11 13:44 a-onl Start_Run_2867,Run_type=Calibration,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=BeO target during Harp scans,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352136 03/16/11 13:38 Fansler RQ2 N2
  • 352135 03/16/11 13:14 Solvignon Beam study done. Waiting for beam.
  • 352134 03/16/11 12:55 A. Camsonne Update on beam studies
  • 352133 03/16/11 12:27 yez Replay Scripts and scripts for batch farm job submitting are all tested and work fine
  • 352132 03/16/11 12:06 igorko timing for the current run
  • 352131 03/16/11 11:55 LeRose Left NMR Lock
  • 352130 03/16/11 11:16 Fansler RQ2 N2
  • 352129 03/16/11 10:51 Fansler L Arm NMR
  • 352128 03/16/11 10:01 Solvignon Hall A in controlled access
  • 352127 03/16/11 09:40 Solvignon Beam study starts
  • 352126 03/16/11 09:35 gilman moved to slanted carbon target
  • 352125 03/16/11 09:34 Solvignon HV trip
  • 352124 03/16/11 09:22 A. Camsonne Hall C target cool down starting.
  • 352123 03/16/11 09:21 A. Camsonne Beam studies moved to today
  • 352122 03/16/11 09:20 a-onl End_of_Run_2866
  • 352121 03/16/11 09:00 israely Correction: Target was He-4 20cm 13uA
  • 352120 03/16/11 08:41 israely Target screenshot
  • 352119 03/16/11 08:40 israely Coda Screenshot
  • 352118 03/16/11 08:39 israely Bigbite scintilator & dcaler screenshot
  • 352117 03/16/11 08:39 israely Neutron Detector Screenshot
  • 352116 03/16/11 08:29 a-onl Start_Run_2866,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=13,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production With scalers PS-1 128, ps6&7-0,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352115 03/16/11 08:29 gilman target screens
  • 352114 03/16/11 08:25 a-onl End_of_Run_2865
  • 352113 03/16/11 08:24 Saha HAND ND N6 Left Problem Run 2865
  • 352112 03/16/11 08:13 V. Sulkosky Re: T5 triggers without a T5 bit
  • 352111 03/16/11 08:11 orchen neutron peak with target vertex cuts
  • 352110 03/16/11 08:02 adaq Beam Time Accounting
  • 352109 03/16/11 08:01 danez Owl Shift Summary
  • 352108 03/16/11 07:59 a-onl Start_Run_2865,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=13,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production With scalers PS-1 128, Ps3 - 3 ps7-2,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352107 03/16/11 07:55 a-onl Start_Run_2864,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=13,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production With scalers PS-1 128, Ps3 - 3 ps7-2,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352106 03/16/11 07:53 a-onl End_of_Run_2863
  • 352105 03/16/11 07:49 a-onl Start_Run_2863,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=13,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production With scalers PS-1 128, Ps3 - 3 ps7-2,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352104 03/16/11 07:48 a-onl End_of_Run_2862
  • 352103 03/16/11 07:44 orchen Low Q2 4He(e,e'pn) TOF peak ???
  • 352102 03/16/11 07:29 Higinbotham T5 vs T6 and T7
  • 352101 03/16/11 07:06 danez Tools Screenshot
  • 352100 03/16/11 07:05 danez Spot++ Screenshot (2862)
  • 352099 03/16/11 06:22 orchen T5 triggers without a T5 bit.
  • 352098 03/16/11 06:18 rbziel Bigbite Scintillator and Scaler Screenshots(2862)
  • 352097 03/16/11 06:16 rbziel HAND Screenshot(2862)
  • 352096 03/16/11 06:00 a-onl Start_Run_2862,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=13,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production With scalers PS-1 128, Ps3 - 3 ps7-2,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352095 03/16/11 05:49 a-onl End_of_Run_2861
  • 352094 03/16/11 04:33 aniol target screens
  • 352093 03/16/11 04:01 a-onl Start_Run_2861,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=13,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production With scalers PS-1 128, Ps3 - 3 ps7-2,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352092 03/16/11 04:01 a-onl End_of_Run_2860
  • 352091 03/16/11 02:48 rbziel Bigbite Scintillator and Scaler Screenshot (2860)
  • 352090 03/16/11 02:46 rbziel HAND Screenshot (2860)
  • 352089 03/16/11 02:02 a-onl Start_Run_2860,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=13,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production With scalers PS-1 128, Ps3 - 3 ps7-2,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352088 03/16/11 02:01 a-onl End_of_Run_2859
  • 352087 03/16/11 01:05 igorko current for the 20 cm He4 target
  • 352086 03/16/11 00:58 igorko T6 trigger problem
  • 352085 03/16/11 00:44 aniol JT valve opened: question
  • 352084 03/16/11 00:39 adaq Beam Time Accounting
  • 352083 03/16/11 00:37 B Norum Shift Summary
  • 352082 03/16/11 00:03 a-onl Start_Run_2859,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=13,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production With scalers PS-1 128, Ps3 - 3 ps7-2,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352081 03/16/11 00:02 a-onl End_of_Run_2858
  • Tuesday

  • 352080 03/15/11 23:52 israely Neutron Detector Screenshot
  • 352079 03/15/11 23:51 israely Bigbite scintilator & dcaler screenshot
  • 352078 03/15/11 23:50 israely Coda screenshot
  • 352077 03/15/11 23:49 israely Analysis screenshot hand 13
  • 352076 03/15/11 23:48 israely Analysis screenshot hand 13
  • 352075 03/15/11 23:48 israely Analysis screenshot hand 12
  • 352074 03/15/11 23:47 israely Analysis screenshot hand 11
  • 352073 03/15/11 23:46 israely Analysis screenshot hand 10
  • 352072 03/15/11 23:46 israely Analysis screenshot hand 9
  • 352071 03/15/11 23:45 israely Analysis screenshot hand 8
  • 352070 03/15/11 23:44 israely Analysis screenshot hand 7
  • 352069 03/15/11 23:43 israely Analysis screenshot hand 6
  • 352068 03/15/11 23:43 israely Analysis screenshot hand 5
  • 352067 03/15/11 23:42 israely Analysis screenshot hand 4
  • 352066 03/15/11 23:42 israely Analysis screenshot hand 3
  • 352065 03/15/11 23:41 israely Analysis screenshot hand 2
  • 352064 03/15/11 23:41 israely Analysis screenshot hand 1
  • 352063 03/15/11 23:39 israely Analysis screenshot bigbite 1
  • 352062 03/15/11 23:39 israely Analysis screenshot bigbite
  • 352061 03/15/11 23:39 israely Analysis screenshot bigbite 1
  • 352060 03/15/11 23:38 israely Analysis screenshot bigbite
  • 352059 03/15/11 23:37 israely Analysis screenshot bigbite 4
  • 352058 03/15/11 23:37 israely Analysis screenshot bigbite 3
  • 352057 03/15/11 23:36 israely Analysis screenshot bigbite 2
  • 352056 03/15/11 23:36 israely Analysis screenshot bigbite 1
  • 352055 03/15/11 23:34 israely Analysis screenshot 11
  • 352054 03/15/11 23:34 israely Analysis screenshot 10
  • 352053 03/15/11 23:33 israely Analysis screenshot 9
  • 352052 03/15/11 23:33 israely Analysis screenshot 8
  • 352051 03/15/11 23:32 israely Analysis screenshot 7
  • 352050 03/15/11 23:31 israely Analysis screenshot 5
  • 352049 03/15/11 23:31 israely Analysis screenshot 4
  • 352048 03/15/11 23:30 israely Analysis screenshot 3
  • 352047 03/15/11 23:29 israely Analysis screenshot 3
  • 352046 03/15/11 23:29 israely Analysis screenshot 2
  • 352045 03/15/11 23:24 israely Analysis screenshots 1
  • 352044 03/15/11 23:06 israely Target screenshot
  • 352043 03/15/11 22:57 israely Neutron Detector Screenshot
  • 352042 03/15/11 22:57 israely Bigbite scintilator & dcaler screenshot
  • 352041 03/15/11 22:56 israely Current is changed to 18uA
  • 352040 03/15/11 22:55 a-onl Start_Run_2858,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=18,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production With scalers PS-1 128, Ps3 - 3 ps7-2,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352039 03/15/11 22:53 a-onl End_of_Run_2857
  • 352038 03/15/11 22:43 igorko LHRS momentum
  • 352037 03/15/11 22:37 israely Neutron Detector Screenshot
  • 352036 03/15/11 22:37 israely Bigbite scintilator & dcaler screenshot
  • 352035 03/15/11 22:36 yawei increase loop1 JT value and decrease loop2 JT value
  • 352034 03/15/11 22:32 a-onl Start_Run_2857,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=20,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production With scalers PS-1 changed to 128, Ps3 - 3 ps7-2,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352033 03/15/11 22:31 a-onl End_of_Run_2856
  • 352032 03/15/11 22:27 a-onl Start_Run_2856,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=20,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production With scalers PS-1 changed to 128, Ps3 - 2,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352031 03/15/11 22:26 a-onl End_of_Run_2855
  • 352030 03/15/11 22:24 a-onl Start_Run_2855,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=20,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production With scalers PS-1 changed to 64,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352029 03/15/11 22:23 israely Current is changed to 20uA and code stoped
  • 352028 03/15/11 22:21 yawei increase the fan's drive frequency from 24Hz to 48Hz, Loop1
  • 352027 03/15/11 22:20 a-onl Start_Run_2854,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=20,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production With scalers PS-1 changed to 64,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352026 03/15/11 22:15 a-onl End_of_Run_2853
  • 352024 03/15/11 22:14 igorko figures in JPG format
  • 352023 03/15/11 21:47 israely Coda screenshot
  • 352022 03/15/11 21:46 V. Sulkosky Re-post plots not ps format
  • 352021 03/15/11 21:42 israely Neutron Detector Screenshot
  • 352020 03/15/11 21:41 israely Bigbite scintilator & dcaler screenshot
  • 352019 03/15/11 21:38 igorko db_DBB database
  • 352018 03/15/11 21:37 israely Scalers from run 2853
  • 352017 03/15/11 21:36 a-onl Start_Run_2853,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=10,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production With scalers PS-1 changed to 20,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352016 03/15/11 21:34 a-onl End_of_Run_2852
  • 352015 03/15/11 21:31 igorko BB scalers
  • 352014 03/15/11 21:19 israely Target screenshot
  • 352013 03/15/11 21:14 igorko timming
  • 352012 03/15/11 21:13 israely Prescales from 2852 run
  • 352011 03/15/11 21:07 a-onl Start_Run_2852,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production With scalers clock reduced to 400,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352010 03/15/11 21:05 a-onl End_of_Run_2851
  • 352009 03/15/11 21:01 israely scalers screenshot for 2851 run
  • 352008 03/15/11 20:52 norum LHRS NMR Regulation
  • 352007 03/15/11 20:50 israely Neutron Detector Screenshot
  • 352006 03/15/11 20:49 israely Production 2851
  • 352005 03/15/11 20:46 Norum LHRS no NMR lock
  • 352004 03/15/11 20:46 israely Clock prescaler is reduced to 20i
  • 352003 03/15/11 20:45 a-onl Start_Run_2851,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production With scalers clock reduced to 20,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352002 03/15/11 20:43 a-onl End_of_Run_2850
  • 352001 03/15/11 20:41 a-onl Start_Run_2850,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production With scalers,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 352000 03/15/11 20:41 A. Camsonne Left NMR unstable
  • 351999 03/15/11 20:40 israely Run 2849 is dead run
  • 351998 03/15/11 20:40 a-onl End_of_Run_2849
  • 351997 03/15/11 20:39 a-onl Start_Run_2849,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.3758,rightarm_theta=33.5,rightarm_sieve=out,comment_text=Production Without scalers,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351996 03/15/11 20:24 israely Neutron Detector Screenshot
  • 351995 03/15/11 20:23 israely Bigbite scintillator & scaler screenshot
  • 351994 03/15/11 20:17 yawei Target moved to 20cm 4He, loop 1
  • 351993 03/15/11 20:15 israely We are moving to production
  • 351992 03/15/11 20:11 Norum Tools snapshot
  • 351991 03/15/11 20:05 yawei Target moved to BeO position.
  • 351990 03/15/11 20:04 a-onl End_of_Run_2848
  • 351989 03/15/11 20:04 israely spot++ results
  • 351988 03/15/11 19:59 a-onl Start_Run_2848,Run_type=Spot++,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=Spot++,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351987 03/15/11 19:59 Jack (Techs) LQ1 power supply saga
  • 351986 03/15/11 19:55 yawei increase JT value flow, loop 1
  • 351985 03/15/11 19:52 a-onl End_of_Run_2847
  • 351984 03/15/11 19:35 a-onl Start_Run_2847,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351983 03/15/11 19:34 yawei Target moved to empty position!
  • 351982 03/15/11 19:34 a-onl End_of_Run_2846
  • 351981 03/15/11 19:24 A. Camsonne Right arm set to 1.3825 GeV/c
  • 351980 03/15/11 19:06 A. Camsonne Issue with Right Q1 power supply
  • 351979 03/15/11 18:54 a-onl Start_Run_2846,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351978 03/15/11 18:52 a-onl End_of_Run_2845
  • 351977 03/15/11 18:36 a-onl Start_Run_2845,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351976 03/15/11 18:35 israely Neutron & bigbite are back to live
  • 351975 03/15/11 18:33 israely Code is rebooted
  • 351974 03/15/11 18:25 a-onl Start_Run_2844,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351973 03/15/11 18:24 a-onl End_of_Run_2843
  • 351972 03/15/11 18:19 israely Neutron Detector Screenshot
  • 351971 03/15/11 18:18 israely Bigbite scintillator screenshotBigbite scintillator & scaler screenshot
  • 351970 03/15/11 18:07 a-onl Start_Run_2843,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351969 03/15/11 18:06 a-onl End_of_Run_2842
  • 351968 03/15/11 17:47 a-onl Start_Run_2842,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351967 03/15/11 17:46 a-onl End_of_Run_2841
  • 351966 03/15/11 17:26 a-onl Start_Run_2841,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351965 03/15/11 17:25 a-onl End_of_Run_2840
  • 351964 03/15/11 17:18 lselvy F1TDC inefficiency
  • 351963 03/15/11 17:09 a-onl Start_Run_2840,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351962 03/15/11 17:07 a-onl End_of_Run_2839
  • 351961 03/15/11 16:58 saw More F1 missing reference time fun
  • 351960 03/15/11 16:51 a-onl Start_Run_2839,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351959 03/15/11 16:50 a-onl End_of_Run_2838
  • 351958 03/15/11 16:47 israely Neutron Detector Screenshot
  • 351957 03/15/11 16:45 israely Bigbite scintillators & scalers screenshot
  • 351956 03/15/11 16:42 israely Cosmics Run
  • 351955 03/15/11 16:38 saw,lselvy More on missing F1 reference times
  • 351954 03/15/11 16:35 a-onl Start_Run_2838,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351953 03/15/11 16:33 a-onl End_of_Run_2837
  • 351952 03/15/11 16:32 a-onl Start_Run_2837,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=Cosmics (With scalers),bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351951 03/15/11 16:06 yawei Target Screens!
  • 351950 03/15/11 16:02 adaq Beam Time Accounting
  • 351949 03/15/11 15:57 ellie Day Shift Summary
  • 351948 03/15/11 15:48 eip so what does those plots mean
  • 351947 03/15/11 15:37 israely Target IOC screenshot
  • 351946 03/15/11 15:27 saw Missing F1 reference times
  • 351945 03/15/11 15:27 nmuangma BigBite TDC spectrum concern for dE lower bars
  • 351944 03/15/11 15:20 A. Camsonne Requested for Restricted Access to check Q1 power supply
  • 351943 03/15/11 15:14 ellie Hall A Tools Screengrab
  • 351942 03/15/11 15:09 lselvy Followup on the start of run change
  • 351941 03/15/11 15:00 a-onl Start_Run_2836,Run_type=Other,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=3.5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=TEST FOR RUN SCRIPT,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351940 03/15/11 14:57 lselvy Start of run script
  • 351939 03/15/11 14:43 Hanretty FPP Straw Chambers off
  • 351938 03/15/11 14:12 igorko timing
  • 351937 03/15/11 13:20 IgorKo, SMB Timing changes in T7
  • 351936 03/15/11 13:12 orchen small evidence for a T6 peak in deutrium when cutting on high momentum
  • 351935 03/15/11 12:57 israely Target Alarm screenshot
  • 351934 03/15/11 12:55 israely Target BDS Positions & Set Absolute Position screenshot
  • 351933 03/15/11 12:51 A. Camsonne HRS moved to production Kinematic angle
  • 351932 03/15/11 12:49 israely WR Valve screenshot
  • 351931 03/15/11 12:46 israely Cryotarget & ERS screenshot
  • 351930 03/15/11 12:32 israely I ment loop 3
  • 351929 03/15/11 12:10 israely Loop 3 Target charts
  • 351928 03/15/11 12:09 israely Loop 2 Target screenshots
  • 351927 03/15/11 12:08 israely Loop 2 Target screenshots
  • 351926 03/15/11 12:01 israely Loop 2 Charts screenshots
  • 351925 03/15/11 12:00 israely Loop 2 Target screenshot
  • 351924 03/15/11 11:40 israely Hall A loop 1 Fan controls screenshot
  • 351923 03/15/11 11:31 israely Loop 1 Charts
  • 351922 03/15/11 11:30 israely Loop 1 Target screenshors
  • 351921 03/15/11 11:08 israely Loop 3 screenshot
  • 351920 03/15/11 11:07 israely Loop 2 screenshots
  • 351919 03/15/11 11:06 orchen semi inclusive T7
  • 351918 03/15/11 11:00 igor semi inclusive T6 problem
  • 351917 03/15/11 10:18 averett reduce JT valve flow-loop 2
  • 351916 03/15/11 10:06 ellie Called for Controlled Access
  • 351915 03/15/11 10:01 Xiaohui RC daily summary
  • 351914 03/15/11 09:58 a-onl End_of_Run_2836
  • 351913 03/15/11 09:46 a-onl Start_Run_2836,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3.5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351912 03/15/11 09:45 a-onl End_of_Run_2835
  • 351911 03/15/11 09:41 israely Target Screenshots
  • 351910 03/15/11 09:30 israely Loop3 schematic diagram
  • 351909 03/15/11 09:10 ellie BTA from March 14 Swing Shift
  • 351908 03/15/11 09:08 Xiaohui BTA log
  • 351907 03/15/11 09:02 ellie Owl Shift BTA
  • 351906 03/15/11 08:57 a-onl Start_Run_2835,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3.5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351905 03/15/11 08:55 israely Target Screenshot
  • 351904 03/15/11 08:52 a-onl End_of_Run_2834
  • 351903 03/15/11 08:51 israely Loop 2 schematic diagram
  • 351902 03/15/11 08:50 israely Loop 1 schematic diagram
  • 351901 03/15/11 08:43 Holmstrom Neutron Rates
  • 351900 03/15/11 08:29 orchen d(e,e'pn) results for RC meating at MCC
  • 351899 03/15/11 08:10 a-onl Start_Run_2834,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3.5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351898 03/15/11 08:10 danez Owl Shift Summary
  • 351897 03/15/11 08:09 a-onl End_of_Run_2833
  • 351896 03/15/11 07:26 a-onl Start_Run_2833,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3.5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351895 03/15/11 07:25 danez ROC10 hung
  • 351894 03/15/11 07:21 a-onl End_of_Run_2832
  • 351893 03/15/11 07:15 danez BigBite Screenshot (2832)
  • 351892 03/15/11 07:14 danez HAND Screenshot (2832)
  • 351891 03/15/11 07:01 aniol target - esr temp above 16K
  • 351890 03/15/11 07:01 danez Tools Screenshot
  • 351889 03/15/11 06:53 danez Spot++ Screenshot (2832)
  • 351888 03/15/11 06:50 a-onl Start_Run_2832,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3.5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351887 03/15/11 06:49 a-onl End_of_Run_2831
  • 351886 03/15/11 06:07 a-onl Start_Run_2831,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3.5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351885 03/15/11 06:06 a-onl End_of_Run_2830
  • 351884 03/15/11 05:43 danez BigBite Screenshot (2830)
  • 351883 03/15/11 05:42 danez HAND Screenshot (2830)
  • 351882 03/15/11 05:23 a-onl Start_Run_2830,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3.5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351881 03/15/11 05:22 a-onl End_of_Run_2829
  • 351880 03/15/11 04:57 aniol target screens
  • 351879 03/15/11 04:40 a-onl Start_Run_2829,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3.5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351878 03/15/11 04:39 a-onl End_of_Run_2828
  • 351877 03/15/11 04:39 danez ROC5 error, run 2828
  • 351876 03/15/11 04:11 danez BigBite Screenshot (2828)
  • 351875 03/15/11 04:10 danez HAND Screenshot
  • 351874 03/15/11 03:56 a-onl Start_Run_2828,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3.5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351873 03/15/11 03:51 a-onl End_of_Run_2827
  • 351872 03/15/11 03:24 nmuangma F1 TDC complain message when replay
  • 351871 03/15/11 03:09 nmuangma 1190 TDC not working for run 2707 to 2799, Work after Vince widen the L1A gate start at run 2800
  • 351870 03/15/11 03:09 a-onl Start_Run_2827,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351869 03/15/11 03:08 a-onl End_of_Run_2826
  • 351868 03/15/11 02:29 danez BigBite Screenshot
  • 351867 03/15/11 02:28 danez HAND Screenshot
  • 351866 03/15/11 02:27 a-onl Start_Run_2826,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351865 03/15/11 02:26 a-onl End_of_Run_2825
  • 351864 03/15/11 02:07 aniol target - adjusted JT valve
  • 351863 03/15/11 01:45 a-onl Start_Run_2825,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351862 03/15/11 01:44 a-onl End_of_Run_2824
  • 351861 03/15/11 01:43 orchen Neutron TOF EDTM location(s)
  • 351860 03/15/11 01:04 danez BigBite Screenshot
  • 351859 03/15/11 01:03 danez HAND Screenshot
  • 351858 03/15/11 01:01 a-onl Start_Run_2824,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351857 03/15/11 01:00 a-onl End_of_Run_2823
  • 351856 03/15/11 00:18 a-onl Start_Run_2823,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351855 03/15/11 00:17 a-onl End_of_Run_2822
  • 351854 03/15/11 00:03 Norum Run Summary
  • Monday

  • 351853 03/14/11 23:59 tireman HAND scaler screen shot
  • 351852 03/14/11 23:57 yawei Target Screens!
  • 351851 03/14/11 23:35 a-onl Start_Run_2822,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351850 03/14/11 23:33 a-onl End_of_Run_2821
  • 351849 03/14/11 22:51 a-onl Start_Run_2821,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351848 03/14/11 22:50 a-onl End_of_Run_2820
  • 351847 03/14/11 22:44 tireman ROC sync plots --- bad scaling with LD2 target
  • 351846 03/14/11 22:40 tireman online replay empty plots that were not empty before
  • 351845 03/14/11 22:38 tireman Online replay issues
  • 351844 03/14/11 22:25 tireman Big bite scintillator singles screen shot
  • 351843 03/14/11 22:23 tireman Hall A scaler data screen shots
  • 351842 03/14/11 22:20 tireman Big bit singles screen snap shot
  • 351841 03/14/11 22:15 tireman Big Bite screen snap shot
  • 351840 03/14/11 22:14 tireman HAND scalar screen snap shot
  • 351839 03/14/11 22:11 yawei Target Screens!
  • 351838 03/14/11 22:08 tireman Back to data collection after CODA restart and ROC 1 and 3 Reboot
  • 351837 03/14/11 22:06 a-onl Start_Run_2820,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351836 03/14/11 22:00 tireman Run 2819 end failed
  • 351835 03/14/11 21:56 a-onl End_of_Run_2819
  • 351834 03/14/11 21:18 tireman Run 2818 is junk (no end run script either)
  • 351833 03/14/11 21:14 a-onl Start_Run_2819,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351832 03/14/11 21:04 a-onl Start_Run_2818,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351831 03/14/11 21:03 a-onl End_of_Run_2817
  • 351830 03/14/11 20:25 tireman Runs 2811 - 2813 are short for rate check on LD2
  • 351829 03/14/11 20:23 tireman Runs 2814 - 2816 are production runs
  • 351828 03/14/11 20:21 a-onl Start_Run_2817,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351827 03/14/11 20:20 a-onl End_of_Run_2816
  • 351826 03/14/11 19:38 a-onl Start_Run_2816,Run_type=Other,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351825 03/14/11 19:37 a-onl End_of_Run_2815
  • 351824 03/14/11 19:10 Norum Hall A Tools
  • 351823 03/14/11 18:53 a-onl Start_Run_2815,Run_type=Other,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=LD2 10 cm 3 microamp PS 100/1/15/1/1/4/4/15 ,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351822 03/14/11 18:52 a-onl End_of_Run_2814
  • 351821 03/14/11 18:45 yez Prepared scripts for data replay on ifarm and jobs submitting to batch farm
  • 351820 03/14/11 18:37 Xiaohui Run plan update
  • 351819 03/14/11 18:16 tireman HAND scalars with LD2 10 cm with 3 micro amp beam
  • 351818 03/14/11 18:10 a-onl Start_Run_2814,Run_type=Other,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=3,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=LD2 10 cm 3 micro amp PS 100/1/15/1/1/4/4/15 change in prescales,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351817 03/14/11 18:08 a-onl End_of_Run_2813
  • 351816 03/14/11 18:07 a-onl Start_Run_2813,Run_type=Other,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=20,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=LD2 10 cm 3 micro amp PS 100/1/12/1/1/4/4/15 change in prescales,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351815 03/14/11 18:05 a-onl End_of_Run_2812
  • 351814 03/14/11 18:05 a-onl Start_Run_2812,Run_type=Other,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=20,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=LD2 10 cm 3 micro amp for rate check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351813 03/14/11 18:02 Norum Spot Check
  • 351812 03/14/11 18:01 tireman Spot++ screen shots after switch to LD2 10 cm target
  • 351811 03/14/11 17:59 a-onl End_of_Run_2811
  • 351810 03/14/11 17:56 a-onl Start_Run_2811,Run_type=Other,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=20,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=LD2 10 cm 1 micro amp for spot check and rates,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351809 03/14/11 17:50 yawei Target to 10cm LD2
  • 351808 03/14/11 17:45 jpchen Adjusted cryo flow to have loop 2 (LD2) ready for beam.
  • 351807 03/14/11 17:42 a-onl End_of_Run_2810
  • 351806 03/14/11 16:54 a-onl Start_Run_2810,Run_type=Other,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=20,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=calibration,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351805 03/14/11 16:53 a-onl End_of_Run_2809
  • 351804 03/14/11 16:28 yawei Target screens.
  • 351803 03/14/11 16:03 adaq Beam Time Accounting
  • 351802 03/14/11 16:02 ellie Day Shift Summary
  • 351801 03/14/11 15:52 a-onl Start_Run_2809,Run_type=Other,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=20,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=calibration,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351800 03/14/11 15:51 V. Sulkosky EDTM was Enabled During Access
  • 351799 03/14/11 15:48 a-onl Start_Run_2808,Run_type=Other,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=20,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351798 03/14/11 15:47 a-onl End_of_Run_2807
  • 351797 03/14/11 15:42 a-onl Start_Run_2807,Run_type=Other,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=20,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351796 03/14/11 15:41 ellie Increasing beam current from 5uA to 20uA
  • 351795 03/14/11 15:38 a-onl End_of_Run_2806
  • 351794 03/14/11 15:37 ellie Hall A Tools Screen Capture
  • 351793 03/14/11 15:36 a-onl Start_Run_2806,Run_type=Other,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351792 03/14/11 15:32 Hanretty target change
  • 351791 03/14/11 15:30 weinstein replay questions
  • 351790 03/14/11 15:24 Hanretty target screens
  • 351789 03/14/11 15:21 a-onl End_of_Run_2805
  • 351788 03/14/11 15:07 a-onl Start_Run_2805,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351787 03/14/11 14:57 ellie Beam Down Temporarily - Hall C's Fault
  • 351786 03/14/11 14:56 a-onl End_of_Run_2804
  • 351785 03/14/11 14:28 pzhu BPM database copied to e07006/DB
  • 351784 03/14/11 14:14 A. Camsonne No more warnings for F1 in ROC5
  • 351783 03/14/11 14:10 pzhu BPM database for leftarm
  • 351782 03/14/11 14:10 pzhu BPM database for leftarm
  • 351781 03/14/11 14:10 Hanretty Heater power
  • 351780 03/14/11 14:04 a-onl Start_Run_2804,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351779 03/14/11 14:03 a-onl End_of_Run_2803
  • 351778 03/14/11 13:31 weinstein big bite rate snapshot
  • 351777 03/14/11 13:29 weinstein hand rate snapshot
  • 351776 03/14/11 13:15 V. Sulkosky Scope Trace BB Retiming
  • 351775 03/14/11 13:07 a-onl Start_Run_2803,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351774 03/14/11 13:06 a-onl End_of_Run_2802
  • 351773 03/14/11 13:06 ellie spot++ for 2802
  • 351772 03/14/11 13:03 a-onl Start_Run_2802,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351771 03/14/11 13:02 ellie BMPA Y - Change from 0.37 to 0.30
  • 351770 03/14/11 13:01 a-onl End_of_Run_2801
  • 351769 03/14/11 12:56 V. Sulkosky Neutron TOF T7 and EDTM
  • 351768 03/14/11 12:56 a-onl Start_Run_2801,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351767 03/14/11 12:54 Hanretty target screens
  • 351766 03/14/11 12:54 ellie Beam Recovery looks good, asking for 5uA
  • 351765 03/14/11 12:52 a-onl End_of_Run_2800
  • 351764 03/14/11 12:51 ellie spot++ for 2800
  • 351763 03/14/11 12:47 a-onl Start_Run_2800,Run_type=Spot++,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351762 03/14/11 12:38 ellie Beam returning, starting Short Beam Recovery
  • 351761 03/14/11 12:28 Hanretty JT valve adjustment
  • 351760 03/14/11 12:28 V. Sulkosky Checked Stops to F1 and 1190
  • 351759 03/14/11 12:27 a-onl End_of_Run_2799
  • 351758 03/14/11 12:26 V. Sulkosky BB Retiming Checked and Veto Enlarged
  • 351757 03/14/11 12:22 V. Sulkosky T6 Now Coincidence between BB and LHRS
  • 351756 03/14/11 12:07 a-onl Start_Run_2799,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=T6,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351755 03/14/11 12:04 Xiaohui RC daily summary
  • 351754 03/14/11 11:56 a-onl End_of_Run_2798
  • 351753 03/14/11 11:36 a-onl Start_Run_2798,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=T6 - redownloaded crls,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351752 03/14/11 11:36 A. Camsonne Redownloaded crls for run 2798
  • 351751 03/14/11 11:34 a-onl End_of_Run_2797
  • 351750 03/14/11 11:32 a-onl Start_Run_2797,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=T7 only with pulser additionnal ,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351749 03/14/11 11:29 a-onl End_of_Run_2796
  • 351748 03/14/11 11:29 lselvy Added Synch Event Check for 1190
  • 351747 03/14/11 11:24 a-onl Start_Run_2796,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=T8 only,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351746 03/14/11 11:12 a-onl Start_Run_2795,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=cosmic,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351745 03/14/11 11:09 a-onl End_of_Run_2794
  • 351744 03/14/11 11:04 a-onl Start_Run_2794,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=cosmic - BB only,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351743 03/14/11 10:40 Xiaohui Run plan update
  • 351742 03/14/11 10:35 ellie Left Dipole Alarm - fixing itself, but happening a few times
  • 351741 03/14/11 10:32 ellie Target moved to Loop 1, 20cm He target
  • 351740 03/14/11 10:30 igorko F1 - 1190 correlation
  • 351739 03/14/11 10:23 Hanretty Target changed
  • 351738 03/14/11 10:20 ellie Asked for Controlled Access
  • 351737 03/14/11 10:12 a-onl End_of_Run_2793
  • 351736 03/14/11 10:01 Hanretty HAND alarm
  • 351735 03/14/11 09:46 a-onl Start_Run_2793,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=cosmic,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351734 03/14/11 09:44 a-onl End_of_Run_2792
  • 351733 03/14/11 09:44 a-onl Start_Run_2792,Run_type=Other,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=20,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=junk,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351732 03/14/11 09:42 a-onl End_of_Run_2791
  • 351731 03/14/11 09:18 a-onl Start_Run_2791,Run_type=Optics,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=20,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=optics run,,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351730 03/14/11 09:17 a-onl End_of_Run_2790
  • 351729 03/14/11 09:16 a-onl Start_Run_2790,Run_type=Optics,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=20,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=optics run, ,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351728 03/14/11 09:07 weinstein coda
  • 351727 03/14/11 09:07 Hanretty Target screens
  • 351726 03/14/11 09:05 ellie Moving Left Sieve to OUT
  • 351725 03/14/11 09:04 Hanretty Hanretty
  • 351724 03/14/11 08:56 a-onl End_of_Run_2789
  • 351723 03/14/11 08:41 ellie Extra numbers in Hall A Tools?
  • 351722 03/14/11 08:05 Weinstein Hand rates
  • 351721 03/14/11 08:05 adaq Beam Time Accounting
  • 351720 03/14/11 08:04 Weinstein Bigbite rates
  • 351719 03/14/11 08:04 danez Owl Shift Summary
  • 351718 03/14/11 08:00 a-onl Start_Run_2789,Run_type=Production SRC,target_type=Optics: multi 1mm C foils,beam_e0=4.46,beam_raster=on,beam_current=2,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=sieve,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=optics run, sieve in on L-HRS, 2 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351717 03/14/11 07:59 a-onl End_of_Run_2788
  • 351716 03/14/11 07:59 a-onl Start_Run_2788,Run_type=Production SRC,target_type=Optics: multi 1mm C foils,beam_e0=4.46,beam_raster=on,beam_current=2,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=sieve,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=optics run, sieve in on L-HRS, 2 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351715 03/14/11 07:56 Hanretty Target switch
  • 351714 03/14/11 07:49 a-onl End_of_Run_2787
  • 351713 03/14/11 07:27 a-onl Start_Run_2787,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=2,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=LD run, 2 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351712 03/14/11 07:26 a-onl End_of_Run_2786
  • 351711 03/14/11 07:17 aniol target screens
  • 351710 03/14/11 07:09 danez Tools Screenshot
  • 351709 03/14/11 07:08 danez Spot++ Screenshot
  • 351708 03/14/11 07:00 a-onl Start_Run_2786,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=2,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=LD run, 2 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351707 03/14/11 06:59 a-onl End_of_Run_2785
  • 351706 03/14/11 06:31 a-onl Start_Run_2785,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=2,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=LD run, 2 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351705 03/14/11 06:30 a-onl End_of_Run_2784
  • 351704 03/14/11 06:21 SMB run 2784, ROCs
  • 351703 03/14/11 06:13 SMB run 2784, LD@2uA, rates on BB
  • 351702 03/14/11 06:11 SMB run 2784, LD@2uA, rates on HAND
  • 351701 03/14/11 06:07 a-onl Start_Run_2784,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=2,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=LD run, 2 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351700 03/14/11 06:03 a-onl End_of_Run_2783
  • 351699 03/14/11 05:38 a-onl Start_Run_2783,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=2,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=LD run, 2 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351698 03/14/11 05:37 a-onl End_of_Run_2782
  • 351697 03/14/11 05:11 a-onl Start_Run_2782,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=2,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=LD run, 2 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351696 03/14/11 04:42 a-onl Start_Run_2781,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=2,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=LD run, 2 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351695 03/14/11 04:41 a-onl End_of_Run_2780
  • 351694 03/14/11 04:36 SMB Run 2777, online GUI, ROCs:
  • 351693 03/14/11 04:29 SMB online GUI, BB - the "step" problem
  • 351692 03/14/11 04:15 a-onl Start_Run_2780,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=2,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=LD run, 2 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351691 03/14/11 04:14 a-onl End_of_Run_2779
  • 351690 03/14/11 04:12 a-onl End_of_Run_2778
  • 351689 03/14/11 03:49 a-onl Start_Run_2779,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=2,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=LD run, 2 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351688 03/14/11 03:46 orchen Updated results on d(e,e'pn) measurement at x>1
  • 351687 03/14/11 03:22 a-onl Start_Run_2778,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=2,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=LD run, 2 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351686 03/14/11 03:21 a-onl End_of_Run_2777
  • 351685 03/14/11 02:55 a-onl Start_Run_2777,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=2,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=LD run, 2 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351684 03/14/11 02:53 a-onl End_of_Run_2776
  • 351683 03/14/11 02:52 aniol target screens
  • 351682 03/14/11 02:28 a-onl Start_Run_2776,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=2,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=LD run, 2 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351681 03/14/11 02:27 a-onl End_of_Run_2775
  • 351680 03/14/11 02:01 a-onl Start_Run_2775,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=2,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=LD run, 2 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351679 03/14/11 02:00 a-onl End_of_Run_2774
  • 351678 03/14/11 01:54 SMB run 2774, rates on hand
  • 351677 03/14/11 01:52 SMB run 2774, rates on BB
  • 351676 03/14/11 01:47 SMB run 2770, online GUI, BB
  • 351675 03/14/11 01:37 SMB Run 2770, online GUI, HRS, hand
  • 351674 03/14/11 01:35 a-onl Start_Run_2774,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=2,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=LD run, 2 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351673 03/14/11 01:33 a-onl End_of_Run_2773
  • 351672 03/14/11 01:08 a-onl Start_Run_2773,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=2,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=LD run, 2 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351671 03/14/11 01:06 a-onl End_of_Run_2772
  • 351670 03/14/11 00:44 a-onl Start_Run_2772,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=2,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=LD run, 2 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351669 03/14/11 00:43 a-onl End_of_Run_2771
  • 351668 03/14/11 00:23 a-onl Start_Run_2771,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1.5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=LD run, 1.5 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351667 03/14/11 00:21 a-onl End_of_Run_2770
  • 351666 03/14/11 00:10 adaq Beam Time Accounting
  • 351665 03/14/11 00:09 Fatiha Shift Summary
  • Sunday

  • 351664 03/13/11 23:53 Ole Hansen Analyzer updated
  • 351663 03/13/11 23:51 a-onl Start_Run_2770,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=ps6=0, ps7=2, ps1=10, ps3=5, production run?? Don't know yet. RC and Igorko still discussing~~~~,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351662 03/13/11 23:32 igorko first tripple for X>1 ?
  • 351661 03/13/11 23:29 Fatiha T5 rates tests
  • 351660 03/13/11 23:24 LeRose Target Screens
  • 351659 03/13/11 23:21 a-onl End_of_Run_2769
  • 351658 03/13/11 23:18 a-onl Start_Run_2769,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=ps6=0, ps7=2, ps1=10, ps3=5, production run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351657 03/13/11 23:16 yawei beam off trigger rates screens
  • 351656 03/13/11 23:09 a-onl End_of_Run_2768
  • 351655 03/13/11 22:50 a-onl Start_Run_2768,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=ps6=0, ps7=2, ps1=10, ps3=5, production run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351654 03/13/11 22:49 a-onl End_of_Run_2767
  • 351653 03/13/11 22:22 yawei rates screen(neutron detector)
  • 351652 03/13/11 22:20 yawei rates screens
  • 351651 03/13/11 22:18 V. Sulkosky EDTM disabled
  • 351650 03/13/11 22:11 a-onl Start_Run_2767,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=ps6=0, ps7=2, ps1=10, ps3=5, production run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351649 03/13/11 22:10 a-onl End_of_Run_2766
  • 351648 03/13/11 21:33 a-onl Start_Run_2766,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=ps6=0, ps7=2, ps1=10, ps3=5, production run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351647 03/13/11 21:32 a-onl End_of_Run_2765
  • 351646 03/13/11 21:25 V. Sulkosky Run 2706 Started on 3/11/2011
  • 351645 03/13/11 20:53 a-onl Start_Run_2765,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=ps6=0, ps7=2, ps1=10, ps3=5, production run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351644 03/13/11 20:53 a-onl End_of_Run_2764
  • 351643 03/13/11 20:17 a-onl Start_Run_2764,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=ps6=0, ps7=2, ps1=10, ps3=5, production run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351642 03/13/11 20:16 a-onl End_of_Run_2763
  • 351641 03/13/11 20:02 yawei production prescale setting
  • 351640 03/13/11 19:54 nmuangma BigBite Problems Unsolve, summary what we have done.
  • 351639 03/13/11 19:38 a-onl Start_Run_2763,Run_type=Production SRC,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=ps6=0, ps7=2, ps1=10, ps3=5, production run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351638 03/13/11 19:35 a-onl End_of_Run_2762
  • 351637 03/13/11 19:31 nmuangma run 2762 after long roc 5 power out.
  • 351636 03/13/11 19:29 a-onl Start_Run_2762,Run_type=Check out,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=roc 5 power cycle test, another one, ps6=1000, ps7=2, ps1=6, ps3=1,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351635 03/13/11 19:18 nmuangma Roc 5 another power off (long 5 mins)
  • 351634 03/13/11 19:16 LeRose Target Screens
  • 351633 03/13/11 19:15 a-onl End_of_Run_2761
  • 351632 03/13/11 18:55 a-onl Start_Run_2761,Run_type=Check out,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=rate test run, ps6=65535, ps7=2, ps1=6, ps3=1,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351631 03/13/11 18:53 a-onl End_of_Run_2760
  • 351630 03/13/11 18:49 yawei rate screens, 1uA, 10cm LD
  • 351629 03/13/11 18:43 A. Camsonne Sync clock ok but still working on ROC 5 TDC ADC correlation
  • 351628 03/13/11 18:33 a-onl Start_Run_2760,Run_type=Check out,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=rate test run, ps6=10000, ps7=65535, ps1=6, ps3=1,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351627 03/13/11 18:32 a-onl End_of_Run_2759
  • 351626 03/13/11 18:14 a-onl Start_Run_2759,Run_type=Check out,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=rate test run, ps6=ps7=65535, ps1=5, ps3=1,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351624 03/13/11 18:12 a-onl End_of_Run_2758
  • 351623 03/13/11 18:12 a-onl Start_Run_2758,Run_type=Check out,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=rate test run, ps6=ps7=65535, ps1=ps2=1,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351622 03/13/11 18:11 a-onl End_of_Run_2757
  • 351621 03/13/11 18:09 yawei similar shape of spot ++
  • 351620 03/13/11 18:04 a-onl Start_Run_2757,Run_type=Check out,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=rate test run, ps6=ps7=65535,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351619 03/13/11 18:02 a-onl End_of_Run_2756
  • 351618 03/13/11 17:59 a-onl Start_Run_2756,Run_type=Check out,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=check rate T5,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351617 03/13/11 17:57 a-onl End_of_Run_2755
  • 351616 03/13/11 17:41 A. Camsonne Also uncommented clear of SIS for ROC5 - sync looks ok now
  • 351615 03/13/11 17:39 yawei spot ++ check 2x2 raster
  • 351614 03/13/11 17:34 LeRose Dipole NMR lock
  • 351613 03/13/11 17:23 Fatiha LHRS D1-NMR not locked while its current looks fine: what you should do?
  • 351612 03/13/11 17:18 nmuangma ADC with TDC cut
  • 351611 03/13/11 17:17 a-onl Start_Run_2755,Run_type=Spot++,target_type=LD 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=spot ++, target is LD 10cm, 1uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351610 03/13/11 17:17 igorko F1 TDC
  • 351609 03/13/11 17:15 nmuangma TDC & ADC still out of sync after the roc 5 power reboot
  • 351608 03/13/11 17:14 LeRose Target: 10 cm LD2
  • 351607 03/13/11 17:09 a-onl End_of_Run_2754
  • 351606 03/13/11 17:03 nmuangma 2753 & 2754 roc 5 power cycle
  • 351605 03/13/11 17:01 a-onl Start_Run_2754,Run_type=Check out,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=roc 5 power cycle, test run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351604 03/13/11 17:00 a-onl End_of_Run_2753
  • 351603 03/13/11 17:00 a-onl Start_Run_2753,Run_type=Check out,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=roc 5 power cycle, test run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351602 03/13/11 16:59 nmuangma power cycle roc 5
  • 351601 03/13/11 16:55 igorko ROC 5 TDCs
  • 351600 03/13/11 16:53 a-onl End_of_Run_2752
  • 351599 03/13/11 16:49 Fatiha Asked for 1uA for rates and spot check
  • 351598 03/13/11 16:48 a-onl Start_Run_2752,Run_type=Spot++,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=spot ++ run, 20 cm he4, 1uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351597 03/13/11 16:38 a-onl End_of_Run_2751
  • 351596 03/13/11 16:31 LeRose Target 20 cm He
  • 351595 03/13/11 16:14 Fatiha Hall A Genaral Tool
  • 351594 03/13/11 16:10 a-onl Start_Run_2751,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=production - After ROCs reboot,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351593 03/13/11 16:09 a-onl End_of_Run_2750
  • 351592 03/13/11 16:03 A. Camsonne commented 0xdeaddead word in ROC1 and ROC2
  • 351591 03/13/11 16:03 adaq Beam Time Accounting
  • 351590 03/13/11 16:02 bogdanw shift summary
  • 351589 03/13/11 15:58 A. Camsonne commented 0xdeaddead word in ROC1 and ROC2
  • 351588 03/13/11 15:57 a-onl Start_Run_2750,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=production - After ROCs reboot,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351587 03/13/11 15:51 Xiaohui RC daily summary
  • 351586 03/13/11 15:32 a-onl End_of_Run_2748
  • 351585 03/13/11 15:23 chench summarization for the beam position scan
  • 351584 03/13/11 14:57 tireman Target Screens
  • 351583 03/13/11 14:57 a-onl Start_Run_2748,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351582 03/13/11 14:55 a-onl End_of_Run_2747
  • 351581 03/13/11 14:52 chench Screenshot of BB scintillator single rate at beam current is 1uA
  • 351580 03/13/11 14:51 chench Screenshot for the neutron detector single rate
  • 351579 03/13/11 14:50 a-onl Start_Run_2747,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351578 03/13/11 14:47 a-onl End_of_Run_2746
  • 351577 03/13/11 14:44 chench Neutron Detector single rates
  • 351576 03/13/11 14:43 chench BB scintillator single rate
  • 351575 03/13/11 14:28 chench ROC5 error
  • 351574 03/13/11 14:28 a-onl Start_Run_2746,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=25,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351573 03/13/11 14:24 V. Sulkosky Fast Clock Roc 2 vs 5
  • 351572 03/13/11 14:23 a-onl End_of_Run_2745
  • 351571 03/13/11 14:21 V. Sulkosky ROC 5 Possible Sync Issue
  • 351570 03/13/11 14:19 chench UN2745 BB wire chamber HVs are on
  • 351569 03/13/11 14:18 a-onl Start_Run_2745,Run_type=Other,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=25,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=BB VDC HV is on,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351568 03/13/11 14:16 chench run 2744 BB mwdc HV is off
  • 351567 03/13/11 14:15 a-onl End_of_Run_2744
  • 351566 03/13/11 14:12 a-onl Start_Run_2744,Run_type=Other,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=25,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=BB VDC HV is off,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351565 03/13/11 14:11 chench Spot_R after beam come back
  • 351564 03/13/11 14:11 V. Sulkosky T6 Now Singles for Wire Chamber checks
  • 351563 03/13/11 14:07 a-onl End_of_Run_2743
  • 351562 03/13/11 14:07 V. Sulkosky Escorted Access to Check T6 LHRS+BB Trigger
  • 351561 03/13/11 14:05 a-onl Start_Run_2743,Run_type=Spot++,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=beam recovery test with beam current 5uA ,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351560 03/13/11 13:54 V. Sulkosky EDTM Re-enabled for T6 check
  • 351559 03/13/11 13:21 eip 10 vs 20 cm target
  • 351558 03/13/11 13:06 orchen test run with BB chambers off
  • 351557 03/13/11 13:02 Xiaohui BPM nominal value adjusted!
  • 351556 03/13/11 13:00 a-onl End_of_Run_2742
  • 351555 03/13/11 13:00 igorko vertex
  • 351554 03/13/11 12:57 igorko vertex reconstruction for 20cm
  • 351553 03/13/11 12:49 a-onl Start_Run_2742,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=25,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=PRODUCTION without BB VDC,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351552 03/13/11 12:47 chench Online GUI run2742
  • 351551 03/13/11 12:45 a-onl End_of_Run_2741
  • 351550 03/13/11 12:42 chench Spot_R from run2741
  • 351549 03/13/11 12:36 a-onl Start_Run_2741,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=25,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=PRODUCTION,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351548 03/13/11 12:34 a-onl End_of_Run_2740
  • 351547 03/13/11 12:23 V. Sulkosky EDTM Disabled
  • 351546 03/13/11 12:15 bogdanw VDC efficiency
  • 351545 03/13/11 12:11 chench spot_R from run2741
  • 351544 03/13/11 12:11 igorko vertex reconstruction
  • 351543 03/13/11 12:09 yez some wires of Left HRS VDC planes have slightly low efficiency
  • 351542 03/13/11 12:08 a-onl Start_Run_2740,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=25,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=PRODUCTION,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351541 03/13/11 12:00 a-onl End_of_Run_2739
  • 351540 03/13/11 11:59 chench spot_R from run2739
  • 351539 03/13/11 11:57 Higinbotham Try L.gold.y
  • 351538 03/13/11 11:56 a-onl Start_Run_2739,Run_type=Spot++,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=beam test run raster 8.7*8.7mcc beam shifted by +2mm on BPMA/B,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351537 03/13/11 11:55 chench spot_R from run2737
  • 351536 03/13/11 11:54 a-onl End_of_Run_2738
  • 351535 03/13/11 11:53 a-onl Start_Run_2738,Run_type=Spot++,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=beam test run raster 8.7*8.7mcc beam shifted by +2mm on BPMA/B ,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351534 03/13/11 11:52 a-onl End_of_Run_2737
  • 351533 03/13/11 11:52 chench spot_R from run2736
  • 351532 03/13/11 11:49 a-onl Start_Run_2737,Run_type=Spot++,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=beam test run raster 8.7*8.7mcc beam shifted by +2mm on BPMA/B junk beam is not stable,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351531 03/13/11 11:46 chench Spot_R run2736
  • 351530 03/13/11 11:46 a-onl End_of_Run_2736
  • 351529 03/13/11 11:45 a-onl Start_Run_2736,Run_type=Spot++,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=beam test run raster 8.7*8.7mcc beam shifted by1mm on BPMA/B,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351528 03/13/11 11:43 a-onl End_of_Run_2735
  • 351527 03/13/11 11:42 igorko vertex reconstruction problem
  • 351526 03/13/11 11:41 chench Spot_R from run2735
  • 351525 03/13/11 11:38 a-onl Start_Run_2735,Run_type=Spot++,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=beam test run raster 8.7*8.7mcc,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351524 03/13/11 11:34 chench Run2732 is junk
  • 351523 03/13/11 11:30 chench Spot_R from run2734
  • 351522 03/13/11 11:28 a-onl End_of_Run_2734
  • 351521 03/13/11 11:27 a-onl Start_Run_2734,Run_type=Spot++,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=beam test run raster 8*5.3mcc,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351520 03/13/11 11:25 chench Spot_R from run2733 beam current is iuA and
  • 351519 03/13/11 11:23 a-onl End_of_Run_2733
  • 351518 03/13/11 11:21 a-onl Start_Run_2733,Run_type=Spot++,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=beam test run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351517 03/13/11 11:21 tireman Screen Shots and a suggestion
  • 351516 03/13/11 11:20 a-onl End_of_Run_2732
  • 351515 03/13/11 11:19 a-onl Start_Run_2732,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=25,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=He4 10cm tgt, current 25uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351514 03/13/11 11:17 a-onl End_of_Run_2731
  • 351513 03/13/11 11:13 a-onl Start_Run_2731,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=25,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=He4 10cm tgt, current 25uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351512 03/13/11 11:12 a-onl End_of_Run_2730
  • 351511 03/13/11 11:03 chench Spot++ from run2730
  • 351510 03/13/11 10:46 a-onl Start_Run_2730,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=25,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=He4 10cm tgt, current 25uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351509 03/13/11 10:45 a-onl End_of_Run_2729
  • 351508 03/13/11 10:21 chench Screenshot of BB scintillator single rate
  • 351507 03/13/11 10:20 chench screenshot of HAND detector single rate
  • 351506 03/13/11 10:19 a-onl Start_Run_2729,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=25,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=He4 10cm tgt, current 25uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351505 03/13/11 10:18 a-onl End_of_Run_2728
  • 351504 03/13/11 09:47 a-onl Start_Run_2728,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=25,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=He4 10cm tgt, current 25uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351503 03/13/11 09:46 a-onl End_of_Run_2727
  • 351502 03/13/11 09:33 A. Camsonne Commented L.* and R.* in HRS.odef
  • 351501 03/13/11 09:26 chench screenshot of BB scitillator single rate
  • 351500 03/13/11 09:25 SMB rates on BB and HAND
  • 351499 03/13/11 09:23 chench screenshot of detectors single rate
  • 351498 03/13/11 09:19 A. Camsonne Disks most likely full, moving files of run less than 2500 to /adaql7/work1/e07006
  • 351497 03/13/11 09:18 a-onl Start_Run_2727,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=25,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=He4 10cm tgt, current 25uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351496 03/13/11 09:13 a-onl End_of_Run_2726
  • 351495 03/13/11 09:06 chench onlie GUI of run2725 for hand
  • 351494 03/13/11 09:03 chench run2725 onlie GUI for BB
  • 351493 03/13/11 08:50 adaq HRS-L NMR locked
  • 351492 03/13/11 08:46 a-onl Start_Run_2726,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=25,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=He4 10cm tgt, current 25uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351491 03/13/11 08:45 a-onl End_of_Run_2725
  • 351490 03/13/11 08:39 adaq screenshot of bigbite scitillators single rate
  • 351489 03/13/11 08:38 adaq screenshot of neutron detectors single rate
  • 351488 03/13/11 08:25 adaq Tools
  • 351487 03/13/11 08:20 a-onl Start_Run_2725,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=25,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=He4 10cm tgt, current 25uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351486 03/13/11 08:17 a-onl End_of_Run_2724
  • 351485 03/13/11 08:17 tireman Target Screens
  • 351484 03/13/11 08:06 chench screenshot of bigbite scintillator single rate
  • 351483 03/13/11 08:04 chench Screenshot of Neutron Detectors single rate
  • 351482 03/13/11 08:02 adaq Beam Time Accounting
  • 351481 03/13/11 07:58 danez Owl Shift Summary
  • 351480 03/13/11 07:43 a-onl Start_Run_2724,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=25,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=He4 10cm tgt, current 25uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351479 03/13/11 07:42 a-onl End_of_Run_2723
  • 351478 03/13/11 07:41 a-onl Start_Run_2723,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=25,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=He4 10cm tgt, current 25uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351477 03/13/11 07:36 a-onl End_of_Run_2722
  • 351476 03/13/11 07:34 SMB run 2717: online GUI, LHRS
  • 351475 03/13/11 07:21 SMB run 2717: online GUI, BB,
  • 351474 03/13/11 07:13 SMB run 2717: online GUI, BB
  • 351473 03/13/11 07:10 danez Tools Screenshot
  • 351472 03/13/11 07:04 danez Spot++ Screenshot
  • 351471 03/13/11 07:04 a-onl Start_Run_2722,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=12,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=He4 20cm tgt, current 12uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351470 03/13/11 07:02 a-onl End_of_Run_2721
  • 351469 03/13/11 06:54 a-onl Start_Run_2721,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=He4 20cm tgt, current 1uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351468 03/13/11 06:50 a-onl End_of_Run_2720
  • 351467 03/13/11 06:49 orchen un reliable reasults from first 4He runs
  • 351466 03/13/11 06:45 a-onl Start_Run_2720,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=12,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=He4 20cm tgt,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351465 03/13/11 06:38 Kai target moved to He4 20cm one
  • 351464 03/13/11 06:34 SMB run 2717: online GUI, HAND
  • 351463 03/13/11 06:30 SMB Run 2717: online Gui, HAND
  • 351462 03/13/11 06:25 SMB online GUI, run 2717
  • 351461 03/13/11 06:21 kai target moved to He4
  • 351460 03/13/11 06:19 a-onl End_of_Run_2719
  • 351459 03/13/11 06:02 a-onl Start_Run_2719,Run_type=Other,target_type=Slanted Carbon,beam_e0=4.46,beam_raster=on,beam_current=25,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=Carbon target for BB test,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351458 03/13/11 05:50 kai target moved to Slanted Carbon
  • 351457 03/13/11 05:49 a-onl End_of_Run_2718
  • 351456 03/13/11 05:48 kai target snapshot
  • 351455 03/13/11 05:36 SMB online replay
  • 351454 03/13/11 05:23 a-onl Start_Run_2718,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=25,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=25 uA beam,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351453 03/13/11 05:21 a-onl End_of_Run_2717
  • 351452 03/13/11 05:17 SMB rates on BB, run 2717
  • 351451 03/13/11 05:16 orchen Very Slow replays on adaq computers
  • 351450 03/13/11 05:16 SMB rates on HAND, run 2717
  • 351449 03/13/11 05:09 SMB online Gui - problems
  • 351448 03/13/11 04:53 a-onl Start_Run_2717,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=25,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=25 uA beam,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351447 03/13/11 04:52 a-onl End_of_Run_2716
  • 351446 03/13/11 04:51 a-onl Start_Run_2716,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=25,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=25 uA beam,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351445 03/13/11 04:49 a-onl End_of_Run_2715
  • 351444 03/13/11 04:46 auto_snap Radiation and Ozone snapshot
  • 351443 03/13/11 04:46 auto_snap Miscellaneous snapshot
  • 351442 03/13/11 04:46 auto_snap Hall A Tools snapshot
  • 351441 03/13/11 04:46 auto_snap Compton Polarimeter snapshot
  • 351440 03/13/11 04:46 auto_snap BigBox Magnet snapshot
  • 351439 03/13/11 04:46 auto_snap Beam Position and Current snapshot
  • 351438 03/13/11 04:45 SMB Run 2715: rates on BB
  • 351437 03/13/11 04:43 SMB rates on hand run 2715
  • 351436 03/13/11 04:18 a-onl Start_Run_2715,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=25,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=25 uA beam,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351435 03/13/11 04:18 danez Left Dipole Momentum Setting Again
  • 351434 03/13/11 04:13 kai step close JT valve
  • 351433 03/13/11 04:11 a-onl End_of_Run_2714
  • 351432 03/13/11 04:10 danez Tools Screenshot
  • 351431 03/13/11 04:10 a-onl Start_Run_2714,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=25,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=25 uA beam,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351430 03/13/11 04:07 orchen Problem with the BB wire chambers(?)
  • 351429 03/13/11 04:05 danez Gas Shed Methane Low Alarm Sounding
  • 351428 03/13/11 04:04 a-onl End_of_Run_2713
  • 351427 03/13/11 04:04 a-onl Start_Run_2713,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=15,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=15 uA beam,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351426 03/13/11 04:02 danez HAND HV Tripping
  • 351425 03/13/11 03:59 orchen Problem with the BB wire chambers
  • 351424 03/13/11 03:57 a-onl End_of_Run_2712
  • 351423 03/13/11 03:54 a-onl Start_Run_2712,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=15,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=15 uA beam,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351422 03/13/11 03:52 orchen No events in T6
  • 351421 03/13/11 03:50 a-onl End_of_Run_2711
  • 351420 03/13/11 03:47 a-onl Start_Run_2711,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=10,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=10 uA beam,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351419 03/13/11 03:46 a-onl End_of_Run_2710
  • 351418 03/13/11 03:45 a-onl Start_Run_2710,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=5 uA beam,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351417 03/13/11 03:40 kai heat power drops to zero due to MCC chamber calibration
  • 351416 03/13/11 03:25 a-onl End_of_Run_2709
  • 351415 03/13/11 03:22 danez Alarm from Left Dipole
  • 351414 03/13/11 03:17 danez Spot++ for larger raster
  • 351413 03/13/11 03:13 a-onl Start_Run_2709,Run_type=Spot++,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=spot++ run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351412 03/13/11 03:10 a-onl End_of_Run_2708
  • 351411 03/13/11 03:07 danez Spot++ 4He 2x2
  • 351410 03/13/11 03:04 a-onl Start_Run_2708,Run_type=Spot++,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=spot++ run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351409 03/13/11 01:47 danez Spot++ from BeO
  • 351408 03/13/11 01:42 a-onl End_of_Run_2707
  • 351407 03/13/11 01:41 a-onl Start_Run_2707,Run_type=Spot++,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.95,leftarm_theta=16,leftarm_sieve=out,rightarm_p=0.952,rightarm_theta=42.5,rightarm_sieve=out,comment_text=spot++ run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351406 03/13/11 01:36 auto_snap Radiation and Ozone snapshot
  • 351405 03/13/11 01:36 auto_snap Miscellaneous snapshot
  • 351404 03/13/11 01:36 auto_snap Hall A Tools snapshot
  • 351403 03/13/11 01:36 auto_snap Compton Polarimeter snapshot
  • 351402 03/13/11 01:36 auto_snap BigBox Magnet snapshot
  • 351401 03/13/11 01:36 auto_snap Beam Position and Current snapshot
  • 351400 03/13/11 00:04 kelleher Swing Shift Summary -- March 12
  • 351399 03/13/11 00:03 adaq Beam Time Accounting
  • Saturday

  • 351398 03/12/11 23:39 orchen HV setting for N5-3L is back to its original value
  • 351397 03/12/11 23:32 orchen current HV settings for BB and HAND
  • 351396 03/12/11 23:27 kelleher target to empty
  • 351395 03/12/11 23:26 orchen curent tresholds
  • 351394 03/12/11 23:18 Xiaohui Left Q1 stays at 3.29 GeV
  • 351393 03/12/11 22:51 eip run plan -commnet
  • 351392 03/12/11 22:38 Xiaohui Run plan update
  • 351391 03/12/11 22:29 Jack (Techs) LQ1 work
  • 351390 03/12/11 22:27 Xiaohui Scattering chamber alarm back on
  • 351389 03/12/11 22:23 Jack (Techs) Left Q1 status
  • 351388 03/12/11 20:12 folts checklist/tech
  • 351387 03/12/11 19:55 kelleher target to BeO
  • 351386 03/12/11 17:36 auto_snap Radiation and Ozone snapshot
  • 351385 03/12/11 17:36 auto_snap Miscellaneous snapshot
  • 351384 03/12/11 17:36 auto_snap Hall A Tools snapshot
  • 351383 03/12/11 17:36 auto_snap Compton Polarimeter snapshot
  • 351382 03/12/11 17:36 auto_snap BigBox Magnet snapshot
  • 351381 03/12/11 17:36 auto_snap Beam Position and Current snapshot
  • 351380 03/12/11 17:15 Xiaohui Run plan (Saturday swing - Sunday day)
  • 351379 03/12/11 16:02 bogdanw shift summary
  • 351378 03/12/11 16:01 adaq Beam Time Accounting
  • 351377 03/12/11 16:01 LeRose Reduced Q1 Optics
  • 351376 03/12/11 15:41 tireman Target Screens
  • 351375 03/12/11 12:36 orchen New trigger setup for T6 and T7
  • 351374 03/12/11 12:34 a-onl End_of_Run_2706
  • 351373 03/12/11 11:53 tireman Target Screens
  • 351372 03/12/11 11:18 jpchen cryotarget Loop 1 PID adjusted
  • 351371 03/12/11 10:05 tireman target motion after camera alignment and focus
  • 351370 03/12/11 09:59 tireman Target Cameras and Target Motion
  • 351369 03/12/11 09:36 auto_snap Radiation and Ozone snapshot
  • 351368 03/12/11 09:36 auto_snap Miscellaneous snapshot
  • 351367 03/12/11 09:36 auto_snap Hall A Tools snapshot
  • 351366 03/12/11 09:36 auto_snap Compton Polarimeter snapshot
  • 351365 03/12/11 09:36 auto_snap BigBox Magnet snapshot
  • 351364 03/12/11 09:36 auto_snap Beam Position and Current snapshot
  • 351363 03/12/11 08:48 Xiaohui Target camera
  • 351362 03/12/11 08:48 Xiaohui Target camera
  • 351361 03/12/11 08:07 tireman Target screen shots and comment
  • 351360 03/12/11 06:31 M Defurne Close JT valve of LD2
  • 351359 03/12/11 05:37 M Defurne screenshots
  • 351358 03/12/11 03:26 M Defurne Maybe Should Increase the minor alarm for heat power
  • 351357 03/12/11 01:49 M Defurne screenshots
  • 351356 03/12/11 01:35 auto_snap Radiation and Ozone snapshot
  • 351355 03/12/11 01:35 auto_snap Miscellaneous snapshot
  • 351354 03/12/11 01:35 auto_snap Hall A Tools snapshot
  • 351353 03/12/11 01:35 auto_snap Compton Polarimeter snapshot
  • 351352 03/12/11 01:35 auto_snap BigBox Magnet snapshot
  • 351351 03/12/11 01:35 auto_snap Beam Position and Current snapshot
  • Friday

  • 351350 03/11/11 23:42 zwzhao target screen
  • 351349 03/11/11 23:36 zwzhao shift summary
  • 351348 03/11/11 23:24 zwzhao gas sensor alarm
  • 351347 03/11/11 21:53 zwzhao target screen
  • 351346 03/11/11 20:19 zwzhao target screen
  • 351345 03/11/11 19:50 a-onl Start_Run_2706,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB HV ON:Threshold (5.0,5.5)V,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351344 03/11/11 19:49 a-onl End_of_Run_2705
  • 351343 03/11/11 19:40 a-onl Start_Run_2705,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB HV OFF: noise study change Threshold from (4.5,5.5)V to (5.0,5.5)V,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351342 03/11/11 19:39 a-onl End_of_Run_2704
  • 351341 03/11/11 19:39 nmuangma BB MWDC threshold set from (4.5V,5.5V) to (5.0V,5V)
  • 351340 03/11/11 19:34 zwzhao target sceen
  • 351339 03/11/11 19:20 a-onl Start_Run_2704,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB HV OFF: noise study,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351338 03/11/11 19:18 a-onl End_of_Run_2703
  • 351337 03/11/11 19:18 a-onl Start_Run_2703,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=cosmics BB HV OFF: noise study,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351336 03/11/11 19:16 a-onl Start_Run_2702,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=cosmics,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351335 03/11/11 19:16 a-onl End_of_Run_2702
  • 351334 03/11/11 19:14 nmuangma BigBite issue: The noise pattern are difference. Might be related to the way coda record data.
  • 351333 03/11/11 18:26 a-onl End_of_Run_2701
  • 351332 03/11/11 18:26 a-onl Start_Run_2701,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=cosmics,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351331 03/11/11 18:22 a-onl End_of_Run_2700
  • 351330 03/11/11 18:07 a-onl Start_Run_2700,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BPM pedestal run 2698 - Bigbite all on now,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351329 03/11/11 18:05 nmuangma(Tai LT7-11 change to LTex3-11
  • 351328 03/11/11 17:52 a-onl Start_Run_2699,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BPM pedestal run 2698 - Bigbite all on now,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351327 03/11/11 17:52 nmuangma LT7-9 change to LTex3-9
  • 351326 03/11/11 17:51 a-onl End_of_Run_2698
  • 351325 03/11/11 17:40 A. Camsonne BPM pedestal runs done - iocse10 set back to autogain
  • 351324 03/11/11 17:36 a-onl Start_Run_2698,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BPM pedestal run 2698 - Bigbite all on now,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351323 03/11/11 17:35 auto_snap Radiation and Ozone snapshot
  • 351322 03/11/11 17:35 auto_snap Miscellaneous snapshot
  • 351321 03/11/11 17:35 a-onl End_of_Run_2697
  • 351320 03/11/11 17:35 auto_snap Hall A Tools snapshot
  • 351319 03/11/11 17:35 auto_snap Compton Polarimeter snapshot
  • 351318 03/11/11 17:35 auto_snap BigBox Magnet snapshot
  • 351317 03/11/11 17:35 auto_snap Beam Position and Current snapshot
  • 351316 03/11/11 17:30 A. Camsonne BPM pedestal run 2697
  • 351315 03/11/11 17:30 nmuangm TEST: ALL BigBite Voltage is "ON": MWDC HV & threshold, dE & E HV
  • 351314 03/11/11 17:30 a-onl Start_Run_2697,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BPM pedestal run 2697 - Bigbite all on now,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351313 03/11/11 17:29 a-onl End_of_Run_2696
  • 351312 03/11/11 17:27 A. Camsonne BPM pedestal run 2696
  • 351311 03/11/11 17:26 a-onl Start_Run_2696,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BPM pedestal run 2696,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351310 03/11/11 16:57 nmuangma TEST: ALL BigBite Voltage is off: MWDC HV & threshold, dE & E HV
  • 351309 03/11/11 16:45 a-onl End_of_Run_2695
  • 351308 03/11/11 16:30 zwzhao target screen
  • 351307 03/11/11 16:18 Jack (Techs) Left Q1
  • 351306 03/11/11 15:43 B Anderson Target screens.
  • 351305 03/11/11 14:35 adaq Beam Time Accounting
  • 351304 03/11/11 14:31 Weinstein Day shift summary
  • 351303 03/11/11 14:08 Xiaohui Update
  • 351302 03/11/11 14:03 B Anderson Target screens
  • 351301 03/11/11 13:36 a-onl Start_Run_2695,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=Cosmic run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351300 03/11/11 13:31 Weinstein HRS set for SRC, except Q1
  • 351299 03/11/11 13:27 a-onl End_of_Run_2694
  • 351298 03/11/11 13:23 a-onl Start_Run_2694,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351297 03/11/11 13:22 a-onl End_of_Run_2693
  • 351296 03/11/11 12:09 B Anderson Target screens
  • 351295 03/11/11 10:56 V. Sulkosky Updated Target Info in db_run.dat Script
  • 351294 03/11/11 10:42 Xiaohui Cosmics check
  • 351293 03/11/11 10:17 B Anderson Target screens
  • 351292 03/11/11 10:11 pzhu Cavity BCM test cabling
  • 351291 03/11/11 09:35 auto_snap Radiation and Ozone snapshot
  • 351290 03/11/11 09:35 auto_snap Miscellaneous snapshot
  • 351289 03/11/11 09:35 auto_snap Hall A Tools snapshot
  • 351288 03/11/11 09:35 auto_snap Compton Polarimeter snapshot
  • 351287 03/11/11 09:35 auto_snap BigBox Magnet snapshot
  • 351286 03/11/11 09:34 auto_snap Beam Position and Current snapshot
  • 351285 03/11/11 09:11 Xiaohui Counting house AC is out
  • 351284 03/11/11 09:03 a-onl Start_Run_2693,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351283 03/11/11 08:55 a-onl End_of_Run_2692
  • 351282 03/11/11 08:39 Weinstein Hall C target cooling starting
  • 351281 03/11/11 08:32 B Anderson Target screens
  • 351280 03/11/11 06:54 aniol target screens
  • 351279 03/11/11 01:34 auto_snap Radiation and Ozone snapshot
  • 351278 03/11/11 01:34 auto_snap Miscellaneous snapshot
  • 351277 03/11/11 01:34 auto_snap Hall A Tools snapshot
  • 351276 03/11/11 01:34 auto_snap Compton Polarimeter snapshot
  • 351275 03/11/11 01:34 auto_snap BigBox Magnet snapshot
  • 351274 03/11/11 01:34 auto_snap Beam Position and Current snapshot
  • 351273 03/11/11 01:10 a-onl Start_Run_2692,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351272 03/11/11 01:03 nmuangma The Noise issue in MWDC is understanded/fixed.
  • 351271 03/11/11 00:39 a-onl End_of_Run_2691
  • 351270 03/11/11 00:22 a-onl Start_Run_2691,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351269 03/11/11 00:21 nmuangma BigBite HV and Threshold Setting (-1600,-1600)V and (4.5V, 5.5V)
  • 351268 03/11/11 00:19 a-onl End_of_Run_2690
  • 351267 03/11/11 00:13 aniol target screens
  • 351266 03/11/11 00:11 a-onl Start_Run_2690,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351265 03/11/11 00:08 nmuangma MWDC noise with MWDC_HV off
  • 351264 03/11/11 00:03 a-onl End_of_Run_2689
  • Thursday

  • 351263 03/10/11 23:52 a-onl Start_Run_2689,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351262 03/10/11 23:51 a-onl End_of_Run_2688
  • 351261 03/10/11 23:37 a-onl Start_Run_2688,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351260 03/10/11 23:32 a-onl End_of_Run_2687
  • 351259 03/10/11 22:57 mhuang Target screens
  • 351258 03/10/11 22:05 nmuangma MWDC partially fixed.
  • 351257 03/10/11 21:52 a-onl Start_Run_2687,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351256 03/10/11 21:51 a-onl End_of_Run_2686
  • 351255 03/10/11 21:46 a-onl Start_Run_2686,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351254 03/10/11 21:38 a-onl End_of_Run_2685
  • 351253 03/10/11 21:31 a-onl Start_Run_2685,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351252 03/10/11 21:30 a-onl End_of_Run_2684
  • 351251 03/10/11 21:07 a-onl Start_Run_2684,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351250 03/10/11 21:07 a-onl End_of_Run_2683
  • 351249 03/10/11 21:04 igorko Counting House Temperature Issue
  • 351248 03/10/11 20:53 a-onl Start_Run_2683,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351247 03/10/11 20:53 a-onl End_of_Run_2682
  • 351246 03/10/11 20:48 a-onl Start_Run_2682,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351245 03/10/11 20:44 a-onl End_of_Run_2681
  • 351244 03/10/11 20:35 a-onl Start_Run_2681,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351243 03/10/11 20:32 a-onl End_of_Run_2680
  • 351242 03/10/11 20:24 a-onl Start_Run_2680,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351241 03/10/11 20:21 a-onl Start_Run_2679,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351240 03/10/11 20:16 a-onl Start_Run_2678,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351239 03/10/11 20:15 a-onl End_of_Run_2677
  • 351238 03/10/11 20:04 a-onl Start_Run_2677,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351237 03/10/11 20:04 a-onl End_of_Run_2676
  • 351236 03/10/11 19:52 a-onl Start_Run_2676,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351235 03/10/11 19:45 a-onl End_of_Run_2675
  • 351234 03/10/11 19:34 a-onl Start_Run_2675,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351233 03/10/11 19:29 mhuang Target screens
  • 351232 03/10/11 19:24 a-onl End_of_Run_2674
  • 351231 03/10/11 19:05 a-onl Start_Run_2674,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351230 03/10/11 18:52 a-onl End_of_Run_2673
  • 351229 03/10/11 18:34 a-onl Start_Run_2673,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351228 03/10/11 18:10 mhuang Target camera
  • 351227 03/10/11 17:51 a-onl End_of_Run_2672
  • 351226 03/10/11 17:38 Xiaohui Summary & plan
  • 351225 03/10/11 17:34 auto_snap Radiation and Ozone snapshot
  • 351224 03/10/11 17:34 auto_snap Miscellaneous snapshot
  • 351223 03/10/11 17:34 auto_snap Hall A Tools snapshot
  • 351222 03/10/11 17:34 auto_snap Compton Polarimeter snapshot
  • 351221 03/10/11 17:34 auto_snap BigBox Magnet snapshot
  • 351220 03/10/11 17:34 auto_snap Beam Position and Current snapshot
  • 351219 03/10/11 17:22 a-onl Start_Run_2672,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351218 03/10/11 16:52 jpchen Cryotarget cool-down completed
  • 351217 03/10/11 16:31 a-onl End_of_Run_2671
  • 351216 03/10/11 16:14 a-onl Start_Run_2671,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351215 03/10/11 16:14 a-onl End_of_Run_2670
  • 351214 03/10/11 15:49 a-onl Start_Run_2670,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351213 03/10/11 15:48 a-onl End_of_Run_2669
  • 351212 03/10/11 15:42 a-onl Start_Run_2669,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351211 03/10/11 15:38 a-onl End_of_Run_2668
  • 351210 03/10/11 15:26 a-onl Start_Run_2668,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check change front threshold back to 4.0V,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351209 03/10/11 15:25 a-onl End_of_Run_2667
  • 351208 03/10/11 15:15 nmuangma Power cycle both Roc9 & Roc10: check LT power supply (Good)
  • 351207 03/10/11 15:13 a-onl Start_Run_2667,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check change front threshold from 4.0 to 4.5V,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351206 03/10/11 15:12 a-onl End_of_Run_2666
  • 351205 03/10/11 15:03 a-onl Start_Run_2666,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB MWDC check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351204 03/10/11 15:02 a-onl End_of_Run_2665
  • 351203 03/10/11 14:53 a-onl Start_Run_2665,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=All detectors are on,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351202 03/10/11 14:52 a-onl End_of_Run_2664
  • 351201 03/10/11 14:51 a-onl Start_Run_2664,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=All detectors are on,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351200 03/10/11 14:40 a-onl End_of_Run_2663
  • 351199 03/10/11 14:17 a-onl Start_Run_2663,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=All detectors are on,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351198 03/10/11 14:03 a-onl End_of_Run_2662
  • 351197 03/10/11 13:55 a-onl Start_Run_2662,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=All detectors are on,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351196 03/10/11 13:31 yez R-HRS: re-arrangement of cables at Patch Panel (both up and downstairs) for PS#19R, Cer#5,#7
  • 351195 03/10/11 13:27 a-onl End_of_Run_2661
  • 351194 03/10/11 12:20 a-onl Start_Run_2661,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=All detectors are on,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351193 03/10/11 12:16 V. Sulkosky Updated DBB Database For ROC 5
  • 351192 03/10/11 12:15 a-onl End_of_Run_2660
  • 351191 03/10/11 11:33 V. Sulkosky F1TDC Issue (fixed)
  • 351190 03/10/11 10:35 a-onl Start_Run_2660,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=All detectors are on,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351189 03/10/11 10:30 a-onl End_of_Run_2659
  • 351188 03/10/11 10:29 lselvy Roc Issues
  • 351187 03/10/11 09:45 R. Michaels replay warnings in recent runs
  • 351186 03/10/11 09:42 lselvy Known problems
  • 351185 03/10/11 09:41 V. Sulkosky BB T0 calibration Definitely needed
  • 351184 03/10/11 09:34 auto_snap Radiation and Ozone snapshot
  • 351183 03/10/11 09:34 auto_snap Miscellaneous snapshot
  • 351182 03/10/11 09:34 auto_snap Hall A Tools snapshot
  • 351181 03/10/11 09:34 auto_snap Compton Polarimeter snapshot
  • 351180 03/10/11 09:34 auto_snap BigBox Magnet snapshot
  • 351179 03/10/11 09:33 auto_snap Beam Position and Current snapshot
  • 351178 03/10/11 09:13 Xiaohui Status update
  • 351177 03/10/11 08:18 igorko cosmic data
  • 351176 03/10/11 08:17 a-onl Start_Run_2659,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=All detectors are on,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351175 03/10/11 08:17 a-onl End_of_Run_2658
  • 351174 03/10/11 07:56 rshneor BigBite wirechambers missing wires looks like T0 calibration problem
  • 351173 03/10/11 07:33 orchen BB:: More indocation that this is not (only) a HW problem
  • 351172 03/10/11 07:00 orchen replay warinings from run 2658
  • 351171 03/10/11 01:33 auto_snap Radiation and Ozone snapshot
  • 351170 03/10/11 01:33 auto_snap Miscellaneous snapshot
  • 351169 03/10/11 01:33 auto_snap Hall A Tools snapshot
  • 351168 03/10/11 01:33 auto_snap Compton Polarimeter snapshot
  • 351167 03/10/11 01:33 auto_snap BigBox Magnet snapshot
  • 351166 03/10/11 01:33 auto_snap Beam Position and Current snapshot
  • Wednesday

  • 351165 03/09/11 22:46 Jack (Techs) Magnet work continues
  • 351164 03/09/11 19:44 eip rock 10 slot 1
  • 351163 03/09/11 18:21 a-onl Start_Run_2658,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=All detectors are on,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351162 03/09/11 18:20 a-onl End_of_Run_2657
  • 351161 03/09/11 17:58 igorko Chamber 1, problem?
  • 351160 03/09/11 17:55 a-onl Start_Run_2657,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=All detectors are on,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351159 03/09/11 17:54 a-onl End_of_Run_2656
  • 351158 03/09/11 17:50 pzhu BCM Calibration
  • 351157 03/09/11 17:33 auto_snap Radiation and Ozone snapshot
  • 351156 03/09/11 17:33 auto_snap Miscellaneous snapshot
  • 351155 03/09/11 17:33 auto_snap Hall A Tools snapshot
  • 351154 03/09/11 17:33 auto_snap Compton Polarimeter snapshot
  • 351153 03/09/11 17:33 auto_snap BigBox Magnet snapshot
  • 351152 03/09/11 17:33 auto_snap Beam Position and Current snapshot
  • 351151 03/09/11 16:36 lselvy Roc 10 out of bounds issue
  • 351150 03/09/11 16:35 Xiaohui Summary & plan
  • 351149 03/09/11 16:27 V. Sulkosky T6 BB Singles at 650 Hz
  • 351148 03/09/11 16:24 V. Sulkosky Scope Traces
  • 351147 03/09/11 16:21 V. Sulkosky T3 Delay Changed this Morning
  • 351146 03/09/11 16:15 yez L->L1A comes ~50ns earlier than retiming, R->L1A comes ~ 25 to 75ns before retiming
  • 351145 03/09/11 16:07 V. Sulkosky T7 Trigger details
  • 351144 03/09/11 15:58 V. Sulkosky T6 Trigger details
  • 351143 03/09/11 15:55 a-onl Start_Run_2656,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=All detectors are on,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351142 03/09/11 15:54 a-onl End_of_Run_2655
  • 351141 03/09/11 15:52 V. Sulkosky Bigbite retiming for Coincidences is setup
  • 351140 03/09/11 15:34 V. Sulkosky Trigger Alignment at TS
  • 351139 03/09/11 15:30 V. Sulkosky T1 and T3 Delay Changed on Weldment
  • 351138 03/09/11 15:12 lselvy ROC 10 Issues
  • 351137 03/09/11 15:06 igorko cosmic data
  • 351136 03/09/11 15:04 a-onl Start_Run_2655,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=All detectors are on,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351135 03/09/11 15:03 a-onl End_of_Run_2654
  • 351134 03/09/11 15:03 a-onl Start_Run_2654,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=All detectors are on,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351133 03/09/11 15:02 a-onl End_of_Run_2653
  • 351132 03/09/11 14:57 Xiaohui Status update
  • 351131 03/09/11 14:54 a-onl Start_Run_2653,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB and HAND prescale to 0,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351130 03/09/11 14:53 a-onl End_of_Run_2652
  • 351129 03/09/11 14:46 a-onl Start_Run_2652,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB test. ,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351128 03/09/11 14:44 a-onl End_of_Run_2651
  • 351127 03/09/11 14:43 a-onl Start_Run_2651,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=HAND test - all bar on except N5-3R,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351126 03/09/11 14:39 a-onl End_of_Run_2650
  • 351125 03/09/11 14:38 a-onl Start_Run_2650,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=HAND singles only from Layer 5 and 6,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351124 03/09/11 14:36 a-onl End_of_Run_2649
  • 351123 03/09/11 14:33 a-onl Start_Run_2649,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=HAND singles only from Layer 6,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351122 03/09/11 14:31 a-onl End_of_Run_2648
  • 351121 03/09/11 14:27 a-onl Start_Run_2648,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=timing checks,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351120 03/09/11 14:26 a-onl End_of_Run_2647
  • 351119 03/09/11 14:16 a-onl Start_Run_2647,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=timing checks,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351118 03/09/11 13:48 a-onl End_of_Run_2646
  • 351117 03/09/11 13:31 a-onl Start_Run_2646,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=timing checks,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351116 03/09/11 13:17 Xiaohui Target update
  • 351115 03/09/11 12:25 V. Sulkosky T6 and T7 Now Coincidence with LHRS
  • 351114 03/09/11 12:05 igorko HAND and BB HV is on
  • 351113 03/09/11 12:03 a-onl End_of_Run_2645
  • 351112 03/09/11 11:15 a-onl Start_Run_2645,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=timing checks,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351111 03/09/11 11:14 igorko HAND HV off
  • 351110 03/09/11 11:13 a-onl End_of_Run_2644
  • 351109 03/09/11 10:41 yez L-HRS: L1A comes 60ns earlier than retiming signal
  • 351108 03/09/11 10:39 a-onl Start_Run_2644,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=timing checks,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351107 03/09/11 10:38 a-onl End_of_Run_2643
  • 351106 03/09/11 10:34 Xiaohui Target update
  • 351105 03/09/11 10:34 a-onl Start_Run_2643,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=timing checks,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351104 03/09/11 10:33 a-onl End_of_Run_2642
  • 351103 03/09/11 09:54 orchen New HAND self trigger
  • 351102 03/09/11 09:33 auto_snap Radiation and Ozone snapshot
  • 351101 03/09/11 09:33 auto_snap Miscellaneous snapshot
  • 351100 03/09/11 09:33 auto_snap Hall A Tools snapshot
  • 351099 03/09/11 09:33 auto_snap Compton Polarimeter snapshot
  • 351098 03/09/11 09:33 auto_snap BigBox Magnet snapshot
  • 351097 03/09/11 09:32 auto_snap Beam Position and Current snapshot
  • 351096 03/09/11 08:57 Xiaohui RC daily summary
  • 351095 03/09/11 08:34 Xiaohui Target vacuum update
  • 351094 03/09/11 08:14 orchen ROC-10 warnings in replay
  • 351093 03/09/11 08:08 a-onl Start_Run_2642,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=timing checks,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351092 03/09/11 08:00 orchen Log cosmic run
  • 351091 03/09/11 07:59 a-onl End_of_Run_2641
  • 351090 03/09/11 06:16 eip nn vs np
  • 351089 03/09/11 01:32 auto_snap Radiation and Ozone snapshot
  • 351088 03/09/11 01:32 auto_snap Miscellaneous snapshot
  • 351087 03/09/11 01:32 auto_snap Hall A Tools snapshot
  • 351086 03/09/11 01:32 auto_snap Compton Polarimeter snapshot
  • 351085 03/09/11 01:32 auto_snap BigBox Magnet snapshot
  • 351084 03/09/11 01:32 auto_snap Beam Position and Current snapshot
  • Tuesday

  • 351083 03/08/11 21:43 V. Sulkosky Cosmic Run 2641
  • 351082 03/08/11 21:41 a-onl Start_Run_2641,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=HRS, BB and HAND cosmics,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351081 03/08/11 21:40 a-onl End_of_Run_2640
  • 351080 03/08/11 21:34 V. Sulkosky BB MWDC HV On
  • 351079 03/08/11 21:19 V. Sulkosky Checked BB ADC Gates against Coincidence
  • 351078 03/08/11 21:10 V. Sulkosky Trigger Arrivals at Trigger supevisor
  • 351077 03/08/11 21:04 V. Sulkosky T3 Delay Changed
  • 351076 03/08/11 21:01 V. Sulkosky T1 Window Reduced
  • 351075 03/08/11 20:59 V. Sulkosky BB HV turned on
  • 351074 03/08/11 20:58 a-onl Start_Run_2640,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=HRS, BB and HAND cosmics,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351073 03/08/11 20:37 a-onl End_of_Run_2639
  • 351072 03/08/11 20:30 a-onl Start_Run_2639,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=timing check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351071 03/08/11 20:19 orchen FIRS triple coincedence (e,e'pp) & (e,e'pn) for 12C x<1 run
  • 351070 03/08/11 20:06 a-onl End_of_Run_2635
  • 351069 03/08/11 19:26 yez HRSs Detectors update
  • 351068 03/08/11 19:07 a-onl Start_Run_2635,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=timing check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351067 03/08/11 18:58 V. Sulkosky CODA Failure
  • 351066 03/08/11 18:55 a-onl End_of_Run_2633
  • 351065 03/08/11 18:22 a-onl Start_Run_2633,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=timing check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351064 03/08/11 18:14 a-onl End_of_Run_2631
  • 351063 03/08/11 18:13 Xiaohui Status update <reformat>
  • 351062 03/08/11 18:06 Xiaohui Status update
  • 351061 03/08/11 18:00 yez L-HRS: Delay retiming signal ~10ns and remove the cable of L1A that go into delay module
  • 351060 03/08/11 17:47 a-onl Start_Run_2631,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=timing check,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351059 03/08/11 17:32 auto_snap Radiation and Ozone snapshot
  • 351058 03/08/11 17:32 auto_snap Miscellaneous snapshot
  • 351057 03/08/11 17:32 auto_snap Hall A Tools snapshot
  • 351056 03/08/11 17:32 auto_snap Compton Polarimeter snapshot
  • 351055 03/08/11 17:32 auto_snap BigBox Magnet snapshot
  • 351054 03/08/11 17:32 auto_snap Beam Position and Current snapshot
  • 351053 03/08/11 17:29 yez L-HRS: Double check that the L1A come 10ns later than retming
  • 351052 03/08/11 17:12 a-onl Start_Run_2629,Run_type=Cosmics,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=Fixing BigBite Rate,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351051 03/08/11 17:11 a-onl End_of_Run_2628
  • 351050 03/08/11 17:11 yez reboot ROC1
  • 351049 03/08/11 17:07 a-onl Start_Run_2628,Run_type=Cosmics,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=Fixing BigBite Rate,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351048 03/08/11 17:03 hsmith Dipole polarity changed right arm
  • 351047 03/08/11 16:45 lselvy Busy Signals
  • 351046 03/08/11 16:44 lselvy F1TDC Issue
  • 351045 03/08/11 15:45 lselvy Scaler Gating on the Right Arm
  • 351044 03/08/11 15:35 V. Sulkosky BB HV turned off
  • 351043 03/08/11 15:26 a-onl Start_Run_2626,Run_type=Cosmics,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=Fixing BigBite Rate,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351042 03/08/11 14:50 a-onl End_of_Run_2624
  • 351041 03/08/11 14:50 a-onl Start_Run_2624,Run_type=Cosmics,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=Fixing BigBite Rate,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351040 03/08/11 14:40 tech andrew polarity change
  • 351039 03/08/11 14:35 a-onl End_of_Run_2622
  • 351038 03/08/11 14:15 saw Analyzer problem, not F1TDC failure
  • 351037 03/08/11 13:00 a-onl Start_Run_2622,Run_type=Cosmics,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=Fixing BigBite Rate,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351036 03/08/11 12:52 a-onl End_of_Run_2621
  • 351035 03/08/11 12:37 a-onl Start_Run_2621,Run_type=Cosmics,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=Fixing BigBite Rate,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351034 03/08/11 12:36 a-onl End_of_Run_2620
  • 351033 03/08/11 12:36 a-onl Start_Run_2620,Run_type=Cosmics,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=Fixing BigBite Rate,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351032 03/08/11 12:05 yez LHRS: L1A comes later then retiming signal
  • 351031 03/08/11 12:05 lselvy Scaler Gating on the Left Arm
  • 351030 03/08/11 11:58 saw Killed a runaway "prescale"
  • 351029 03/08/11 11:56 yez L-HRS: All Cerenkov ADC and TDC signals are back. T4 with Cer is enabled
  • 351028 03/08/11 11:56 R. Michaels clearing scalers: try at go
  • 351027 03/08/11 11:56 V. Sulkosky T1 and T3 Window Widths
  • 351026 03/08/11 11:53 V. Sulkosky Re-enabled EDTM for SRC setup correction
  • 351025 03/08/11 11:45 a-onl Start_Run_2619,Run_type=Cosmics,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=COSMIC RUN.,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351024 03/08/11 11:27 pzhu HAPPEX cavity electronic cable bcm part exchanged
  • 351023 03/08/11 11:24 adaq L-HRS: Turn on HVs of Cerenkov
  • 351022 03/08/11 11:21 V. Sulkosky Scaler Gating
  • 351021 03/08/11 11:11 a-onl Start_Run_2618,Run_type=Cosmics,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=COSMIC RUN.,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351020 03/08/11 11:09 a-onl End_of_Run_2617
  • 351019 03/08/11 11:06 Jack Gas Cherenkov is in and S0 is out of Left Detector Stack
  • 351018 03/08/11 10:54 a-onl Start_Run_2617,Run_type=Cosmics,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=COSMIC RUN.,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351017 03/08/11 10:43 a-onl Start_Run_2616,Run_type=Cosmics,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=COSMIC RUN.,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351016 03/08/11 10:42 a-onl End_of_Run_2615
  • 351015 03/08/11 10:35 a-onl Start_Run_2615,Run_type=Cosmics,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=COSMIC RUN.,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351014 03/08/11 10:34 Xiaohui HRS polarity change
  • 351013 03/08/11 10:32 a-onl End_of_Run_2614
  • 351012 03/08/11 10:30 a-onl Start_Run_2614,Run_type=Cosmics,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=t8 run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351011 03/08/11 10:29 a-onl End_of_Run_2613
  • 351010 03/08/11 10:23 a-onl Start_Run_2613,Run_type=Cosmics,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=junk,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351009 03/08/11 10:21 a-onl End_of_Run_2612
  • 351008 03/08/11 10:20 a-onl Start_Run_2612,Run_type=Cosmics,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=Cosmic - new HAND self trigger,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351007 03/08/11 10:15 a-onl Start_Run_2611,Run_type=Cosmics,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=Cosmic - new HAND self trigger,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351006 03/08/11 10:10 a-onl Start_Run_2610,Run_type=Cosmics,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=Cosmic - new HAND self trigger,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351005 03/08/11 10:10 a-onl End_of_Run_2610
  • 351004 03/08/11 10:08 a-onl End_of_Run_2609
  • 351003 03/08/11 10:06 a-onl Start_Run_2609,Run_type=Cosmics,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=JUNK,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 351002 03/08/11 10:04 a-onl End_of_Run_2608
  • 351001 03/08/11 10:03 jmammei target scattering chamber alarm
  • 351000 03/08/11 09:32 auto_snap Radiation and Ozone snapshot
  • 350999 03/08/11 09:32 auto_snap Miscellaneous snapshot
  • 350998 03/08/11 09:32 auto_snap Hall A Tools snapshot
  • 350997 03/08/11 09:32 auto_snap Compton Polarimeter snapshot
  • 350996 03/08/11 09:31 auto_snap BigBox Magnet snapshot
  • 350995 03/08/11 09:31 auto_snap Beam Position and Current snapshot
  • 350994 03/08/11 09:00 jmammei plan for overnight Tuesday
  • 350993 03/08/11 09:00 V. Sulkosky Re-enabled EDTM for SRC setup
  • 350992 03/08/11 08:50 V. Sulkosky Zeroed T1 and T3 delays
  • 350991 03/08/11 08:46 Igorko HAND HV Trip S13.2
  • 350990 03/08/11 08:32 jmammei RC Daily Summary
  • 350989 03/08/11 08:28 Sharon Beck run 2608 online analysis
  • 350988 03/08/11 08:08 danez Owl Shift Summary
  • 350987 03/08/11 08:03 adaq Beam Time Accounting
  • 350986 03/08/11 07:06 a-onl Start_Run_2608,Run_type=Cosmics,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=Noo Beam, cosmic run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350985 03/08/11 07:00 a-onl End_of_Run_2607
  • 350984 03/08/11 06:51 danez Target Screenshot
  • 350983 03/08/11 06:03 Sharon Beck single rates in HAND
  • 350982 03/08/11 06:00 a-onl Start_Run_2607,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=10,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=Carbon run,beam 10uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350981 03/08/11 05:58 a-onl End_of_Run_2606
  • 350980 03/08/11 05:58 Sharon Beck online analysis
  • 350979 03/08/11 05:53 nmuangma F1TDC Failure
  • 350978 03/08/11 05:32 danez Tools Screenshot
  • 350977 03/08/11 05:13 a-onl Start_Run_2606,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=10,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=Carbon run,beam 10uA, prescales changed to prev. values (#2603),bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350976 03/08/11 05:10 a-onl End_of_Run_2605
  • 350975 03/08/11 05:08 a-onl Start_Run_2605,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=Carbon run, reduce beam to 5 uA, prescales changed,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350974 03/08/11 05:07 a-onl End_of_Run_2604
  • 350973 03/08/11 05:04 a-onl Start_Run_2604,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=Carbon run, reduce beam to 5 uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350972 03/08/11 04:59 a-onl End_of_Run_2603
  • 350971 03/08/11 04:32 a-onl Start_Run_2603,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=10,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=Carbon run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350970 03/08/11 04:31 a-onl End_of_Run_2602
  • 350969 03/08/11 04:13 Sharon Beck changes in online histograms
  • 350968 03/08/11 04:09 danez Target Screenshot
  • 350967 03/08/11 03:56 danez Left HRS S1 HV Adjustment
  • 350966 03/08/11 03:52 danez Prescales on White Board
  • 350965 03/08/11 03:43 a-onl Start_Run_2602,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=10,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=Carbon run ,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350964 03/08/11 03:39 a-onl End_of_Run_2601
  • 350963 03/08/11 03:24 Sharon Beck HAND
  • 350962 03/08/11 03:20 nmuangma Current BigBite Setting: threshold(4.0,4.5) -1650,-1600
  • 350961 03/08/11 03:14 a-onl Start_Run_2601,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=Carbon run with rates ~80Hz,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350960 03/08/11 03:12 a-onl End_of_Run_2600
  • 350959 03/08/11 03:11 a-onl Start_Run_2600,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=Carbon run with rates ~20Hz,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350958 03/08/11 03:09 a-onl End_of_Run_2599
  • 350957 03/08/11 03:00 danez Prescales for Run 2599
  • 350956 03/08/11 02:47 a-onl Start_Run_2599,Run_type=Production SRC,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=Carbon run with rates ~300Hz,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350955 03/08/11 02:35 a-onl End_of_Run_2598
  • 350954 03/08/11 02:30 a-onl Start_Run_2598,Run_type=Other,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB front chamber scan 1650,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350953 03/08/11 02:29 a-onl End_of_Run_2597
  • 350952 03/08/11 02:24 a-onl Start_Run_2597,Run_type=Other,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB front chamber scan 1550,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350951 03/08/11 02:23 a-onl End_of_Run_2596
  • 350950 03/08/11 02:18 a-onl Start_Run_2596,Run_type=Other,target_type=Slanted Carbon,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=BB front chamber scan,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350949 03/08/11 02:12 danez Move to Slanted Carbon
  • 350948 03/08/11 02:11 a-onl End_of_Run_2595
  • 350947 03/08/11 02:09 danez Spot++ Screenshot
  • 350946 03/08/11 02:07 a-onl Start_Run_2595,Run_type=Spot++,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=spot++, BeO target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350945 03/08/11 01:51 danez Target Screenshots
  • 350944 03/08/11 01:42 lselvy hv map change
  • 350943 03/08/11 01:31 auto_snap Radiation and Ozone snapshot
  • 350942 03/08/11 01:31 auto_snap Miscellaneous snapshot
  • 350941 03/08/11 01:31 auto_snap Hall A Tools snapshot
  • 350940 03/08/11 01:31 auto_snap Compton Polarimeter snapshot
  • 350939 03/08/11 01:31 auto_snap BigBox Magnet snapshot
  • 350938 03/08/11 01:31 auto_snap Beam Position and Current snapshot
  • 350937 03/08/11 01:03 danez Vacuum Pressure above 2^-4 torr
  • 350936 03/08/11 00:55 danez Moved to BeO
  • 350935 03/08/11 00:42 orchen neutron detection efficiency after analyser changes
  • 350934 03/08/11 00:01 adaq Beam Time Accounting
  • Monday

  • 350933 03/07/11 23:57 flay Monday March 7 Swing Shift Summary
  • 350932 03/07/11 23:51 a-onl End_of_Run_2594
  • 350931 03/07/11 23:51 orchen Analyzer change in HAND lib
  • 350930 03/07/11 23:32 a-onl Start_Run_2594,Run_type=Calibration,target_type=LD2 4 cm cell,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=deutron, X<1 500MeV/c HAND efficiency calibration, EDTM disabled,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350929 03/07/11 23:31 a-onl End_of_Run_2593
  • 350928 03/07/11 23:30 Punjabi PM scalers for run 2593
  • 350927 03/07/11 23:29 Punjabi HRS scalers for run 2593
  • 350926 03/07/11 23:17 flay Hall A Tools
  • 350925 03/07/11 23:03 Punjabi coda problems
  • 350924 03/07/11 22:59 a-onl Start_Run_2593,Run_type=Calibration,target_type=LD2 4 cm cell,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=deutron, X<1 500MeV/c HAND efficiency calibration, EDTM disabled,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350923 03/07/11 22:35 a-onl End_of_Run_2592
  • 350922 03/07/11 22:34 Punjabi PM scalers
  • 350921 03/07/11 22:33 Punjabi Hrs scalers
  • 350920 03/07/11 22:08 a-onl Start_Run_2592,Run_type=Calibration,target_type=LD2 4 cm cell,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=deutron, X<1 500MeV/c HAND efficiency calibration, EDTM disabled,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350919 03/07/11 21:58 a-onl End_of_Run_2591
  • 350918 03/07/11 21:52 Punjabi PM scalers
  • 350917 03/07/11 21:49 flay Hall A Tools
  • 350916 03/07/11 21:31 a-onl Start_Run_2591,Run_type=Calibration,target_type=LD2 4 cm cell,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=deutron, X<1 500MeV/c HAND efficiency calibration, EDTM disabled,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350915 03/07/11 21:29 a-onl End_of_Run_2590
  • 350914 03/07/11 21:27 Punjabi PM scalers
  • 350913 03/07/11 21:27 Punjabi HRS scalers
  • 350912 03/07/11 21:09 a-onl Start_Run_2590,Run_type=Calibration,target_type=LD2 4 cm cell,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=deutron, X<1 500MeV/c HAND efficiency calibration, EDTM disabled,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350911 03/07/11 21:07 a-onl End_of_Run_2589
  • 350910 03/07/11 21:07 Punjabi PM scalers
  • 350909 03/07/11 21:06 Punjabi HRS scalers
  • 350908 03/07/11 20:47 a-onl Start_Run_2589,Run_type=Calibration,target_type=LD2 4 cm cell,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=deutron, X<1 500MeV/c HAND efficiency calibration, EDTM disabled,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350907 03/07/11 20:47 a-onl End_of_Run_2588
  • 350906 03/07/11 20:44 Punjabi Pmscalers
  • 350905 03/07/11 20:43 Punjabi HRS scalers
  • 350904 03/07/11 20:24 a-onl Start_Run_2588,Run_type=Calibration,target_type=LD2 4 cm cell,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=deutron, X<1 500MeV/c HAND efficiency calibration, EDTM disabled,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350903 03/07/11 20:22 a-onl End_of_Run_2587
  • 350902 03/07/11 20:22 Punjabi PM scalers
  • 350901 03/07/11 20:21 Punjabi HRS scalers
  • 350900 03/07/11 20:17 V. Sulkosky Checked HAND Thresholds Okay (-30 mV)
  • 350899 03/07/11 20:16 orchen first results from 5 and 10 uA deutron efficiency runs
  • 350898 03/07/11 20:00 a-onl Start_Run_2587,Run_type=Calibration,target_type=LD2 4 cm cell,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=deutron, X<1 500MeV/c HAND efficiency calibration, EDTM disabled,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350897 03/07/11 19:58 a-onl End_of_Run_2586
  • 350896 03/07/11 19:58 Punjabi HRS scalers
  • 350895 03/07/11 19:57 Punjabi PM scalers
  • 350894 03/07/11 19:48 a-onl Start_Run_2580,Run_type=Production SRC,target_type=LD2 4 cm cell,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=deutron, X<1 500MeV/c HAND efficiency calibration,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350893 03/07/11 19:36 a-onl Start_Run_2586,Run_type=Calibration,target_type=LD2 4 cm cell,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=deutron, X<1 500MeV/c HAND efficiency calibration, EDTM disabled,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350892 03/07/11 19:34 a-onl End_of_Run_2585
  • 350891 03/07/11 19:33 Punjabi Hrs scalers
  • 350890 03/07/11 19:32 Punjabi scalers
  • 350889 03/07/11 19:17 Reimer Target: Loop 2 hal2_cernox_2_r
  • 350888 03/07/11 19:11 flay Run 2585: Move back to 5 uA
  • 350887 03/07/11 18:59 a-onl Start_Run_2585,Run_type=Calibration,target_type=LD2 4 cm cell,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=deutron, X<1 500MeV/c HAND efficiency calibration, EDTM disabled,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350886 03/07/11 18:56 a-onl End_of_Run_2584
  • 350885 03/07/11 18:51 V. Sulkosky Updated BCM Coefficients in EOR Script
  • 350884 03/07/11 18:44 flay Hall A Tools
  • 350883 03/07/11 18:41 igorko updated current value
  • 350882 03/07/11 18:40 igorko Beam current
  • 350881 03/07/11 18:38 a-onl Start_Run_2584,Run_type=Calibration,target_type=LD2 4 cm cell,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=deutron, X<1 500MeV/c HAND efficiency calibration, EDTM disabled,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350880 03/07/11 18:37 a-onl End_of_Run_2583
  • 350879 03/07/11 18:33 a-onl Start_Run_2583,Run_type=Calibration,target_type=LD2 4 cm cell,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=deutron, X<1 500MeV/c HAND efficiency calibration, EDTM disabled,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350878 03/07/11 18:32 a-onl End_of_Run_2582
  • 350877 03/07/11 18:19 V. Sulkosky Disabled EDTM
  • 350876 03/07/11 18:17 a-onl Start_Run_2582,Run_type=Calibration,target_type=LD2 4 cm cell,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=deutron, X<1 500MeV/c HAND efficiency calibration, EDTM disabled,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350875 03/07/11 18:13 a-onl End_of_Run_2581
  • 350874 03/07/11 18:11 V. Sulkosky Updated Calibration Constants BCM
  • 350873 03/07/11 18:01 a-onl Start_Run_2581,Run_type=Calibration,target_type=LD2 4 cm cell,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=deutron, X<1 500MeV/c HAND efficiency calibration,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350872 03/07/11 17:59 a-onl End_of_Run_2580
  • 350871 03/07/11 17:50 V. Sulkosky Coda Configuration Updated
  • 350870 03/07/11 17:44 a-onl End_of_Run_2579
  • 350869 03/07/11 17:43 a-onl Start_Run_2579,Run_type=Check out,target_type=LD2 4 cm cell,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=test run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350868 03/07/11 17:32 a-onl End_of_Run_2578
  • 350867 03/07/11 17:31 auto_snap Radiation and Ozone snapshot
  • 350866 03/07/11 17:31 auto_snap Miscellaneous snapshot
  • 350865 03/07/11 17:31 auto_snap Hall A Tools snapshot
  • 350864 03/07/11 17:31 auto_snap Compton Polarimeter snapshot
  • 350863 03/07/11 17:31 auto_snap BigBox Magnet snapshot
  • 350862 03/07/11 17:31 auto_snap Beam Position and Current snapshot
  • 350861 03/07/11 17:30 a-onl Start_Run_2578,Run_type=Calibration,target_type=LD2 4 cm cell,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=4 cm Deuterium target, x<1. 500 MeV/c,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350860 03/07/11 17:24 a-onl End_of_Run_2577
  • 350859 03/07/11 17:22 Punjabi Spot++
  • 350858 03/07/11 17:14 a-onl Start_Run_2577,Run_type=Spot++,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.7642,leftarm_theta=54.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=14.,rightarm_sieve=out,comment_text=Spot++. Beo target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350857 03/07/11 17:03 a-onl End_of_Run_2576
  • 350856 03/07/11 16:38 jpchen Don't remove the Screens directory.
  • 350855 03/07/11 16:34 jpchen Access to hall to reset Loop 1 fan controller
  • 350854 03/07/11 16:31 V. Sulkosky Re: SRC configuration Unbuffered
  • 350853 03/07/11 16:24 V. Sulkosky Fixed T7 Trigger (escorted access)
  • 350852 03/07/11 16:17 a-onl Start_Run_2576,Run_type=Cosmics,target_type=4cm dummy,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.621,leftarm_theta=34.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=21.,rightarm_sieve=out,comment_text=cosmics,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350851 03/07/11 16:00 adaq Beam Time Accounting
  • 350850 03/07/11 15:48 a-onl Start_Run_2575,Run_type=Cosmics,target_type=4cm dummy,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.621,leftarm_theta=34.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=21.,rightarm_sieve=out,comment_text=cosmics,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350849 03/07/11 15:47 averett Day Shift Summary
  • 350848 03/07/11 15:44 a-onl Start_Run_2574,Run_type=Cosmics,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.621,leftarm_theta=34.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=21.,rightarm_sieve=out,comment_text=cosmics,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350847 03/07/11 15:26 orchen BB sieve out
  • 350846 03/07/11 15:09 B Anderson Beam off JT closed 1 step
  • 350845 03/07/11 15:02 a-onl End_of_Run_2573
  • 350844 03/07/11 14:51 a-onl Start_Run_2573,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.621,leftarm_theta=34.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=21.,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350843 03/07/11 14:51 a-onl End_of_Run_2572
  • 350842 03/07/11 14:39 a-onl Start_Run_2572,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.621,leftarm_theta=34.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=21.,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350841 03/07/11 14:38 a-onl End_of_Run_2571
  • 350840 03/07/11 14:33 B Anderson Target charts
  • 350839 03/07/11 14:32 B Anderson Target screen
  • 350838 03/07/11 14:28 a-onl Start_Run_2571,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.621,leftarm_theta=34.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=21.,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350837 03/07/11 14:27 a-onl End_of_Run_2570
  • 350836 03/07/11 14:27 jmammei call for escorted access at 11:50 pm
  • 350835 03/07/11 14:15 a-onl Start_Run_2570,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.621,leftarm_theta=34.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=21.,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350834 03/07/11 14:15 jmammei shifts canceled
  • 350833 03/07/11 14:14 a-onl End_of_Run_2569
  • 350832 03/07/11 14:11 averett Hall A Tools
  • 350831 03/07/11 14:05 jmammei run plan for overnight Monday and beyond...
  • 350830 03/07/11 14:04 a-onl Start_Run_2569,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.621,leftarm_theta=34.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=21.,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350829 03/07/11 14:03 a-onl End_of_Run_2568
  • 350828 03/07/11 13:52 a-onl Start_Run_2568,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.621,leftarm_theta=34.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=21.,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350827 03/07/11 13:51 a-onl End_of_Run_2567
  • 350826 03/07/11 13:40 danez Kinematic #13 N-Delta plot
  • 350825 03/07/11 13:39 a-onl Start_Run_2567,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.621,leftarm_theta=34.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=21.,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350824 03/07/11 13:38 a-onl End_of_Run_2566
  • 350823 03/07/11 13:38 danez Kinematic #14 N-Delta plot
  • 350822 03/07/11 13:27 B Anderson Target Screen
  • 350821 03/07/11 13:26 a-onl Start_Run_2566,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.621,leftarm_theta=34.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=21.,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350820 03/07/11 13:25 a-onl End_of_Run_2565
  • 350819 03/07/11 13:20 lselvy SRC configuration
  • 350818 03/07/11 13:20 B Anderson H target Open JT 1 step
  • 350817 03/07/11 13:13 a-onl Start_Run_2565,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.621,leftarm_theta=34.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=21.,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350816 03/07/11 13:04 a-onl End_of_Run_2564
  • 350815 03/07/11 12:49 a-onl Start_Run_2564,Run_type=Other,target_type=15cm dummy,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.621,leftarm_theta=34.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=21.,rightarm_sieve=out,comment_text=junk,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350814 03/07/11 12:47 a-onl End_of_Run_2563
  • 350813 03/07/11 12:46 a-onl Start_Run_2563,Run_type=Other,target_type=15cm dummy,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.621,leftarm_theta=34.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=21.,rightarm_sieve=out,comment_text=junk,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350812 03/07/11 12:45 a-onl End_of_Run_2562
  • 350811 03/07/11 12:39 a-onl Start_Run_2562,Run_type=Optics,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.621,leftarm_theta=34.5,leftarm_sieve=out,rightarm_p=0.789,rightarm_theta=21.,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350810 03/07/11 12:15 danez Spectrometer Momenta Changed
  • 350809 03/07/11 12:14 danez Right Bogies Controls Acting Up
  • 350808 03/07/11 12:12 danez Spectrometers Moved (LHRS = 34.5, RHRS = 21)
  • 350807 03/07/11 12:08 B Anderson BeO. Close JT 2 steps.
  • 350806 03/07/11 12:00 a-onl End_of_Run_2561
  • 350805 03/07/11 11:46 a-onl Start_Run_2561,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.574,leftarm_theta=37.5,leftarm_sieve=out,rightarm_p=0.826,rightarm_theta=20.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350804 03/07/11 11:46 a-onl End_of_Run_2560
  • 350803 03/07/11 11:34 a-onl Start_Run_2560,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.574,leftarm_theta=37.5,leftarm_sieve=out,rightarm_p=0.826,rightarm_theta=20.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350802 03/07/11 11:33 a-onl End_of_Run_2559
  • 350801 03/07/11 11:19 a-onl Start_Run_2559,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.574,leftarm_theta=37.5,leftarm_sieve=out,rightarm_p=0.826,rightarm_theta=20.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350800 03/07/11 11:18 a-onl End_of_Run_2558
  • 350799 03/07/11 11:07 a-onl Start_Run_2558,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.574,leftarm_theta=37.5,leftarm_sieve=out,rightarm_p=0.826,rightarm_theta=20.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350798 03/07/11 11:06 a-onl End_of_Run_2557
  • 350797 03/07/11 10:54 a-onl Start_Run_2557,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.574,leftarm_theta=37.5,leftarm_sieve=out,rightarm_p=0.826,rightarm_theta=20.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350796 03/07/11 10:54 a-onl End_of_Run_2556
  • 350795 03/07/11 10:51 orchen wire chamber efficiency results from HV & treshold scans (RESULTS ARE NOT CLEAR!!)
  • 350794 03/07/11 10:42 a-onl Start_Run_2556,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.574,leftarm_theta=37.5,leftarm_sieve=out,rightarm_p=0.826,rightarm_theta=20.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350793 03/07/11 10:41 a-onl End_of_Run_2555
  • 350792 03/07/11 10:32 B Anderson Target Plots
  • 350791 03/07/11 10:30 a-onl Start_Run_2555,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.574,leftarm_theta=37.5,leftarm_sieve=out,rightarm_p=0.826,rightarm_theta=20.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350790 03/07/11 10:29 a-onl End_of_Run_2554
  • 350789 03/07/11 10:16 a-onl Start_Run_2554,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.574,leftarm_theta=37.5,leftarm_sieve=out,rightarm_p=0.826,rightarm_theta=20.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350788 03/07/11 10:16 a-onl End_of_Run_2553
  • 350787 03/07/11 10:05 B Anderson Target screen.
  • 350786 03/07/11 10:03 a-onl Start_Run_2553,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.574,leftarm_theta=37.5,leftarm_sieve=out,rightarm_p=0.826,rightarm_theta=20.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350785 03/07/11 10:03 a-onl End_of_Run_2552
  • 350784 03/07/11 09:51 a-onl Start_Run_2552,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.574,leftarm_theta=37.5,leftarm_sieve=out,rightarm_p=0.826,rightarm_theta=20.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350783 03/07/11 09:50 a-onl End_of_Run_2551
  • 350782 03/07/11 09:49 A. Camsonne Run 2500 sync check - higher statistics
  • 350781 03/07/11 09:39 a-onl Start_Run_2551,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.574,leftarm_theta=37.5,leftarm_sieve=out,rightarm_p=0.826,rightarm_theta=20.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350780 03/07/11 09:35 B Anderson JT valve open 2 steps
  • 350779 03/07/11 09:31 auto_snap Radiation and Ozone snapshot
  • 350778 03/07/11 09:31 auto_snap Miscellaneous snapshot
  • 350777 03/07/11 09:31 auto_snap Hall A Tools snapshot
  • 350776 03/07/11 09:31 auto_snap Compton Polarimeter snapshot
  • 350775 03/07/11 09:31 auto_snap BigBox Magnet snapshot
  • 350774 03/07/11 09:30 auto_snap Beam Position and Current snapshot
  • 350773 03/07/11 09:25 a-onl End_of_Run_2550
  • 350772 03/07/11 09:19 A. Camsonne /adaql10/work1 disk full removing older files
  • 350771 03/07/11 09:13 jmammei RC Daily Summary
  • 350770 03/07/11 09:10 a-onl Start_Run_2550,Run_type=Other,target_type=15cm dummy,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.574,leftarm_theta=37.5,leftarm_sieve=out,rightarm_p=0.826,rightarm_theta=20.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350769 03/07/11 09:05 a-onl End_of_Run_2549
  • 350768 03/07/11 09:02 a-onl Start_Run_2549,Run_type=Optics,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.574,leftarm_theta=37.5,leftarm_sieve=out,rightarm_p=0.826,rightarm_theta=20.5,rightarm_sieve=out,comment_text=Spot,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350767 03/07/11 09:01 danez Spot++ Screenshot
  • 350766 03/07/11 08:58 a-onl End_of_Run_2548
  • 350765 03/07/11 08:58 a-onl Start_Run_2548,Run_type=Spot++,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.574,leftarm_theta=37.5,leftarm_sieve=out,rightarm_p=0.826,rightarm_theta=20.5,rightarm_sieve=out,comment_text=Spot,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350764 03/07/11 08:56 a-onl End_of_Run_2547
  • 350763 03/07/11 08:55 a-onl Start_Run_2547,Run_type=Spot++,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.574,leftarm_theta=37.5,leftarm_sieve=out,rightarm_p=0.826,rightarm_theta=20.5,rightarm_sieve=out,comment_text=Spot,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350762 03/07/11 08:54 orchen two bugs in the start of run scripts
  • 350761 03/07/11 08:52 orchen loading HV file for BB and HAND
  • 350760 03/07/11 08:46 B Anderson JT Valve opened. BeO in.
  • 350759 03/07/11 08:42 eip radiation level near the scattering chamber
  • 350758 03/07/11 08:01 adaq Beam Time Accounting
  • 350757 03/07/11 08:00 riordan Owl Shift Summary
  • 350756 03/07/11 07:58 M Friend JT valve stepped closed
  • 350755 03/07/11 07:55 a-onl End_of_Run_2546
  • 350754 03/07/11 07:42 a-onl Start_Run_2546,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350753 03/07/11 07:41 a-onl End_of_Run_2545
  • 350752 03/07/11 07:27 a-onl Start_Run_2545,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350751 03/07/11 07:26 a-onl End_of_Run_2544
  • 350750 03/07/11 07:12 a-onl Start_Run_2544,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350749 03/07/11 07:11 a-onl End_of_Run_2543
  • 350748 03/07/11 07:08 danez Runs 2541 and 2452 are junk
  • 350747 03/07/11 06:59 a-onl Start_Run_2543,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350746 03/07/11 06:50 a-onl End_of_Run_2540
  • 350745 03/07/11 06:45 M Friend Target screenshots
  • 350744 03/07/11 06:40 a-onl Start_Run_2540,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350743 03/07/11 06:39 a-onl End_of_Run_2539
  • 350742 03/07/11 06:27 a-onl Start_Run_2539,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350741 03/07/11 06:26 a-onl End_of_Run_2538
  • 350740 03/07/11 06:11 a-onl Start_Run_2538,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350739 03/07/11 06:10 a-onl End_of_Run_2537
  • 350738 03/07/11 06:00 a-onl Start_Run_2537,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350737 03/07/11 05:59 a-onl End_of_Run_2536
  • 350736 03/07/11 05:48 a-onl Start_Run_2536,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350735 03/07/11 05:47 a-onl End_of_Run_2535
  • 350734 03/07/11 05:36 a-onl Start_Run_2535,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350733 03/07/11 05:36 a-onl End_of_Run_2534
  • 350732 03/07/11 05:25 a-onl Start_Run_2534,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350731 03/07/11 05:22 M Friend Target to 15cm LH2
  • 350730 03/07/11 05:18 a-onl End_of_Run_2533
  • 350729 03/07/11 04:58 a-onl Start_Run_2533,Run_type=Other,target_type=15cm dummy,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=15cm Aluminum Dummy Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350728 03/07/11 04:55 M Friend Target to 15cm Al dummy
  • 350727 03/07/11 04:55 M Friend Target high powered heater yellow alarm
  • 350726 03/07/11 04:53 a-onl End_of_Run_2532
  • 350725 03/07/11 04:48 a-onl Start_Run_2532,Run_type=Optics,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=BeO Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350724 03/07/11 04:45 M Friend Target to BeO
  • 350723 03/07/11 04:39 a-onl End_of_Run_2531
  • 350722 03/07/11 04:28 a-onl Start_Run_2531,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350721 03/07/11 04:27 a-onl End_of_Run_2530
  • 350720 03/07/11 04:22 M Friend Target snapshots
  • 350719 03/07/11 04:17 a-onl Start_Run_2530,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350718 03/07/11 04:16 a-onl End_of_Run_2529
  • 350717 03/07/11 04:13 riordan Hall A Tools
  • 350716 03/07/11 04:12 riordan spot++
  • 350715 03/07/11 04:05 a-onl Start_Run_2529,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350714 03/07/11 04:03 a-onl End_of_Run_2528
  • 350713 03/07/11 03:51 a-onl Start_Run_2528,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350712 03/07/11 03:50 a-onl End_of_Run_2527
  • 350711 03/07/11 03:42 danez Synch Check Plot for Run 2480
  • 350710 03/07/11 03:39 a-onl Start_Run_2527,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350709 03/07/11 03:39 a-onl End_of_Run_2526
  • 350708 03/07/11 03:28 a-onl Start_Run_2526,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350707 03/07/11 03:27 a-onl End_of_Run_2525
  • 350706 03/07/11 03:14 a-onl Start_Run_2525,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350705 03/07/11 03:14 a-onl End_of_Run_2524
  • 350704 03/07/11 02:51 a-onl Start_Run_2524,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350703 03/07/11 02:50 a-onl End_of_Run_2523
  • 350702 03/07/11 02:39 M Friend Target screenshots
  • 350701 03/07/11 02:35 M Friend JT valve stepped open
  • 350700 03/07/11 02:29 a-onl Start_Run_2523,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production - test run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350699 03/07/11 02:29 a-onl End_of_Run_2522
  • 350698 03/07/11 02:28 a-onl Start_Run_2522,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.646,leftarm_theta=41.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production - test run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350697 03/07/11 02:22 riordan LHRS at 41.5 degrees
  • 350696 03/07/11 02:07 M Friend Target snapshots
  • 350695 03/07/11 02:00 a-onl End_of_Run_2521
  • 350694 03/07/11 02:00 riordan Finish Setting #9 - changing to setting #10
  • 350693 03/07/11 01:49 a-onl Start_Run_2521,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350692 03/07/11 01:48 a-onl End_of_Run_2520
  • 350691 03/07/11 01:37 a-onl Start_Run_2520,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350690 03/07/11 01:37 a-onl End_of_Run_2519
  • 350689 03/07/11 01:36 danez Synch Roc Screenshot - Run 2452
  • 350688 03/07/11 01:35 nmuangma BCM cavity constants
  • 350687 03/07/11 01:30 auto_snap Radiation and Ozone snapshot
  • 350686 03/07/11 01:30 auto_snap Miscellaneous snapshot
  • 350685 03/07/11 01:30 auto_snap Hall A Tools snapshot
  • 350684 03/07/11 01:30 auto_snap Compton Polarimeter snapshot
  • 350683 03/07/11 01:30 auto_snap BigBox Magnet snapshot
  • 350682 03/07/11 01:30 auto_snap Beam Position and Current snapshot
  • 350681 03/07/11 01:26 a-onl Start_Run_2519,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350680 03/07/11 01:25 a-onl End_of_Run_2518
  • 350679 03/07/11 01:14 a-onl Start_Run_2518,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350678 03/07/11 01:10 a-onl End_of_Run_2517
  • 350677 03/07/11 01:00 a-onl Start_Run_2517,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350676 03/07/11 00:59 a-onl End_of_Run_2516
  • 350675 03/07/11 00:49 a-onl Start_Run_2516,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350674 03/07/11 00:48 a-onl End_of_Run_2515
  • 350673 03/07/11 00:37 a-onl Start_Run_2515,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350672 03/07/11 00:34 a-onl End_of_Run_2514
  • 350671 03/07/11 00:24 a-onl Start_Run_2514,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350670 03/07/11 00:23 a-onl End_of_Run_2513
  • 350669 03/07/11 00:12 a-onl Start_Run_2513,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350668 03/07/11 00:11 a-onl End_of_Run_2512
  • 350667 03/07/11 00:05 adaq Beam Time Accounting
  • 350666 03/07/11 00:03 zwzhao swing shift summary
  • 350665 03/07/11 00:00 a-onl Start_Run_2512,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • Sunday

  • 350664 03/06/11 23:59 a-onl End_of_Run_2511
  • 350663 03/06/11 23:51 zwzhao HallA tools snapshot
  • 350662 03/06/11 23:47 a-onl Start_Run_2511,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350661 03/06/11 23:46 a-onl End_of_Run_2510
  • 350660 03/06/11 23:39 LeRose Target Screens
  • 350659 03/06/11 23:38 LeRose Target 15cm LH2
  • 350658 03/06/11 23:38 flay RMS values are not reducing, staying at ~ 50.
  • 350657 03/06/11 23:30 a-onl Start_Run_2510,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350656 03/06/11 23:29 a-onl End_of_Run_2509
  • 350655 03/06/11 23:18 a-onl Start_Run_2509,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350654 03/06/11 23:18 a-onl End_of_Run_2508
  • 350653 03/06/11 23:11 flay run 2508: ROC sync differences RMS are back down to acceptable levels
  • 350652 03/06/11 23:04 a-onl Start_Run_2508,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350651 03/06/11 23:03 a-onl End_of_Run_2507
  • 350650 03/06/11 22:52 a-onl Start_Run_2507,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350649 03/06/11 22:42 a-onl End_of_Run_2506
  • 350648 03/06/11 22:37 a-onl Start_Run_2506,Run_type=Optics,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=BeO Optics,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350647 03/06/11 22:35 LeRose Target BeO
  • 350646 03/06/11 22:31 a-onl End_of_Run_2505
  • 350645 03/06/11 22:17 a-onl Start_Run_2505,Run_type=Optics,target_type=15cm dummy,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Optics,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350644 03/06/11 22:15 LeRose Target to 15 cm Dummy
  • 350643 03/06/11 22:09 flay Run 2504 is Junk
  • 350642 03/06/11 22:08 a-onl End_of_Run_2504
  • 350641 03/06/11 22:08 a-onl Start_Run_2504,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350640 03/06/11 22:07 a-onl End_of_Run_2503
  • 350639 03/06/11 21:55 a-onl Start_Run_2503,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350638 03/06/11 21:54 a-onl End_of_Run_2502
  • 350637 03/06/11 21:54 A. Camsonne Set sync plots to log scale - Sync differences fluctuations.
  • 350636 03/06/11 21:43 a-onl Start_Run_2502,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350635 03/06/11 21:42 a-onl End_of_Run_2501
  • 350634 03/06/11 21:36 flay Run 2500: Sync ROC Differences Rise Again
  • 350633 03/06/11 21:31 a-onl Start_Run_2501,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350632 03/06/11 21:31 a-onl End_of_Run_2500
  • 350631 03/06/11 21:20 a-onl Start_Run_2500,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350630 03/06/11 21:19 a-onl End_of_Run_2499
  • 350629 03/06/11 21:18 flay Run 2499: ROC RMS Plots Are Calming Down
  • 350628 03/06/11 21:07 a-onl Start_Run_2499,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350627 03/06/11 21:06 a-onl End_of_Run_2498
  • 350626 03/06/11 20:54 a-onl Start_Run_2498,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350625 03/06/11 20:54 a-onl End_of_Run_2497
  • 350624 03/06/11 20:42 a-onl Start_Run_2497,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350623 03/06/11 20:41 a-onl End_of_Run_2496
  • 350622 03/06/11 20:32 flay Run 2495: Sync ROC Plots Back to Normal (Mostly)
  • 350621 03/06/11 20:32 jmammei doing 85% of the charge at setting #9
  • 350620 03/06/11 20:30 a-onl Start_Run_2496,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350619 03/06/11 20:29 a-onl End_of_Run_2495
  • 350618 03/06/11 20:19 flay Run 2493: Sync rocs are EMTPY!
  • 350617 03/06/11 20:18 a-onl Start_Run_2495,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350616 03/06/11 20:18 a-onl End_of_Run_2494
  • 350615 03/06/11 20:17 flay sync roc differences are still showing an RMS > 10...
  • 350614 03/06/11 20:07 a-onl Start_Run_2494,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350613 03/06/11 20:06 a-onl End_of_Run_2493
  • 350612 03/06/11 19:55 LeRose Target Screens
  • 350611 03/06/11 19:53 a-onl Start_Run_2493,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350610 03/06/11 19:52 a-onl End_of_Run_2492
  • 350609 03/06/11 19:51 flay Sync ROC Plots for Run 2491
  • 350608 03/06/11 19:46 orchen Updated Results for BB HV and Treshold Scan.
  • 350607 03/06/11 19:35 a-onl Start_Run_2492,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350606 03/06/11 19:34 a-onl End_of_Run_2491
  • 350605 03/06/11 19:24 flay Run 2490 is Junk
  • 350604 03/06/11 19:23 a-onl Start_Run_2491,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350603 03/06/11 19:22 a-onl End_of_Run_2490
  • 350602 03/06/11 19:22 a-onl Start_Run_2490,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350601 03/06/11 19:21 a-onl End_of_Run_2489
  • 350600 03/06/11 19:11 flay N->Delta plots for I = 25 uA on 15cm LH2 target
  • 350599 03/06/11 19:05 flay Run 2488 is at I = 30 uA
  • 350598 03/06/11 19:04 a-onl Start_Run_2489,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=test,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350597 03/06/11 19:03 a-onl End_of_Run_2488
  • 350596 03/06/11 19:01 flay N->Delta plots for I = 30 uA on 15cm LH2 target
  • 350595 03/06/11 18:58 flay N->Delta plots for I = 40 uA on 15cm LH2 target
  • 350594 03/06/11 18:55 a-onl Start_Run_2488,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=test,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350593 03/06/11 18:54 a-onl End_of_Run_2487
  • 350592 03/06/11 18:52 a-onl Start_Run_2487,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.646,leftarm_theta=21.0,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=test,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350591 03/06/11 18:40 jmammei techs: LHRS now moving
  • 350590 03/06/11 18:34 jmammei techs: LHRS won't move
  • 350589 03/06/11 18:20 a-onl End_of_Run_2486
  • 350588 03/06/11 18:17 igorko BigBite databases
  • 350587 03/06/11 17:59 a-onl Start_Run_2486,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.670,leftarm_theta=31.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350586 03/06/11 17:58 a-onl End_of_Run_2485
  • 350585 03/06/11 17:51 jmammei Run plan for overnight Sunday / Monday day
  • 350584 03/06/11 17:32 a-onl Start_Run_2485,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.670,leftarm_theta=31.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350583 03/06/11 17:32 a-onl End_of_Run_2484
  • 350582 03/06/11 17:30 auto_snap Radiation and Ozone snapshot
  • 350581 03/06/11 17:30 auto_snap Miscellaneous snapshot
  • 350580 03/06/11 17:30 auto_snap Hall A Tools snapshot
  • 350579 03/06/11 17:30 auto_snap Compton Polarimeter snapshot
  • 350578 03/06/11 17:30 auto_snap BigBox Magnet snapshot
  • 350577 03/06/11 17:30 auto_snap Beam Position and Current snapshot
  • 350576 03/06/11 17:07 V. Sulkosky Re: HAND Timing Measurements (Late Entry 2/24/2011)
  • 350575 03/06/11 17:04 flay Run 2483 is Junk
  • 350574 03/06/11 17:03 a-onl Start_Run_2484,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.670,leftarm_theta=31.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350573 03/06/11 16:59 a-onl End_of_Run_2482
  • 350572 03/06/11 16:42 LeRose Target Screens
  • 350571 03/06/11 16:35 a-onl Start_Run_2482,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.670,leftarm_theta=31.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350570 03/06/11 16:35 a-onl End_of_Run_2481
  • 350569 03/06/11 16:16 orchen BB tresholds
  • 350568 03/06/11 16:12 orchen HAND tresholds
  • 350567 03/06/11 16:12 orchen HAND tresholds
  • 350566 03/06/11 16:12 adaq Beam Time Accounting
  • 350565 03/06/11 16:11 M Defurne Day Shift Summary
  • 350564 03/06/11 16:10 a-onl Start_Run_2481,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.670,leftarm_theta=31.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350563 03/06/11 16:09 a-onl End_of_Run_2480
  • 350562 03/06/11 15:59 adaq screen shot of ndelta for run2480
  • 350561 03/06/11 15:47 M Defurne Start production run with LH2 4cm at 15:45
  • 350560 03/06/11 15:45 a-onl Start_Run_2480,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.670,leftarm_theta=31.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350559 03/06/11 15:38 adaq spot++ from run2479
  • 350558 03/06/11 15:31 a-onl End_of_Run_2479
  • 350557 03/06/11 15:30 A. Camsonne Finished booting the iocha22
  • 350556 03/06/11 15:25 a-onl Start_Run_2479,Run_type=Optics,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.670,leftarm_theta=31.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=5mim Optics run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350555 03/06/11 15:23 nmuangma Correction: BCM current calibration
  • 350554 03/06/11 15:17 a-onl End_of_Run_2478
  • 350553 03/06/11 15:16 orchen ROC-8 restart
  • 350552 03/06/11 15:16 M Defurne screenshots of Hall A general Tools
  • 350551 03/06/11 15:14 V. Sulkosky Re: can't connect to bbps1.2001 (HAND right side HV control)
  • 350550 03/06/11 15:04 orchen can't connect to bbps1.2001 (HAND right side HV control)
  • 350549 03/06/11 15:02 orchen HV change in veto
  • 350548 03/06/11 14:53 M Defurne run 2473 Al 15cm Ok
  • 350547 03/06/11 14:49 adaq coda works again
  • 350546 03/06/11 14:49 a-onl Start_Run_2478,Run_type=Calibration,target_type=4cm dummy,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.670,leftarm_theta=31.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Al 1/2hour calibration run with 20uA Beam,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350545 03/06/11 14:49 A. Camsonne iocha22 error during boot
  • 350544 03/06/11 14:45 M Defurne white alarm LeCroy_HV because of iocha22
  • 350543 03/06/11 14:43 A. Camsonne iocha22 Transfer Timed Out, might have network issues
  • 350542 03/06/11 14:30 adaq the error of coda
  • 350541 03/06/11 14:27 adaq er2 of coda is pused
  • 350540 03/06/11 14:18 adaq run#2474 is junk becuase EB2 of coda does not response
  • 350539 03/06/11 14:14 M Defurne Late Entry for target changed
  • 350538 03/06/11 14:03 a-onl End_of_Run_2473
  • 350537 03/06/11 13:51 M Defurne LeCroy HV screen
  • 350536 03/06/11 13:36 M Defurne start run calibration with alumimium 15cm
  • 350535 03/06/11 13:25 M Defurne white alarm hall a LECROY-HV
  • 350534 03/06/11 13:12 a-onl End_of_Run_2472
  • 350533 03/06/11 13:01 a-onl Start_Run_2472,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.670,leftarm_theta=31.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350532 03/06/11 12:59 a-onl End_of_Run_2471
  • 350531 03/06/11 12:49 a-onl Start_Run_2471,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.670,leftarm_theta=31.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350530 03/06/11 12:47 a-onl End_of_Run_2470
  • 350529 03/06/11 12:37 a-onl Start_Run_2470,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.670,leftarm_theta=31.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350528 03/06/11 12:36 a-onl End_of_Run_2469
  • 350527 03/06/11 12:26 a-onl Start_Run_2469,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.670,leftarm_theta=31.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350526 03/06/11 12:22 a-onl End_of_Run_2468
  • 350525 03/06/11 12:13 a-onl Start_Run_2468,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.670,leftarm_theta=31.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350524 03/06/11 12:12 a-onl End_of_Run_2467
  • 350523 03/06/11 12:06 a-onl Start_Run_2467,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.670,leftarm_theta=31.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350522 03/06/11 12:05 a-onl End_of_Run_2466
  • 350521 03/06/11 12:02 adaq run#2466 is production run
  • 350520 03/06/11 11:59 a-onl Start_Run_2466,Run_type=Other,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=35,leftarm_p=0.670,leftarm_theta=31.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350519 03/06/11 11:55 a-onl End_of_Run_2465
  • 350518 03/06/11 11:55 danez Run 2463 - Incorrect Angle Data
  • 350517 03/06/11 11:50 adaq run#2464 is junk since TS1 does not response.
  • 350516 03/06/11 11:47 a-onl Start_Run_2465,Run_type=Other,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.670,leftarm_theta=31.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350515 03/06/11 11:36 a-onl End_of_Run_2463
  • 350514 03/06/11 11:34 M Defurne Finish setting #12 moving to setting #8
  • 350513 03/06/11 11:31 a-onl Start_Run_2463,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.670,leftarm_theta=31.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350512 03/06/11 11:27 a-onl End_of_Run_2462
  • 350511 03/06/11 11:11 a-onl Start_Run_2462,Run_type=Other,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.670,leftarm_theta=31.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=BEAM current test run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350510 03/06/11 11:11 M Defurne open JT valve for LH2 15cm
  • 350509 03/06/11 11:05 M Defurne change target to LH2 15cm at 11:00
  • 350508 03/06/11 10:58 a-onl End_of_Run_2461
  • 350507 03/06/11 10:46 a-onl Start_Run_2461,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350506 03/06/11 10:45 a-onl End_of_Run_2460
  • 350505 03/06/11 10:39 igorko BigBite TOF spectra
  • 350504 03/06/11 10:31 igorko BigBite ADCs dE/E
  • 350503 03/06/11 10:25 a-onl Start_Run_2460,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350502 03/06/11 10:24 a-onl End_of_Run_2459
  • 350501 03/06/11 10:06 a-onl Start_Run_2459,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350500 03/06/11 10:05 a-onl End_of_Run_2458
  • 350499 03/06/11 09:47 a-onl Start_Run_2458,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350498 03/06/11 09:47 a-onl End_of_Run_2457
  • 350497 03/06/11 09:30 auto_snap Radiation and Ozone snapshot
  • 350496 03/06/11 09:30 auto_snap Miscellaneous snapshot
  • 350495 03/06/11 09:30 auto_snap Hall A Tools snapshot
  • 350494 03/06/11 09:30 auto_snap Compton Polarimeter snapshot
  • 350493 03/06/11 09:30 auto_snap BigBox Magnet snapshot
  • 350492 03/06/11 09:29 auto_snap Beam Position and Current snapshot
  • 350491 03/06/11 09:28 a-onl Start_Run_2457,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350490 03/06/11 09:28 a-onl End_of_Run_2456
  • 350489 03/06/11 09:09 a-onl Start_Run_2456,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350488 03/06/11 09:08 a-onl End_of_Run_2455
  • 350487 03/06/11 09:05 M Defurne screenshots
  • 350486 03/06/11 08:50 a-onl Start_Run_2455,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350485 03/06/11 08:49 a-onl End_of_Run_2454
  • 350484 03/06/11 08:29 a-onl Start_Run_2454,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350483 03/06/11 08:25 a-onl End_of_Run_2453
  • 350482 03/06/11 08:11 jmammei RC Daily Summary
  • 350481 03/06/11 08:08 a-onl Start_Run_2453,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350480 03/06/11 08:07 a-onl End_of_Run_2452
  • 350479 03/06/11 08:00 adaq Beam Time Accounting
  • 350478 03/06/11 07:58 riordan Owl Shift Summary
  • 350477 03/06/11 07:54 riordan Increased Beam Ramp Rate
  • 350476 03/06/11 07:48 a-onl Start_Run_2452,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350475 03/06/11 07:48 a-onl End_of_Run_2451
  • 350474 03/06/11 07:47 jmammei correction: setting #8, not 38
  • 350473 03/06/11 07:46 danez Synch Check - Run 2437
  • 350472 03/06/11 07:30 a-onl Start_Run_2451,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350471 03/06/11 07:27 jmammei Run plan for Suday day/ beginning of swing
  • 350470 03/06/11 07:26 M Friend Target to 4cm LH2
  • 350469 03/06/11 07:22 a-onl End_of_Run_2450
  • 350468 03/06/11 07:11 M Friend Target screenshots
  • 350467 03/06/11 07:02 a-onl Start_Run_2450,Run_type=Other,target_type=4cm dummy,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Aluminum Dummy Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350466 03/06/11 07:01 riordan Zone Recovered - resume Al dummy
  • 350465 03/06/11 06:43 riordan No Beam - RF Zone 2L14 down
  • 350464 03/06/11 06:43 a-onl End_of_Run_2449
  • 350463 03/06/11 06:03 a-onl Start_Run_2449,Run_type=Other,target_type=4cm dummy,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Aluminum Dummy Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350462 03/06/11 06:00 M Friend Target to 4cm Al dummy
  • 350461 03/06/11 05:58 a-onl End_of_Run_2448
  • 350460 03/06/11 05:53 a-onl Start_Run_2448,Run_type=Optics,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=BeO Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350459 03/06/11 05:50 M Friend Target to BeO
  • 350458 03/06/11 05:45 a-onl End_of_Run_2447
  • 350457 03/06/11 05:40 a-onl Start_Run_2447,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Boiling Tests,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350456 03/06/11 05:33 a-onl End_of_Run_2446
  • 350455 03/06/11 05:28 a-onl Start_Run_2446,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=70,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Boiling Tests,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350454 03/06/11 05:25 a-onl End_of_Run_2445
  • 350453 03/06/11 05:20 a-onl Start_Run_2445,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=60,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Boiling Tests,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350452 03/06/11 05:18 a-onl End_of_Run_2444
  • 350451 03/06/11 05:13 a-onl Start_Run_2444,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Boiling Tests,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350450 03/06/11 05:12 a-onl End_of_Run_2443
  • 350449 03/06/11 05:10 a-onl Start_Run_2443,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=50,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Boiling Tests,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350448 03/06/11 05:07 a-onl End_of_Run_2442
  • 350447 03/06/11 05:01 a-onl Start_Run_2442,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Boiling Tests,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350446 03/06/11 04:59 a-onl End_of_Run_2441
  • 350445 03/06/11 04:54 a-onl Start_Run_2441,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Boiling Tests,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350444 03/06/11 04:52 a-onl End_of_Run_2440
  • 350443 03/06/11 04:47 a-onl Start_Run_2440,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Boiling Tests,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350442 03/06/11 04:46 a-onl End_of_Run_2439
  • 350441 03/06/11 04:43 riordan Hall A Tools
  • 350440 03/06/11 04:42 riordan spot++
  • 350439 03/06/11 04:40 a-onl Start_Run_2439,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=10,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Boiling Tests,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350438 03/06/11 04:33 a-onl End_of_Run_2438
  • 350437 03/06/11 04:16 M Friend Target screenshots
  • 350436 03/06/11 04:15 a-onl Start_Run_2438,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350435 03/06/11 04:14 a-onl End_of_Run_2437
  • 350434 03/06/11 03:56 a-onl Start_Run_2437,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350433 03/06/11 03:56 a-onl End_of_Run_2436
  • 350432 03/06/11 03:42 danez Kinematic 12 Screenshot - 80 microamps
  • 350431 03/06/11 03:31 a-onl Start_Run_2436,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.606,leftarm_theta=48,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350430 03/06/11 03:29 danez Synch Check - Run 2399
  • 350429 03/06/11 03:25 riordan LHRS at 48 degrees
  • 350428 03/06/11 02:52 riordan Finish Setting #11 - moving to setting #12
  • 350427 03/06/11 02:51 a-onl End_of_Run_2435
  • 350426 03/06/11 02:33 riordan Gas Shed Flammable Gas Sensor Low Alarm
  • 350425 03/06/11 02:23 a-onl Start_Run_2435,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=55,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350424 03/06/11 02:22 a-onl End_of_Run_2434
  • 350423 03/06/11 02:08 M Friend Target screenshots
  • 350422 03/06/11 01:58 a-onl Start_Run_2434,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=55,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350421 03/06/11 01:56 a-onl End_of_Run_2433
  • 350420 03/06/11 01:39 danez Runs 2431 and 2432 are junk
  • 350419 03/06/11 01:31 a-onl Start_Run_2433,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=55,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350418 03/06/11 01:29 auto_snap Radiation and Ozone snapshot
  • 350417 03/06/11 01:29 auto_snap Miscellaneous snapshot
  • 350416 03/06/11 01:29 auto_snap Hall A Tools snapshot
  • 350415 03/06/11 01:29 auto_snap Compton Polarimeter snapshot
  • 350414 03/06/11 01:29 auto_snap BigBox Magnet snapshot
  • 350413 03/06/11 01:29 auto_snap Beam Position and Current snapshot
  • 350412 03/06/11 01:23 a-onl Start_Run_2432,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=55,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350411 03/06/11 01:17 a-onl Start_Run_2431,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=55,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350410 03/06/11 01:16 a-onl End_of_Run_2430
  • 350409 03/06/11 00:51 a-onl Start_Run_2430,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=55,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350408 03/06/11 00:50 a-onl End_of_Run_2429
  • 350407 03/06/11 00:26 a-onl Start_Run_2429,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=55,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350406 03/06/11 00:24 a-onl End_of_Run_2428
  • 350405 03/06/11 00:03 bogdanw Shift sumamry
  • 350404 03/06/11 00:03 adaq Beam Time Accounting
  • Saturday

  • 350403 03/05/11 23:59 a-onl Start_Run_2428,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=55,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350402 03/05/11 23:59 a-onl End_of_Run_2427
  • 350401 03/05/11 23:33 a-onl Start_Run_2427,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=55,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350400 03/05/11 23:33 a-onl End_of_Run_2426
  • 350399 03/05/11 23:01 a-onl End_of_Run_2425
  • 350398 03/05/11 22:49 flay Target Strip Charts [4cm LH2]
  • 350397 03/05/11 22:47 bogdanw Hall A tools
  • 350396 03/05/11 22:32 flay Moved target back to 4cm LH2 for production at I = 55 uA.
  • 350395 03/05/11 22:26 a-onl End_of_Run_2424
  • 350394 03/05/11 21:52 a-onl Start_Run_2424,Run_type=Optics,target_type=4cm dummy,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Optics Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350393 03/05/11 21:48 flay Moved to 4cm Aluminum target position
  • 350392 03/05/11 21:46 a-onl End_of_Run_2423
  • 350391 03/05/11 21:41 a-onl Start_Run_2423,Run_type=Optics,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Optics Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350390 03/05/11 21:39 flay Target moved to BeO position for 5 min. I - 20 uA run.
  • 350389 03/05/11 21:32 a-onl End_of_Run_2422
  • 350388 03/05/11 21:03 a-onl Start_Run_2422,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=55,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350387 03/05/11 21:03 a-onl End_of_Run_2421
  • 350386 03/05/11 20:38 a-onl Start_Run_2421,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=55,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350385 03/05/11 20:37 a-onl End_of_Run_2420
  • 350384 03/05/11 20:09 a-onl Start_Run_2420,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=55,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350383 03/05/11 20:08 a-onl End_of_Run_2419
  • 350382 03/05/11 19:55 jmammei HPH settings for different targets
  • 350381 03/05/11 19:43 a-onl Start_Run_2419,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=55,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350380 03/05/11 19:42 a-onl End_of_Run_2418
  • 350379 03/05/11 19:38 flay Error on Number of Events for Run 2415
  • 350378 03/05/11 19:28 jmammei boiling study
  • 350377 03/05/11 19:26 boiling study addition to overnight run plan
  • 350376 03/05/11 19:23 flay Summary of Runs 2409--2417
  • 350375 03/05/11 19:18 jmammei run plan for overnight Saturday
  • 350374 03/05/11 19:11 a-onl Start_Run_2418,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=55,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350373 03/05/11 19:09 a-onl End_of_Run_2417
  • 350372 03/05/11 18:51 flay Target Strip Charts [4cm LH2]
  • 350371 03/05/11 18:47 flay Test Runs 2409--2414
  • 350370 03/05/11 18:43 a-onl Start_Run_2417,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=55,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350369 03/05/11 18:41 a-onl Start_Run_2416,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=55,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350368 03/05/11 18:40 a-onl End_of_Run_2415
  • 350367 03/05/11 18:37 orchen HAND tresholds per PMT in MeVee
  • 350366 03/05/11 18:33 a-onl Start_Run_2415,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=55,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350365 03/05/11 18:28 a-onl End_of_Run_2414
  • 350364 03/05/11 18:26 flay N->Delta plots for I = 60 uA on 4cm LH2 target
  • 350363 03/05/11 18:20 flay N->Delta plots for I = 40 uA on 4cm LH2 target
  • 350362 03/05/11 18:15 flay N->Delta plots on 4cm LH2 target at I = 20 uA
  • 350361 03/05/11 18:12 a-onl Start_Run_2414,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=60,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Test,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350360 03/05/11 18:10 a-onl End_of_Run_2413
  • 350359 03/05/11 18:09 a-onl Start_Run_2413,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Test,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350358 03/05/11 18:07 a-onl End_of_Run_2412
  • 350357 03/05/11 18:02 a-onl Start_Run_2412,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Test,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350356 03/05/11 17:58 flay N->Delta plots at I = 30 uA
  • 350355 03/05/11 17:56 jmammei moving to 4cm target
  • 350354 03/05/11 17:55 a-onl End_of_Run_2411
  • 350353 03/05/11 17:51 flay N->Delta Plots for I = 10 uA
  • 350352 03/05/11 17:50 a-onl Start_Run_2411,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Test,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350351 03/05/11 17:49 a-onl End_of_Run_2410
  • 350350 03/05/11 17:45 a-onl Start_Run_2410,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Test,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350349 03/05/11 17:44 a-onl End_of_Run_2409
  • 350348 03/05/11 17:40 a-onl Start_Run_2409,Run_type=Production N -> Delta,target_type=LH2 15cm cell,beam_e0=1.16,beam_raster=on,beam_current=10,leftarm_p=0.606,leftarm_theta=14.5,leftarm_sieve=out,rightarm_p=0.750,rightarm_theta=22,rightarm_sieve=out,comment_text=Test,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350347 03/05/11 17:39 rshneor BigBite VDC eff per HV
  • 350346 03/05/11 17:38 bogdanw Hall A tools for kin #11
  • 350345 03/05/11 17:35 jmammei changed HP heater low high alarm
  • 350344 03/05/11 17:29 auto_snap Radiation and Ozone snapshot
  • 350343 03/05/11 17:29 auto_snap Miscellaneous snapshot
  • 350342 03/05/11 17:29 auto_snap Hall A Tools snapshot
  • 350341 03/05/11 17:29 auto_snap Compton Polarimeter snapshot
  • 350340 03/05/11 17:29 auto_snap BigBox Magnet snapshot
  • 350339 03/05/11 17:29 auto_snap Beam Position and Current snapshot
  • 350338 03/05/11 16:29 flay Target Moved To: 15cm LH2
  • 350337 03/05/11 16:24 a-onl End_of_Run_2408
  • 350336 03/05/11 16:19 flay spot_R for run 2407
  • 350335 03/05/11 16:19 a-onl Start_Run_2408,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.626,leftarm_theta=30,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350334 03/05/11 16:18 a-onl End_of_Run_2407
  • 350333 03/05/11 16:15 M Defurne Day Shift Summary
  • 350332 03/05/11 16:05 a-onl Start_Run_2407,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.626,leftarm_theta=30,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350331 03/05/11 16:04 a-onl End_of_Run_2406
  • 350330 03/05/11 16:02 adaq Beam Time Accounting
  • 350329 03/05/11 15:54 R. Michaels Deadtime vs Rate
  • 350328 03/05/11 15:51 a-onl Start_Run_2406,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.626,leftarm_theta=30,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350327 03/05/11 15:50 a-onl End_of_Run_2405
  • 350326 03/05/11 15:49 M Defurne forgotten screenshot with last emtry
  • 350325 03/05/11 15:46 M Defurne variation of temperature when changing
  • 350324 03/05/11 15:39 a-onl Start_Run_2405,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.626,leftarm_theta=30,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350323 03/05/11 15:38 a-onl End_of_Run_2404
  • 350322 03/05/11 15:30 M Defurne Back to production for setting 5
  • 350321 03/05/11 15:27 a-onl Start_Run_2404,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.626,leftarm_theta=30,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350320 03/05/11 15:27 M Defurne move target to LH2 4cm
  • 350319 03/05/11 15:24 a-onl End_of_Run_2403
  • 350318 03/05/11 15:23 a-onl Start_Run_2403,Run_type=Optics,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.626,leftarm_theta=30,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=junk,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350317 03/05/11 15:06 a-onl End_of_Run_2402
  • 350316 03/05/11 15:00 a-onl Start_Run_2402,Run_type=Optics,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.626,leftarm_theta=30,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Optics Run for 5min with Target BeO,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350315 03/05/11 14:57 adaq run#2401 was junk
  • 350314 03/05/11 14:54 M Defurne BeO target
  • 350313 03/05/11 14:52 a-onl Start_Run_2401,Run_type=Optics,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.626,leftarm_theta=30,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Optics Run for 5min with Target BeO,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350312 03/05/11 14:50 M Defurne move target from dummy 4cm to BeO
  • 350311 03/05/11 14:46 a-onl End_of_Run_2400
  • 350310 03/05/11 14:20 orchen HAND veto gains and tresholds
  • 350309 03/05/11 14:19 M Defurne Start calibration run with dummy 4cm for 30 min
  • 350308 03/05/11 14:16 a-onl Start_Run_2400,Run_type=Calibration,target_type=4cm dummy,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.626,leftarm_theta=30,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=calibration run with Target Al,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350307 03/05/11 14:14 M Defurne move target to dummy 4cm
  • 350306 03/05/11 14:00 a-onl End_of_Run_2399
  • 350305 03/05/11 13:48 a-onl Start_Run_2399,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.626,leftarm_theta=30,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350304 03/05/11 13:46 a-onl End_of_Run_2398
  • 350303 03/05/11 13:35 a-onl Start_Run_2398,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.626,leftarm_theta=30,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350302 03/05/11 13:34 a-onl End_of_Run_2397
  • 350301 03/05/11 13:19 a-onl Start_Run_2397,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.626,leftarm_theta=30,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350300 03/05/11 13:18 a-onl End_of_Run_2396
  • 350299 03/05/11 13:07 a-onl Start_Run_2396,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.626,leftarm_theta=30,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350298 03/05/11 13:06 a-onl End_of_Run_2395
  • 350297 03/05/11 12:55 a-onl Start_Run_2395,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.626,leftarm_theta=30,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350296 03/05/11 12:54 a-onl End_of_Run_2394
  • 350295 03/05/11 12:52 M Defurne Late entry (move LHRS)
  • 350294 03/05/11 12:43 a-onl Start_Run_2394,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.626,leftarm_theta=30,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350293 03/05/11 12:35 a-onl End_of_Run_2393
  • 350292 03/05/11 12:23 a-onl Start_Run_2393,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=70,leftarm_p=0.626,leftarm_theta=30,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350291 03/05/11 12:19 a-onl End_of_Run_2392
  • 350290 03/05/11 12:19 a-onl Start_Run_2392,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=70,leftarm_p=0.626,leftarm_theta=30,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=test run for beam current = 70uA prescale was modified,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350289 03/05/11 12:15 a-onl End_of_Run_2391
  • 350288 03/05/11 12:15 a-onl Start_Run_2391,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=70,leftarm_p=0.626,leftarm_theta=30,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=test run for beam current = 70uA,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350287 03/05/11 12:11 a-onl End_of_Run_2390
  • 350286 03/05/11 12:11 a-onl Start_Run_2390,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=60,leftarm_p=0.626,leftarm_theta=30,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350285 03/05/11 12:03 adaq corrections: run#2389 is test run after left arm moved, mom of left arm=0.626Gev/C and the angle=30degree
  • 350284 03/05/11 11:57 a-onl End_of_Run_2389
  • 350283 03/05/11 11:57 a-onl Start_Run_2389,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350282 03/05/11 11:23 a-onl End_of_Run_2388
  • 350281 03/05/11 11:06 a-onl Start_Run_2388,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350280 03/05/11 11:05 a-onl End_of_Run_2387
  • 350279 03/05/11 10:48 a-onl Start_Run_2387,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350278 03/05/11 10:43 a-onl End_of_Run_2386
  • 350277 03/05/11 10:35 danez Another duplicate file
  • 350276 03/05/11 10:24 a-onl Start_Run_2386,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350275 03/05/11 10:23 a-onl End_of_Run_2385
  • 350274 03/05/11 10:10 M Defurne screenshots
  • 350273 03/05/11 10:06 orchen HAND TOF resolution using Hydrogen Elastic run
  • 350272 03/05/11 10:04 a-onl Start_Run_2385,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350271 03/05/11 10:03 a-onl End_of_Run_2384
  • 350270 03/05/11 09:46 a-onl Start_Run_2384,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350269 03/05/11 09:45 a-onl End_of_Run_2383
  • 350268 03/05/11 09:29 auto_snap Radiation and Ozone snapshot
  • 350267 03/05/11 09:29 auto_snap Miscellaneous snapshot
  • 350266 03/05/11 09:29 auto_snap Hall A Tools snapshot
  • 350265 03/05/11 09:29 auto_snap Compton Polarimeter snapshot
  • 350264 03/05/11 09:29 auto_snap BigBox Magnet snapshot
  • 350263 03/05/11 09:29 auto_snap Beam Position and Current snapshot
  • 350262 03/05/11 09:24 a-onl Start_Run_2383,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350261 03/05/11 09:24 a-onl End_of_Run_2382
  • 350260 03/05/11 09:07 a-onl Start_Run_2382,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350259 03/05/11 09:06 a-onl End_of_Run_2381
  • 350258 03/05/11 09:05 jmammei RC Daily Summary
  • 350257 03/05/11 08:47 a-onl Start_Run_2381,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350256 03/05/11 08:43 a-onl End_of_Run_2380
  • 350255 03/05/11 08:23 a-onl Start_Run_2380,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350254 03/05/11 08:22 a-onl End_of_Run_2379
  • 350253 03/05/11 08:04 M Defurne screenshots at 8 am
  • 350252 03/05/11 08:04 a-onl Start_Run_2379,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350251 03/05/11 08:01 adaq Beam Time Accounting
  • 350250 03/05/11 07:56 riordan Owl Shift Summary
  • 350249 03/05/11 07:53 a-onl End_of_Run_2378
  • 350248 03/05/11 07:35 a-onl Start_Run_2378,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350247 03/05/11 07:34 a-onl End_of_Run_2377
  • 350246 03/05/11 07:15 a-onl Start_Run_2377,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350245 03/05/11 07:11 a-onl End_of_Run_2376
  • 350244 03/05/11 07:04 M Friend Target screenshots
  • 350243 03/05/11 06:54 a-onl Start_Run_2376,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350242 03/05/11 06:54 a-onl End_of_Run_2375
  • 350241 03/05/11 06:33 a-onl Start_Run_2375,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350240 03/05/11 06:29 a-onl End_of_Run_2374
  • 350239 03/05/11 06:05 a-onl Start_Run_2374,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350238 03/05/11 06:04 a-onl End_of_Run_2373
  • 350237 03/05/11 05:43 a-onl Start_Run_2373,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350236 03/05/11 05:42 a-onl End_of_Run_2372
  • 350235 03/05/11 05:25 a-onl Start_Run_2372,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350234 03/05/11 05:24 M Friend Target snapshots
  • 350233 03/05/11 05:23 M Friend Target to 4cm LH2
  • 350232 03/05/11 05:21 a-onl End_of_Run_2371
  • 350231 03/05/11 05:12 a-onl Start_Run_2371,Run_type=Optics,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=BeO run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350230 03/05/11 05:10 M Friend Target to BeO
  • 350229 03/05/11 05:08 a-onl End_of_Run_2370
  • 350228 03/05/11 04:07 a-onl Start_Run_2370,Run_type=Calibration,target_type=4cm dummy,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Al dummy test,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350227 03/05/11 04:04 M Friend Target to 4cm Al dummy
  • 350226 03/05/11 04:00 a-onl End_of_Run_2369
  • 350225 03/05/11 03:53 riordan Hall A Tools
  • 350224 03/05/11 03:42 a-onl Start_Run_2369,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350223 03/05/11 03:42 a-onl End_of_Run_2368
  • 350222 03/05/11 03:39 M Friend JT valve stepped open
  • 350221 03/05/11 03:25 a-onl Start_Run_2368,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350220 03/05/11 03:24 a-onl End_of_Run_2367
  • 350219 03/05/11 03:14 nmuangma BCM calibration against OLO2 and OLO2 against Faraday Cup
  • 350218 03/05/11 03:07 a-onl Start_Run_2367,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350217 03/05/11 03:07 a-onl End_of_Run_2366
  • 350216 03/05/11 02:56 riordan spot++
  • 350215 03/05/11 02:56 danez N-Delta Screenshot
  • 350214 03/05/11 02:50 a-onl Start_Run_2366,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350213 03/05/11 02:49 a-onl End_of_Run_2365
  • 350212 03/05/11 02:44 a-onl Start_Run_2365,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=80,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Test Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350211 03/05/11 02:36 a-onl End_of_Run_2364
  • 350210 03/05/11 02:35 a-onl Start_Run_2364,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=70,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Test Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350209 03/05/11 02:26 a-onl End_of_Run_2363
  • 350208 03/05/11 02:25 a-onl Start_Run_2363,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=60,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Test Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350207 03/05/11 02:20 a-onl End_of_Run_2362
  • 350206 03/05/11 02:19 a-onl Start_Run_2362,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Test Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350205 03/05/11 02:11 a-onl End_of_Run_2361
  • 350204 03/05/11 02:10 a-onl Start_Run_2361,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=46,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Test Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350203 03/05/11 02:10 riordan FFB On
  • 350202 03/05/11 02:07 M Friend Target screenshots
  • 350201 03/05/11 02:00 riordan LHRS at 46 degrees
  • 350200 03/05/11 01:48 M Friend Target to 4cm LH2
  • 350199 03/05/11 01:34 riordan Controlled Access for Spectrometer movement
  • 350198 03/05/11 01:28 auto_snap Radiation and Ozone snapshot
  • 350197 03/05/11 01:28 auto_snap Miscellaneous snapshot
  • 350196 03/05/11 01:28 auto_snap Hall A Tools snapshot
  • 350195 03/05/11 01:28 auto_snap Compton Polarimeter snapshot
  • 350194 03/05/11 01:28 auto_snap BigBox Magnet snapshot
  • 350193 03/05/11 01:28 auto_snap Beam Position and Current snapshot
  • 350192 03/05/11 01:02 M Friend Target to BeO
  • 350191 03/05/11 00:57 a-onl End_of_Run_2360
  • 350189 03/05/11 00:03 adaq Beam Time Accounting
  • Friday

  • 350188 03/04/11 23:34 a-onl Start_Run_2360,Run_type=Other,target_type=4cm dummy,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Dummy Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350187 03/04/11 23:32 H. Albataine Target moved to Al
  • 350186 03/04/11 23:26 a-onl End_of_Run_2359
  • 350185 03/04/11 23:21 a-onl Start_Run_2359,Run_type=Spot++,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Spot++ run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350184 03/04/11 23:19 H. Albataine Target moved to BeO
  • 350183 03/04/11 23:14 a-onl End_of_Run_2358
  • 350182 03/04/11 23:13 a-onl Start_Run_2358,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350181 03/04/11 23:12 a-onl End_of_Run_2357
  • 350180 03/04/11 23:01 Chao spot_R result for run 2356
  • 350179 03/04/11 23:00 a-onl Start_Run_2357,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350178 03/04/11 23:00 a-onl End_of_Run_2356
  • 350177 03/04/11 22:56 yez I fixed it by removing FPP,BB,and HAND out of the replay script
  • 350176 03/04/11 22:54 R. Michaels two datafiles for 2349
  • 350175 03/04/11 22:44 a-onl Start_Run_2356,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350174 03/04/11 22:44 a-onl End_of_Run_2355
  • 350173 03/04/11 22:25 a-onl Start_Run_2355,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350172 03/04/11 22:24 a-onl End_of_Run_2354
  • 350171 03/04/11 22:13 Chao Missing Data File
  • 350170 03/04/11 22:11 a-onl Start_Run_2354,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350169 03/04/11 22:05 a-onl End_of_Run_2353
  • 350168 03/04/11 21:54 a-onl Start_Run_2353,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350167 03/04/11 21:53 a-onl End_of_Run_2352
  • 350166 03/04/11 21:38 a-onl Start_Run_2352,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350165 03/04/11 21:37 a-onl End_of_Run_2351
  • 350164 03/04/11 21:23 a-onl Start_Run_2351,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350163 03/04/11 21:21 a-onl End_of_Run_2350
  • 350162 03/04/11 21:08 a-onl Start_Run_2350,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350161 03/04/11 21:06 a-onl End_of_Run_2349
  • 350160 03/04/11 20:53 a-onl Start_Run_2349,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350159 03/04/11 20:52 a-onl End_of_Run_2348
  • 350158 03/04/11 20:37 a-onl Start_Run_2348,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350157 03/04/11 20:36 a-onl End_of_Run_2347
  • 350156 03/04/11 20:24 a-onl Start_Run_2347,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350155 03/04/11 20:22 a-onl End_of_Run_2346
  • 350154 03/04/11 20:10 jmammei Al and BeO runs
  • 350153 03/04/11 20:09 a-onl Start_Run_2346,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350152 03/04/11 20:08 a-onl End_of_Run_2345
  • 350151 03/04/11 19:54 a-onl Start_Run_2345,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350150 03/04/11 19:54 a-onl End_of_Run_2344
  • 350149 03/04/11 19:41 a-onl Start_Run_2344,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350148 03/04/11 19:40 a-onl End_of_Run_2343
  • 350147 03/04/11 19:29 a-onl Start_Run_2343,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350146 03/04/11 19:28 a-onl End_of_Run_2342
  • 350145 03/04/11 19:15 a-onl Start_Run_2342,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350144 03/04/11 19:14 a-onl End_of_Run_2341
  • 350143 03/04/11 18:58 a-onl Start_Run_2341,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350142 03/04/11 18:57 a-onl End_of_Run_2340
  • 350141 03/04/11 18:44 a-onl Start_Run_2340,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350140 03/04/11 18:43 a-onl End_of_Run_2339
  • 350139 03/04/11 18:29 a-onl Start_Run_2339,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350138 03/04/11 18:28 a-onl End_of_Run_2338
  • 350137 03/04/11 18:16 a-onl Start_Run_2338,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350136 03/04/11 18:04 a-onl Start_Run_2337,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350135 03/04/11 18:03 a-onl End_of_Run_2336
  • 350134 03/04/11 17:51 a-onl Start_Run_2336,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350133 03/04/11 17:50 a-onl End_of_Run_2335
  • 350132 03/04/11 17:39 a-onl Start_Run_2335,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350131 03/04/11 17:38 a-onl End_of_Run_2334
  • 350130 03/04/11 17:28 auto_snap Radiation and Ozone snapshot
  • 350129 03/04/11 17:28 auto_snap Miscellaneous snapshot
  • 350128 03/04/11 17:28 auto_snap Hall A Tools snapshot
  • 350127 03/04/11 17:28 auto_snap Compton Polarimeter snapshot
  • 350126 03/04/11 17:28 auto_snap BigBox Magnet snapshot
  • 350125 03/04/11 17:28 auto_snap Beam Position and Current snapshot
  • 350124 03/04/11 17:26 a-onl Start_Run_2334,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350123 03/04/11 17:25 a-onl End_of_Run_2333
  • 350122 03/04/11 17:18 A. Camsonne Additionnal sync plots
  • 350121 03/04/11 17:14 V. Sulkosky Work Disk getting full
  • 350120 03/04/11 17:13 a-onl Start_Run_2333,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350119 03/04/11 17:12 a-onl End_of_Run_2332
  • 350118 03/04/11 16:52 a-onl Start_Run_2332,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350117 03/04/11 16:52 a-onl End_of_Run_2331
  • 350116 03/04/11 16:38 H. Albataine Target snap shots
  • 350115 03/04/11 16:34 a-onl Start_Run_2331,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350114 03/04/11 16:33 a-onl End_of_Run_2330
  • 350113 03/04/11 16:20 orchen this is from run number 2270
  • 350112 03/04/11 16:19 a-onl Start_Run_2330,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350111 03/04/11 16:19 a-onl End_of_Run_2329
  • 350110 03/04/11 16:13 adaq Beam Time Accounting
  • 350109 03/04/11 16:11 yez PRL2 ADC are fine but something wrong with the repaly script
  • 350108 03/04/11 16:06 a-onl Start_Run_2329,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350107 03/04/11 16:05 a-onl End_of_Run_2328
  • 350106 03/04/11 15:53 a-onl Start_Run_2328,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350105 03/04/11 15:52 zwzhao Shift Summary
  • 350104 03/04/11 15:52 a-onl End_of_Run_2327
  • 350103 03/04/11 15:40 a-onl Start_Run_2327,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350102 03/04/11 15:39 a-onl End_of_Run_2326
  • 350101 03/04/11 15:31 flay Sync roc plots for Run 2324
  • 350100 03/04/11 15:26 a-onl Start_Run_2326,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350099 03/04/11 15:25 a-onl End_of_Run_2325
  • 350098 03/04/11 15:14 a-onl Start_Run_2325,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350097 03/04/11 15:13 a-onl End_of_Run_2324
  • 350096 03/04/11 15:06 zwzhao turn pion rejector HV on
  • 350095 03/04/11 15:00 a-onl Start_Run_2324,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350094 03/04/11 14:59 a-onl End_of_Run_2323
  • 350093 03/04/11 14:51 jmammei Run Plan for overnight Friday
  • 350092 03/04/11 14:50 flay Run 2322 is Junk
  • 350091 03/04/11 14:48 a-onl Start_Run_2323,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350090 03/04/11 14:44 a-onl Start_Run_2322,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350089 03/04/11 14:43 a-onl End_of_Run_2321
  • 350088 03/04/11 14:32 a-onl Start_Run_2321,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350087 03/04/11 14:31 a-onl End_of_Run_2320
  • 350086 03/04/11 14:30 zwzhao Halla Tolls sanpshot
  • 350085 03/04/11 14:17 a-onl End_of_Run_2319
  • 350084 03/04/11 14:12 flay LHRS Pion Rejector ADCs
  • 350083 03/04/11 14:11 a-onl End_of_Run_2318
  • 350082 03/04/11 14:01 orchen missing mass and conincedence time
  • 350081 03/04/11 13:58 a-onl End_of_Run_2317
  • 350080 03/04/11 13:40 a-onl Start_Run_2317,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350079 03/04/11 13:39 a-onl End_of_Run_2316
  • 350078 03/04/11 13:27 a-onl Start_Run_2316,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350077 03/04/11 13:26 a-onl End_of_Run_2315
  • 350076 03/04/11 13:23 V. Sulkosky Run 2314 TOF/Mmiss 40 uA
  • 350075 03/04/11 13:15 a-onl Start_Run_2315,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350074 03/04/11 13:14 a-onl End_of_Run_2314
  • 350073 03/04/11 13:03 a-onl Start_Run_2314,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Test,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350072 03/04/11 13:01 a-onl End_of_Run_2313
  • 350071 03/04/11 12:59 a-onl Start_Run_2313,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=40,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Test,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350070 03/04/11 12:59 a-onl End_of_Run_2312
  • 350069 03/04/11 12:58 X. Zheng Target snap shot
  • 350068 03/04/11 12:57 X. Zheng Numbers for target cooling power and beam heating
  • 350067 03/04/11 12:46 a-onl Start_Run_2312,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Test,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350066 03/04/11 12:45 a-onl End_of_Run_2311
  • 350065 03/04/11 12:35 a-onl Start_Run_2311,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=20,leftarm_p=0.576,leftarm_theta=13.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=18.0,rightarm_sieve=out,comment_text=Junk,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350064 03/04/11 12:32 X. Zheng Target moved to 4cm loop2
  • 350063 03/04/11 12:30 zwzhao spot_R result
  • 350062 03/04/11 12:27 a-onl End_of_Run_2310
  • 350061 03/04/11 12:24 yez HRS Detectors Update:
  • 350060 03/04/11 12:22 a-onl Start_Run_2310,Run_type=Spot++,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.576,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.770,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Spot++ on BeO Target,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350059 03/04/11 12:15 X. Zheng Target to BeO
  • 350058 03/04/11 12:10 danez FPP Information from CC White Board
  • 350057 03/04/11 12:06 X. Zheng target moved to empty
  • 350056 03/04/11 12:04 a-onl End_of_Run_2309
  • 350055 03/04/11 12:03 a-onl Start_Run_2309,Run_type=Check out,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=deadtime chceck,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350054 03/04/11 11:52 a-onl End_of_Run_2308
  • 350053 03/04/11 11:45 saw bbts1 scaler event format
  • 350052 03/04/11 11:34 adaq L-HRS: S1 HVs and scaler before changing HVs
  • 350051 03/04/11 11:12 Higinbotham [techs] Right Q1 Tripped
  • 350050 03/04/11 11:07 a-onl Start_Run_2308,Run_type=Check out,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=deadtime chceck,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350049 03/04/11 11:04 V. Sulkosky HAND HV OFF
  • 350048 03/04/11 10:58 R. Michaels retored trigger
  • 350047 03/04/11 10:54 atrig Trigger-Download
  • 350046 03/04/11 10:52 Bob / Alex modification to bb_sfi3.crl
  • 350045 03/04/11 10:52 a-onl End_of_Run_2307
  • 350044 03/04/11 10:44 a-onl Start_Run_2307,Run_type=Check out,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=deadtime chceck,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350043 03/04/11 10:40 a-onl End_of_Run_2306
  • 350042 03/04/11 10:35 a-onl Start_Run_2306,Run_type=Check out,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=deadtime chceck,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350041 03/04/11 10:33 a-onl End_of_Run_2305
  • 350040 03/04/11 10:32 a-onl Start_Run_2305,Run_type=Check out,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=deadtime chceck,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350039 03/04/11 10:31 a-onl End_of_Run_2304
  • 350038 03/04/11 10:27 a-onl Start_Run_2304,Run_type=Check out,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=deadtime chceck,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350037 03/04/11 10:25 a-onl End_of_Run_2303
  • 350036 03/04/11 10:13 Bob/Alex deadtime checks, part III
  • 350035 03/04/11 10:09 Higibotham Left Arm @ 13.5
  • 350034 03/04/11 10:08 a-onl Start_Run_2303,Run_type=Check out,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=deadtime chceck,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350033 03/04/11 10:06 a-onl End_of_Run_2302
  • 350032 03/04/11 09:59 danez T4 Not Connected Correctly 2/27 - 3/4
  • 350031 03/04/11 09:55 a-onl Start_Run_2302,Run_type=Check out,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=deadtime chceck,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350030 03/04/11 09:41 danez Cross-Connected Strobe Signals
  • 350029 03/04/11 09:40 a-onl End_of_Run_2300
  • 350028 03/04/11 09:31 a-onl Start_Run_2300,Run_type=Check out,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=deadtime chceck,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350027 03/04/11 09:28 auto_snap Radiation and Ozone snapshot
  • 350026 03/04/11 09:28 auto_snap Miscellaneous snapshot
  • 350025 03/04/11 09:28 auto_snap Hall A Tools snapshot
  • 350024 03/04/11 09:28 auto_snap Compton Polarimeter snapshot
  • 350023 03/04/11 09:28 auto_snap BigBox Magnet snapshot
  • 350022 03/04/11 09:28 auto_snap Beam Position and Current snapshot
  • 350021 03/04/11 09:26 a-onl End_of_Run_2299
  • 350020 03/04/11 09:10 A. Camsonne Busy ROC3 and ROC4
  • 350019 03/04/11 08:59 igorko BigBite momentum
  • 350018 03/04/11 08:58 V. Sulkosky Widen T1 trigger to 50 ns
  • 350017 03/04/11 08:55 V. Sulkosky Changed EDTM Setup for new Kinematic
  • 350016 03/04/11 08:52 R. Michaels deadtime checks, part II
  • 350015 03/04/11 08:52 zwzhao Q1 ramp down
  • 350014 03/04/11 08:49 jmammei RC Daily Summary
  • 350013 03/04/11 08:30 a-onl Start_Run_2299,Run_type=Check out,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=deadtime chceck,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 350012 03/04/11 08:29 R. Michaels download loose triggers, deadtime check
  • 350011 03/04/11 08:23 zwzhao HV on for neutron detector
  • 350010 03/04/11 08:13 V. Sulkosky Turned bbps1:2008 HAND HV On
  • 350009 03/04/11 08:02 adaq Beam Time Accounting
  • 350008 03/04/11 08:02 riordan Owl Shift Summary
  • 350007 03/04/11 08:02 riordan Owl Shift Summary
  • 350006 03/04/11 07:44 danez Deadtime Monitor moved to adaql1
  • 350005 03/04/11 07:37 riordan Sweep done - Hall in controlled access
  • 350004 03/04/11 07:10 riordan Spectrometer Magnets to Setting #6
  • 350003 03/04/11 07:08 riordan Controlled Access
  • 350002 03/04/11 07:06 jmammei call for access at 7am
  • 350001 03/04/11 07:03 danez Run 2297 Ended with CODA crash
  • 350000 03/04/11 07:00 a-onl End_of_Run_2298
  • 349999 03/04/11 06:48 a-onl Start_Run_2298,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349998 03/04/11 06:19 a-onl Start_Run_2297,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349997 03/04/11 06:18 a-onl End_of_Run_2296
  • 349996 03/04/11 06:10 M Friend Target screenshots
  • 349995 03/04/11 05:56 M Friend Compton EPICS reboot
  • 349994 03/04/11 05:52 a-onl Start_Run_2296,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349993 03/04/11 05:51 a-onl End_of_Run_2295
  • 349992 03/04/11 05:24 a-onl Start_Run_2295,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349991 03/04/11 05:24 a-onl End_of_Run_2294
  • 349990 03/04/11 04:56 a-onl Start_Run_2294,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349989 03/04/11 04:56 a-onl End_of_Run_2293
  • 349988 03/04/11 04:30 a-onl Start_Run_2293,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349987 03/04/11 04:29 a-onl End_of_Run_2292
  • 349986 03/04/11 04:05 riordan Hall A Tools
  • 349985 03/04/11 04:03 danez CODA Crash - Runs 2289-2291 are junk
  • 349984 03/04/11 04:02 M Friend Target snapshots
  • 349983 03/04/11 04:01 a-onl Start_Run_2292,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349982 03/04/11 03:44 a-onl Start_Run_2289,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349981 03/04/11 03:44 a-onl End_of_Run_2288
  • 349980 03/04/11 03:24 yez Put a bigger plot to show the detail: x->TDC, y->ADC
  • 349979 03/04/11 03:18 yez R-HRS: Cerenkov double TDC peaks issue
  • 349978 03/04/11 03:18 a-onl Start_Run_2288,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349977 03/04/11 03:16 a-onl End_of_Run_2287
  • 349976 03/04/11 02:49 a-onl Start_Run_2287,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349975 03/04/11 02:48 a-onl End_of_Run_2286
  • 349974 03/04/11 02:21 a-onl Start_Run_2286,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production - changed prescales,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349973 03/04/11 02:19 a-onl End_of_Run_2285
  • 349972 03/04/11 02:14 yez L-HRS: S1 HV adjustment plan
  • 349971 03/04/11 02:02 M Friend Target screenshots
  • 349970 03/04/11 01:51 a-onl Start_Run_2285,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349969 03/04/11 01:50 a-onl End_of_Run_2284
  • 349968 03/04/11 01:27 auto_snap Radiation and Ozone snapshot
  • 349967 03/04/11 01:27 auto_snap Miscellaneous snapshot
  • 349966 03/04/11 01:27 auto_snap Hall A Tools snapshot
  • 349965 03/04/11 01:27 auto_snap Compton Polarimeter snapshot
  • 349964 03/04/11 01:27 auto_snap BigBox Magnet snapshot
  • 349963 03/04/11 01:27 auto_snap Beam Position and Current snapshot
  • 349962 03/04/11 01:21 a-onl Start_Run_2284,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production - DTMon off,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349961 03/04/11 01:20 yez R-HRS: S1 HV adjustment plan (wait for conform)
  • 349960 03/04/11 01:19 a-onl End_of_Run_2283
  • 349959 03/04/11 01:09 lselvy End of run beep restored
  • 349958 03/04/11 00:53 yez RHRS HV setting before doing detectors ADC alignment
  • 349957 03/04/11 00:50 a-onl Start_Run_2283,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production - DTMon off,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349956 03/04/11 00:49 a-onl End_of_Run_2282
  • 349955 03/04/11 00:46 riordan spot++ Run 2282
  • 349954 03/04/11 00:45 yez LHRS HV setting before detectors alignment
  • 349953 03/04/11 00:38 yez Verification of Shower and PreShower map (Look OK)
  • 349952 03/04/11 00:19 a-onl Start_Run_2282,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=BeO run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349951 03/04/11 00:17 M Friend Target to 4cm LH2
  • 349950 03/04/11 00:13 a-onl End_of_Run_2281
  • 349949 03/04/11 00:07 a-onl Start_Run_2281,Run_type=Optics,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=BeO run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349948 03/04/11 00:06 bogdanw shift summary
  • 349947 03/04/11 00:04 adaq Beam Time Accounting
  • 349946 03/04/11 00:04 bogdanw kin #1 HRSs
  • 349945 03/04/11 00:03 M Friend Target to BeO
  • Thursday

  • 349944 03/03/11 23:12 a-onl End_of_Run_2280
  • 349943 03/03/11 22:56 mhuang N-Delta run2279 Check
  • 349942 03/03/11 22:56 jmammei pedestals in ped1 and 2
  • 349941 03/03/11 22:54 pzhu target status
  • 349940 03/03/11 22:54 a-onl Start_Run_2280,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=setting#3,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349939 03/03/11 22:54 a-onl End_of_Run_2279
  • 349938 03/03/11 22:27 a-onl Start_Run_2279,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=setting#3,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349937 03/03/11 22:26 a-onl End_of_Run_2278
  • 349936 03/03/11 22:22 yez Not all channels in ROC1,2 ADC have pedestal subpression on
  • 349935 03/03/11 22:21 mhuang N-Delta Run 2278 Check
  • 349934 03/03/11 22:17 mhuang N-Delta run2277 check
  • 349933 03/03/11 22:13 bogdanw Hall A tools
  • 349932 03/03/11 22:05 a-onl End_of_Run_2277
  • 349931 03/03/11 21:46 yez RHRS HV setting during APEX
  • 349930 03/03/11 21:44 adaq N-Delta Run Check
  • 349929 03/03/11 21:42 pzhu target status
  • 349928 03/03/11 21:40 a-onl End_of_Run_2276
  • 349927 03/03/11 21:33 jmammei start of runs not logging again?
  • 349926 03/03/11 21:27 jmammei prescale gui not top user, but java is at ~15%
  • 349925 03/03/11 21:26 jmammei prescale gui not top user, but java is at ~15%
  • 349924 03/03/11 21:25 jmammei N->Delta HowToPage
  • 349923 03/03/11 21:21 jmammei run 2273 was high DT test run
  • 349922 03/03/11 21:17 a-onl End_of_Run_2275
  • 349921 03/03/11 20:57 bogdanw Max DAQ rate
  • 349920 03/03/11 20:55 R. Michaels re: prescale GUI (run on adaql1 ?)
  • 349919 03/03/11 20:41 a-onl End_of_Run_2273
  • 349918 03/03/11 20:39 R. Michaels re: deadtime scan data
  • 349917 03/03/11 20:38 a-onl End_of_Run_2272
  • 349916 03/03/11 20:36 pzhu target status
  • 349915 03/03/11 20:15 a-onl Start_Run_2272,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=setting#3,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349914 03/03/11 20:14 a-onl End_of_Run_2271
  • 349913 03/03/11 20:01 igorko Software gains for BB ADCs
  • 349912 03/03/11 19:52 Higinbotham High Rate Test
  • 349911 03/03/11 19:49 a-onl Start_Run_2271,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=setting#3,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349910 03/03/11 19:48 a-onl End_of_Run_2270
  • 349909 03/03/11 19:40 pzhu target status
  • 349908 03/03/11 19:24 a-onl Start_Run_2270,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=setting#3,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349907 03/03/11 19:22 a-onl End_of_Run_2269
  • 349906 03/03/11 18:51 adaq online analysis GUI
  • 349905 03/03/11 18:49 pzhu Hall A tool screen
  • 349904 03/03/11 18:42 a-onl Start_Run_2269,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DAQ TESTING,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349903 03/03/11 18:40 a-onl End_of_Run_2268
  • 349902 03/03/11 18:10 lselvy Deadtime Tests
  • 349901 03/03/11 18:08 Kalyan Prescale GUI is working
  • 349900 03/03/11 18:03 pzhu target status
  • 349899 03/03/11 18:03 a-onl Start_Run_2268,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DAQ TESTING,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349898 03/03/11 17:58 a-onl End_of_Run_2267
  • 349897 03/03/11 17:58 a-onl Start_Run_2267,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DAQ TESTING,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349896 03/03/11 17:57 a-onl End_of_Run_2266
  • 349895 03/03/11 17:56 a-onl Start_Run_2266,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DAQ TESTING,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349894 03/03/11 17:55 a-onl End_of_Run_2265
  • 349893 03/03/11 17:54 a-onl Start_Run_2265,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DAQ TESTING,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349892 03/03/11 17:54 a-onl End_of_Run_2264
  • 349891 03/03/11 17:53 a-onl Start_Run_2264,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DAQ TESTING,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349890 03/03/11 17:52 a-onl End_of_Run_2263
  • 349889 03/03/11 17:51 a-onl Start_Run_2263,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DAQ TESTING,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349888 03/03/11 17:51 a-onl End_of_Run_2262
  • 349887 03/03/11 17:50 a-onl Start_Run_2262,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DAQ TESTING,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349886 03/03/11 17:49 a-onl End_of_Run_2261
  • 349885 03/03/11 17:49 a-onl Start_Run_2261,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DAQ TESTING,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349884 03/03/11 17:48 a-onl End_of_Run_2260
  • 349883 03/03/11 17:47 benesch puzzle on hall A beam size
  • 349882 03/03/11 17:44 a-onl Start_Run_2260,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DAQ TESTING,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349881 03/03/11 17:41 jmammei prescale gui not working
  • 349880 03/03/11 17:27 auto_snap Radiation and Ozone snapshot
  • 349879 03/03/11 17:27 auto_snap Miscellaneous snapshot
  • 349878 03/03/11 17:27 auto_snap Hall A Tools snapshot
  • 349877 03/03/11 17:27 auto_snap Compton Polarimeter snapshot
  • 349876 03/03/11 17:27 auto_snap BigBox Magnet snapshot
  • 349875 03/03/11 17:27 auto_snap Beam Position and Current snapshot
  • 349874 03/03/11 16:34 a-onl End_of_Run_2259
  • 349873 03/03/11 16:21 a-onl Start_Run_2259,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349872 03/03/11 16:20 pzhu target status
  • 349871 03/03/11 16:19 a-onl End_of_Run_2258
  • 349870 03/03/11 16:16 lselvy Pedestal Supression
  • 349869 03/03/11 16:01 a-onl Start_Run_2258,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349868 03/03/11 16:01 adaq Beam Time Accounting
  • 349867 03/03/11 16:01 Saha Shift Summary
  • 349866 03/03/11 16:00 a-onl End_of_Run_2257
  • 349865 03/03/11 15:42 a-onl Start_Run_2257,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349864 03/03/11 15:40 flay Rn 2256 is a Pedestal Run
  • 349863 03/03/11 15:39 a-onl End_of_Run_2255
  • 349862 03/03/11 15:30 orchen HRS timing calibration
  • 349861 03/03/11 15:27 a-onl End_of_Run_2255
  • 349860 03/03/11 15:09 Holmstrom Target Move
  • 349859 03/03/11 15:07 a-onl Start_Run_2255,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349858 03/03/11 15:03 lselvy Readout time for ROC 10
  • 349857 03/03/11 14:54 Saha Hall A BCM Linearity Test
  • 349856 03/03/11 14:46 a-onl End_of_Run_2254
  • 349855 03/03/11 14:25 flay Hall A Tools
  • 349854 03/03/11 14:21 flay Spot++ for Run 2254
  • 349853 03/03/11 14:16 R. Michaels added locftm to config string for pedrun
  • 349852 03/03/11 14:12 a-onl Start_Run_2254,Run_type=Calibration,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=BCM Calibration Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349851 03/03/11 13:39 a-onl End_of_Run_2253
  • 349850 03/03/11 13:32 flay Target Type for Run 2252
  • 349849 03/03/11 13:24 a-onl Start_Run_2253,Run_type=Calibration,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=BCM Calibration Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349848 03/03/11 12:49 Saha BCM OLO2 - FC 2 linearity calibration
  • 349847 03/03/11 12:47 a-onl End_of_Run_2252
  • 349846 03/03/11 12:28 a-onl Start_Run_2252,Run_type=Calibration,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=BCM Calibration Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349845 03/03/11 12:16 Saha BCM: OLO2 -FC2 calibration
  • 349844 03/03/11 11:51 jmammei settings for kinematic setting #1
  • 349843 03/03/11 11:45 jmammei do another BeO run at setting #1
  • 349842 03/03/11 11:26 Holmstrom Move Target to Empty
  • 349841 03/03/11 11:23 jmammei Ole's script auto-magically removes the Al runs
  • 349840 03/03/11 11:16 jmammei run plan for overnight Thursday
  • 349839 03/03/11 11:13 a-onl End_of_Run_2251
  • 349838 03/03/11 10:52 a-onl End_of_Run_2250
  • 349837 03/03/11 10:45 jmammei RC Daily Summary
  • 349836 03/03/11 10:33 a-onl Start_Run_2250,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349835 03/03/11 10:31 a-onl End_of_Run_2249
  • 349834 03/03/11 10:16 lselvy Readout time for BBTS1
  • 349833 03/03/11 10:13 a-onl Start_Run_2249,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349832 03/03/11 10:12 a-onl End_of_Run_2248
  • 349831 03/03/11 10:08 orchen the change
  • 349830 03/03/11 10:03 orchen BB optics DB change
  • 349829 03/03/11 09:53 a-onl Start_Run_2248,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349828 03/03/11 09:53 a-onl End_of_Run_2247
  • 349827 03/03/11 09:44 V. Sulkosky Trigger supervisor deadtime
  • 349826 03/03/11 09:43 V. Sulkosky Trigger supervisor deadtime
  • 349825 03/03/11 09:33 a-onl Start_Run_2247,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349824 03/03/11 09:32 a-onl End_of_Run_2246
  • 349823 03/03/11 09:27 auto_snap Radiation and Ozone snapshot
  • 349822 03/03/11 09:27 auto_snap Miscellaneous snapshot
  • 349821 03/03/11 09:27 auto_snap Hall A Tools snapshot
  • 349820 03/03/11 09:27 auto_snap Compton Polarimeter snapshot
  • 349819 03/03/11 09:27 auto_snap BigBox Magnet snapshot
  • 349818 03/03/11 09:27 auto_snap Beam Position and Current snapshot
  • 349817 03/03/11 09:12 a-onl Start_Run_2246,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349816 03/03/11 09:10 a-onl End_of_Run_2245
  • 349815 03/03/11 08:52 a-onl Start_Run_2245,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349814 03/03/11 08:51 a-onl End_of_Run_2244
  • 349813 03/03/11 08:33 a-onl Start_Run_2244,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349812 03/03/11 08:32 a-onl End_of_Run_2243
  • 349811 03/03/11 08:13 a-onl Start_Run_2243,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349810 03/03/11 08:13 a-onl End_of_Run_2242
  • 349809 03/03/11 08:01 adaq Beam Time Accounting
  • 349808 03/03/11 07:58 riordan Owl Shift Summary
  • 349807 03/03/11 07:58 Holmstrom Target sanpshot
  • 349806 03/03/11 07:52 a-onl Start_Run_2242,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349805 03/03/11 07:51 a-onl End_of_Run_2241
  • 349804 03/03/11 07:32 a-onl Start_Run_2241,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349803 03/03/11 07:32 a-onl End_of_Run_2240
  • 349802 03/03/11 07:14 a-onl Start_Run_2240,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349801 03/03/11 07:13 a-onl End_of_Run_2239
  • 349800 03/03/11 06:54 a-onl Start_Run_2239,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349799 03/03/11 06:53 a-onl End_of_Run_2238
  • 349798 03/03/11 06:33 a-onl Start_Run_2238,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349797 03/03/11 06:28 a-onl End_of_Run_2237
  • 349796 03/03/11 06:10 a-onl Start_Run_2237,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349795 03/03/11 06:09 a-onl End_of_Run_2236
  • 349794 03/03/11 06:07 M Friend Target snapshots
  • 349793 03/03/11 05:51 a-onl Start_Run_2236,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349792 03/03/11 05:50 a-onl End_of_Run_2235
  • 349791 03/03/11 05:32 a-onl Start_Run_2235,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349790 03/03/11 05:31 a-onl End_of_Run_2234
  • 349789 03/03/11 05:14 a-onl Start_Run_2234,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349788 03/03/11 05:12 a-onl End_of_Run_2233
  • 349787 03/03/11 04:53 a-onl Start_Run_2233,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349786 03/03/11 04:53 a-onl End_of_Run_2232
  • 349785 03/03/11 04:51 riordan RF Feedback off for ~40min
  • 349784 03/03/11 04:42 M Friend Target alarm
  • 349783 03/03/11 04:34 a-onl Start_Run_2232,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349782 03/03/11 04:33 a-onl End_of_Run_2231
  • 349781 03/03/11 04:14 a-onl End_of_Run_2230
  • 349780 03/03/11 04:10 M Friend Target screenshots
  • 349779 03/03/11 03:54 a-onl End_of_Run_2229
  • 349778 03/03/11 03:34 a-onl End_of_Run_2228
  • 349777 03/03/11 03:15 a-onl End_of_Run_2227
  • 349776 03/03/11 02:56 a-onl End_of_Run_2226
  • 349775 03/03/11 02:38 riordan Start of runs not logging
  • 349774 03/03/11 02:38 a-onl End_of_Run_2225
  • 349773 03/03/11 02:17 a-onl End_of_Run_2224
  • 349772 03/03/11 02:15 M Friend Target screenshots
  • 349771 03/03/11 01:57 M Friend Target to 4cm LH2
  • 349770 03/03/11 01:52 a-onl End_of_Run_2223
  • 349769 03/03/11 01:27 auto_snap Radiation and Ozone snapshot
  • 349768 03/03/11 01:26 auto_snap Miscellaneous snapshot
  • 349767 03/03/11 01:26 auto_snap Hall A Tools snapshot
  • 349766 03/03/11 01:26 auto_snap Compton Polarimeter snapshot
  • 349765 03/03/11 01:26 auto_snap BigBox Magnet snapshot
  • 349764 03/03/11 01:26 auto_snap Beam Position and Current snapshot
  • 349763 03/03/11 01:16 riordan Running on 4cm Al dummy since ~00:55
  • 349762 03/03/11 00:51 M Friend Target to 4cm Al dummy
  • 349761 03/03/11 00:47 a-onl End_of_Run_2222
  • 349760 03/03/11 00:27 a-onl End_of_Run_2221
  • 349759 03/03/11 00:16 riordan Hall A Tools
  • 349758 03/03/11 00:16 riordan spot++ Run 2221
  • 349757 03/03/11 00:12 Nikos N->Delta true events rate (Kin # 3)
  • 349756 03/03/11 00:08 a-onl Start_Run_2221,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349755 03/03/11 00:08 adaq Beam Time Accounting
  • 349754 03/03/11 00:07 a-onl End_of_Run_2220
  • Wednesday

  • 349753 03/02/11 23:58 Ole Hansen Swing Shift Summary
  • 349752 03/02/11 23:49 a-onl Start_Run_2220,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349751 03/02/11 23:48 a-onl End_of_Run_2219
  • 349750 03/02/11 23:30 whit target screens
  • 349749 03/02/11 23:29 a-onl Start_Run_2219,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349748 03/02/11 23:29 a-onl End_of_Run_2218
  • 349747 03/02/11 23:21 Ole Hansen Beam current for kine #3
  • 349746 03/02/11 23:10 a-onl Start_Run_2218,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349745 03/02/11 23:09 a-onl End_of_Run_2217
  • 349744 03/02/11 23:05 a-onl Start_Run_2217,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Short Production Run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349743 03/02/11 23:04 a-onl End_of_Run_2216
  • 349742 03/02/11 22:57 a-onl Start_Run_2216,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=25,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Setting prescales,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349741 03/02/11 22:55 a-onl End_of_Run_2215
  • 349740 03/02/11 22:54 nmuangma Trigger Dead time calculation after the end of run
  • 349739 03/02/11 22:51 a-onl End_of_Run_2214
  • 349738 03/02/11 22:48 a-onl Start_Run_2214,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Setting prescales,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349737 03/02/11 22:47 a-onl End_of_Run_2213
  • 349736 03/02/11 22:44 lselvy Readout time for ROCs 1,2,3,4
  • 349735 03/02/11 22:43 a-onl Start_Run_2213,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Setting prescales,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349734 03/02/11 22:33 a-onl Start_Run_2212,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Setting prescales,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349733 03/02/11 22:32 a-onl End_of_Run_2211
  • 349732 03/02/11 22:31 Ole Hansen [Techs] Gas alarm
  • 349731 03/02/11 22:28 a-onl Start_Run_2211,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Setting prescales,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349730 03/02/11 22:25 whit Target moved L2 4cm
  • 349729 03/02/11 22:20 a-onl End_of_Run_2210
  • 349728 03/02/11 22:14 narbe BeO Raster Check
  • 349727 03/02/11 22:12 a-onl Start_Run_2210,Run_type=Optics,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Optics run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349726 03/02/11 22:11 a-onl End_of_Run_2209
  • 349725 03/02/11 22:03 a-onl Start_Run_2209,Run_type=Production N -> Delta,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=5,leftarm_p=0.528,leftarm_theta=38.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Spot ++,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349724 03/02/11 22:01 whit target moved to BeO
  • 349723 03/02/11 22:00 Ole Hansen Recovering beam
  • 349722 03/02/11 21:49 whit moved target to empty
  • 349721 03/02/11 21:14 Ole Hansen LHRS move finished
  • 349720 03/02/11 21:14 narbe Run 2208 Cosmics
  • 349719 03/02/11 21:09 a-onl End_of_Run_2208
  • 349718 03/02/11 21:01 Higinbotham Left HRS 14 degree stop installed
  • 349717 03/02/11 20:44 narbe Run 2207 Junk
  • 349716 03/02/11 20:33 a-onl Start_Run_2208,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production; unbuffered, cosmics,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349715 03/02/11 20:31 a-onl Start_Run_2207,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production; unbuffered,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349714 03/02/11 20:31 a-onl End_of_Run_2207
  • 349713 03/02/11 20:26 Ole Hansen Moving LHRS
  • 349712 03/02/11 20:20 a-onl End_of_Run_2206
  • 349711 03/02/11 19:57 a-onl Start_Run_2206,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production; unbuffered,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349710 03/02/11 19:57 a-onl End_of_Run_2205
  • 349709 03/02/11 19:34 a-onl Start_Run_2205,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production; unbuffered,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349708 03/02/11 19:34 a-onl End_of_Run_2204
  • 349707 03/02/11 19:11 a-onl Start_Run_2204,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production; unbuffered,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349706 03/02/11 19:10 a-onl End_of_Run_2203
  • 349705 03/02/11 18:49 whit target screens
  • 349704 03/02/11 18:47 a-onl Start_Run_2203,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production; unbuffered,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349703 03/02/11 18:47 a-onl End_of_Run_2202
  • 349702 03/02/11 18:23 a-onl Start_Run_2202,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production; unbuffered,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349701 03/02/11 18:23 yez RHRS:R.cer.#1 show its s.p.e peak after increasing HVs
  • 349700 03/02/11 18:21 a-onl End_of_Run_2201
  • 349699 03/02/11 18:02 a-onl Start_Run_2201,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production; unbuffered,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349698 03/02/11 18:02 a-onl End_of_Run_2200
  • 349697 03/02/11 18:01 lselvy Deadtime code added to ROCs 1,2,3, and 4
  • 349696 03/02/11 17:40 a-onl Start_Run_2200,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production; unbuffered,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349695 03/02/11 17:39 a-onl End_of_Run_2199
  • 349694 03/02/11 17:37 Ole Hansen Charge summing script
  • 349693 03/02/11 17:35 yez increase HV of R.cer.#1 from -2050 to -2100
  • 349692 03/02/11 17:35 Ole Hansen Switched to unbuffered DAQ
  • 349691 03/02/11 17:33 a-onl Start_Run_2199,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production; unbuffered,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349690 03/02/11 17:31 a-onl End_of_Run_2198
  • 349689 03/02/11 17:26 auto_snap Radiation and Ozone snapshot
  • 349688 03/02/11 17:26 auto_snap Miscellaneous snapshot
  • 349687 03/02/11 17:26 auto_snap Hall A Tools snapshot
  • 349686 03/02/11 17:26 auto_snap Compton Polarimeter snapshot
  • 349685 03/02/11 17:26 auto_snap BigBox Magnet snapshot
  • 349684 03/02/11 17:26 auto_snap Beam Position and Current snapshot
  • 349683 03/02/11 17:19 V. Sulkosky New CODA Config unbuffered
  • 349682 03/02/11 17:10 a-onl Start_Run_2198,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production; unbuffered,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349681 03/02/11 17:04 a-onl Start_Run_2197,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production; buffered,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349680 03/02/11 17:03 a-onl End_of_Run_2197
  • 349679 03/02/11 17:02 a-onl End_of_Run_2197
  • 349678 03/02/11 17:00 A. Camsonne Macro location
  • 349677 03/02/11 16:57 jmammei run plan for overnight Wednesday
  • 349676 03/02/11 16:56 A. Camsonne Simple physics macro for signal check
  • 349675 03/02/11 16:39 a-onl Start_Run_2197,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production; buffered,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349674 03/02/11 16:36 a-onl Start_Run_2196,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production; dThresh2arm config unbuffered,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349673 03/02/11 16:28 a-onl Start_Run_2195,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production; dThresh2arm config unbuffered,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349672 03/02/11 16:21 adaq Beam Time Accounting
  • 349671 03/02/11 16:21 a-onl End_of_Run_2194
  • 349670 03/02/11 15:59 Wednesday, March 2: Day Shift Summary
  • 349669 03/02/11 15:58 a-onl Start_Run_2194,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349668 03/02/11 15:55 A. Camsonne ROC3 error message slot 25 ROC 3
  • 349667 03/02/11 15:53 a-onl Start_Run_2193,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Error message from ROC3 Fastbus Data Error,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349666 03/02/11 15:53 a-onl End_of_Run_2193
  • 349665 03/02/11 15:47 a-onl End_of_Run_2192
  • 349664 03/02/11 15:43 yez RHRS: A better plot for Cerenkov ADC distribution with TDC cuts
  • 349663 03/02/11 15:35 a-onl Start_Run_2192,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349662 03/02/11 15:33 a-onl End_of_Run_2191
  • 349661 03/02/11 15:19 flay Spot++ for Run 2191
  • 349660 03/02/11 15:12 a-onl Start_Run_2191,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349659 03/02/11 15:10 flay Beam is Coming Back Now
  • 349658 03/02/11 14:59 Hanretty Target Screens
  • 349657 03/02/11 14:59 yez RHRS: Cerenkov rate from scaler
  • 349656 03/02/11 14:56 yez RHRS: A better plot for Cerenkov ADC distribution and s.p.e peaks.
  • 349655 03/02/11 14:46 flay Beam is Down for 20 Minutes
  • 349654 03/02/11 14:45 a-onl End_of_Run_2190
  • 349653 03/02/11 14:41 flay Runs 2186--2189 are Junk
  • 349652 03/02/11 14:40 a-onl Start_Run_2190,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349651 03/02/11 14:25 Hanretty Hall A Tools
  • 349650 03/02/11 14:06 a-onl End_of_Run_2185
  • 349649 03/02/11 13:44 flay Run 2184 is Junk
  • 349648 03/02/11 13:44 a-onl Start_Run_2185,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349647 03/02/11 13:40 a-onl Start_Run_2184,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349646 03/02/11 13:39 a-onl End_of_Run_2183
  • 349645 03/02/11 13:15 a-onl Start_Run_2183,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349644 03/02/11 13:14 a-onl End_of_Run_2182
  • 349643 03/02/11 12:52 Hanretty Target Screens
  • 349642 03/02/11 12:50 a-onl Start_Run_2182,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349641 03/02/11 12:49 a-onl End_of_Run_2181
  • 349640 03/02/11 12:31 rshneor padestal suppresion for R-HRS
  • 349639 03/02/11 12:23 a-onl Start_Run_2181,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349638 03/02/11 12:20 a-onl End_of_Run_2180
  • 349637 03/02/11 12:14 R. Michaels pedestals still off, so no wonder DT big
  • 349636 03/02/11 12:07 orchen Padestal suppresion for R-HRS
  • 349635 03/02/11 12:04 R. Michaels adaql2: system check
  • 349634 03/02/11 12:03 a-onl Start_Run_2180,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349633 03/02/11 12:01 a-onl End_of_Run_2179
  • 349632 03/02/11 11:53 flay Run 2178 is Junk
  • 349631 03/02/11 11:51 a-onl Start_Run_2179,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349630 03/02/11 11:50 a-onl End_of_Run_2178
  • 349629 03/02/11 11:49 a-onl Start_Run_2178,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349628 03/02/11 11:47 a-onl End_of_Run_2177
  • 349627 03/02/11 11:44 yez RHRS: Cerenkov detectors do not have clear s.p.e. peaks,increase HVs?
  • 349626 03/02/11 11:42 Hanretty FPP HV briefly turned off to see effect on deadtime
  • 349625 03/02/11 11:38 Hanretty Target Computer reboot
  • 349624 03/02/11 11:25 a-onl Start_Run_2177,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349623 03/02/11 11:23 a-onl End_of_Run_2176
  • 349622 03/02/11 11:11 yez RHRS: shower, preshower and electron ID
  • 349621 03/02/11 11:01 a-onl Start_Run_2176,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349620 03/02/11 10:59 a-onl End_of_Run_2175
  • 349619 03/02/11 10:44 flay Run 2174 is Junk
  • 349618 03/02/11 10:37 a-onl Start_Run_2175,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349617 03/02/11 10:34 a-onl End_of_Run_2174
  • 349616 03/02/11 10:33 a-onl Start_Run_2174,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349615 03/02/11 10:32 a-onl End_of_Run_2173
  • 349614 03/02/11 10:24 jmammei RC daily summary
  • 349613 03/02/11 10:22 R. Michaels DAQ deadtime, theory, culprit: no ped suppress
  • 349612 03/02/11 10:21 a-onl Start_Run_2173,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349611 03/02/11 10:19 a-onl End_of_Run_2172
  • 349610 03/02/11 09:57 Hanretty Target screens
  • 349609 03/02/11 09:57 a-onl Start_Run_2172,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349608 03/02/11 09:56 a-onl End_of_Run_2171
  • 349607 03/02/11 09:33 a-onl Start_Run_2171,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349606 03/02/11 09:32 a-onl End_of_Run_2170
  • 349604 03/02/11 09:26 auto_snap Radiation and Ozone snapshot
  • 349603 03/02/11 09:26 auto_snap Miscellaneous snapshot
  • 349602 03/02/11 09:26 auto_snap Hall A Tools snapshot
  • 349601 03/02/11 09:26 auto_snap Compton Polarimeter snapshot
  • 349600 03/02/11 09:26 auto_snap BigBox Magnet snapshot
  • 349599 03/02/11 09:26 auto_snap Beam Position and Current snapshot
  • 349598 03/02/11 09:09 flay CODA Rebooted
  • 349597 03/02/11 09:09 a-onl Start_Run_2170,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349596 03/02/11 09:01 flay Run 2169 is Junk
  • 349595 03/02/11 08:57 a-onl Start_Run_2169,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349594 03/02/11 08:56 a-onl End_of_Run_2168
  • 349593 03/02/11 08:32 a-onl Start_Run_2168,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349592 03/02/11 08:31 a-onl End_of_Run_2167
  • 349591 03/02/11 08:04 a-onl Start_Run_2167,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349590 03/02/11 08:03 a-onl End_of_Run_2166
  • 349589 03/02/11 08:01 adaq Beam Time Accounting
  • 349588 03/02/11 08:00 riordan Owl Shift Summary
  • 349587 03/02/11 07:40 mahbub Target screens
  • 349586 03/02/11 07:40 a-onl Start_Run_2166,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349585 03/02/11 07:39 a-onl End_of_Run_2165
  • 349584 03/02/11 07:15 a-onl Start_Run_2165,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349583 03/02/11 07:15 a-onl End_of_Run_2164
  • 349582 03/02/11 06:48 a-onl Start_Run_2164,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349581 03/02/11 06:47 a-onl End_of_Run_2163
  • 349580 03/02/11 06:24 a-onl Start_Run_2163,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349579 03/02/11 06:23 a-onl End_of_Run_2162
  • 349578 03/02/11 05:59 mahbub Target screens
  • 349577 03/02/11 05:59 a-onl Start_Run_2162,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349576 03/02/11 05:58 a-onl End_of_Run_2161
  • 349575 03/02/11 05:36 a-onl Start_Run_2161,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349574 03/02/11 05:35 a-onl End_of_Run_2160
  • 349573 03/02/11 05:13 a-onl Start_Run_2160,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349572 03/02/11 05:12 a-onl End_of_Run_2159
  • 349571 03/02/11 04:49 a-onl Start_Run_2159,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349570 03/02/11 04:48 a-onl End_of_Run_2158
  • 349569 03/02/11 04:35 riordan Hall A Tools
  • 349568 03/02/11 04:26 a-onl Start_Run_2158,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349567 03/02/11 04:25 a-onl End_of_Run_2157
  • 349566 03/02/11 04:03 a-onl Start_Run_2157,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349565 03/02/11 04:02 a-onl End_of_Run_2156
  • 349564 03/02/11 03:51 mahbub Target screens
  • 349563 03/02/11 03:39 a-onl Start_Run_2156,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349562 03/02/11 03:38 a-onl End_of_Run_2155
  • 349561 03/02/11 03:17 M Friend 4cm LH2 prescales
  • 349560 03/02/11 03:15 M Friend coda reboot
  • 349559 03/02/11 03:08 a-onl Start_Run_2154,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349558 03/02/11 03:03 mahbub Target change to 4cm LH2
  • 349557 03/02/11 03:01 a-onl End_of_Run_2153
  • 349556 03/02/11 02:11 M Friend spot++ BeO, run 2152
  • 349555 03/02/11 02:06 a-onl Start_Run_2153,Run_type=Calibration,target_type=4cm dummy,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=background run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349554 03/02/11 02:05 M Friend Al prescales
  • 349553 03/02/11 02:02 mahbub Target change to 4cm Al dummy
  • 349552 03/02/11 02:00 a-onl End_of_Run_2152
  • 349551 03/02/11 01:58 mahbub Target screens
  • 349550 03/02/11 01:56 M Friend BeO prescales
  • 349549 03/02/11 01:55 a-onl Start_Run_2152,Run_type=Optics,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=pointint run,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349548 03/02/11 01:48 a-onl End_of_Run_2151
  • 349547 03/02/11 01:41 a-onl Start_Run_2151,Run_type=Spot++,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=0.5,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=spot++, 12x8 raster, ps3=1,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349546 03/02/11 01:41 M Friend spot++ run 2150
  • 349545 03/02/11 01:40 a-onl End_of_Run_2150
  • 349544 03/02/11 01:37 a-onl Start_Run_2150,Run_type=Spot++,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=0.5,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=spot++, 12x8 raster,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349543 03/02/11 01:36 M Friend spot++ run 2149
  • 349542 03/02/11 01:35 a-onl End_of_Run_2149
  • 349541 03/02/11 01:32 a-onl Start_Run_2149,Run_type=Spot++,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=0.5,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=spot++,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349540 03/02/11 01:31 a-onl Start_Run_2148,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349538 03/02/11 01:26 auto_snap Radiation and Ozone snapshot
  • 349537 03/02/11 01:26 auto_snap Miscellaneous snapshot
  • 349536 03/02/11 01:25 auto_snap Hall A Tools snapshot
  • 349535 03/02/11 01:25 auto_snap Compton Polarimeter snapshot
  • 349534 03/02/11 01:25 auto_snap BigBox Magnet snapshot
  • 349533 03/02/11 01:25 auto_snap Beam Position and Current snapshot
  • 349532 03/02/11 01:24 mahbub Target change to BeO
  • 349531 03/02/11 01:22 a-onl End_of_Run_2147
  • 349530 03/02/11 01:01 M Friend Extra peak in TDC spectrum?
  • 349529 03/02/11 00:59 a-onl Start_Run_2147,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349528 03/02/11 00:58 a-onl End_of_Run_2146
  • 349527 03/02/11 00:48 riordan Increased LHRS VDC HV by ~1%
  • 349526 03/02/11 00:36 a-onl Start_Run_2146,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349525 03/02/11 00:35 a-onl End_of_Run_2145
  • 349524 03/02/11 00:27 V. Sulkosky CT Peak 2144
  • 349523 03/02/11 00:23 A. Camsonne Retiming ok on left arm
  • 349522 03/02/11 00:17 A. Camsonne Added sync plots for ROC sync clocks
  • 349521 03/02/11 00:11 a-onl Start_Run_2145,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349520 03/02/11 00:10 a-onl End_of_Run_2144
  • 349519 03/02/11 00:07 mahbub Target screens
  • 349518 03/02/11 00:04 kalyan Re: CT spectra from this evening
  • 349517 03/02/11 00:01 adaq Beam Time Accounting
  • 349516 03/02/11 00:01 efuchey 3/01/11 swing shift summary
  • 349515 03/02/11 00:01 A. Camsonne Test timer library
  • Tuesday

  • 349514 03/01/11 23:58 a-onl Start_Run_2144,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349513 03/01/11 23:55 efuchey beam back
  • 349512 03/01/11 23:53 V. Sulkosky Reduced width on T3 Window to 70 ns
  • 349511 03/01/11 23:52 V. Sulkosky Re: DBB BigBite TDC Info Fixed
  • 349510 03/01/11 23:45 V. Sulkosky CT spectra from this evening
  • 349509 03/01/11 23:42 efuchey Going back to beam permit.
  • 349508 03/01/11 23:22 whit target screens
  • 349507 03/01/11 23:22 a-onl End_of_Run_2143
  • 349506 03/01/11 23:21 efuchey Switching to escorted access
  • 349505 03/01/11 23:17 V. Sulkosky DBB BigBite TDC Info Issue
  • 349504 03/01/11 23:14 jmammei second prescale change was to 3
  • 349503 03/01/11 23:12 a-onl Start_Run_2143,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349502 03/01/11 23:08 a-onl End_of_Run_2142
  • 349501 03/01/11 23:03 lselvy High Current High Prescale
  • 349500 03/01/11 22:58 a-onl Start_Run_2142,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349499 03/01/11 22:57 a-onl End_of_Run_2141
  • 349498 03/01/11 22:56 a-onl Start_Run_2141,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349497 03/01/11 22:47 a-onl End_of_Run_2140
  • 349496 03/01/11 22:45 efuchey Hall A tools screenshot
  • 349495 03/01/11 22:40 a-onl Start_Run_2140,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349494 03/01/11 22:35 a-onl Start_Run_2139,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349493 03/01/11 22:29 a-onl Start_Run_2138,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=junk,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349492 03/01/11 22:14 a-onl Start_Run_2136,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=junk,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349491 03/01/11 22:13 a-onl Start_Run_2135,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=junk,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349490 03/01/11 22:11 a-onl End_of_Run_2134
  • 349489 03/01/11 21:55 a-onl Start_Run_2134,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=junk,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349488 03/01/11 21:53 a-onl Start_Run_2133,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=junk,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349487 03/01/11 21:52 a-onl End_of_Run_2132
  • 349486 03/01/11 21:50 efuchey Beam back
  • 349485 03/01/11 21:49 a-onl Start_Run_2132,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DT test, full production, BBTS1 out of readout,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349484 03/01/11 21:29 V. Sulkosky Changed width on T3 Window
  • 349483 03/01/11 21:28 efuchey Got back to beam permit
  • 349482 03/01/11 21:08 efuchey Switching to controlled access
  • 349481 03/01/11 21:08 a-onl End_of_Run_2131
  • 349480 03/01/11 20:50 a-onl Start_Run_2131,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DT test, full production, BBTS1 out of readout,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349479 03/01/11 20:48 a-onl End_of_Run_2130
  • 349478 03/01/11 20:37 lselvy Trigger timing
  • 349477 03/01/11 20:37 a-onl Start_Run_2130,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DT test, full production, BBTS1 out of readout,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349476 03/01/11 20:35 a-onl End_of_Run_2129
  • 349475 03/01/11 20:19 a-onl Start_Run_2129,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DT test, full production, BBTS1 out of readout,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349474 03/01/11 20:17 a-onl End_of_Run_2128
  • 349473 03/01/11 20:17 lselvy DT Tests
  • 349472 03/01/11 20:13 a-onl Start_Run_2128,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DT test, full production, BBTS1 out of readout,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349471 03/01/11 20:08 a-onl End_of_Run_2127
  • 349470 03/01/11 19:58 a-onl Start_Run_2127,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DT test, full production, BBTS1 out of readout,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349469 03/01/11 19:36 a-onl Start_Run_2125,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DT test, both arms coincidence at 500 Hz,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349468 03/01/11 19:31 a-onl Start_Run_2124,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DT test, both arms coincidence at 500 Hz,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349467 03/01/11 19:25 a-onl Start_Run_2123,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DT test, both arms at 500 Hz,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349466 03/01/11 19:21 a-onl Start_Run_2122,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DT test, left arm only in full 2 arm configuration,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349465 03/01/11 19:19 a-onl Start_Run_2121,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DT test, left arm only,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349464 03/01/11 19:19 R. Michaels comments re: deadtime
  • 349463 03/01/11 19:17 a-onl End_of_Run_2120
  • 349462 03/01/11 19:16 a-onl Start_Run_2120,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DT test, right arm only ,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349461 03/01/11 19:13 a-onl End_of_Run_2119
  • 349460 03/01/11 18:48 a-onl Start_Run_2119,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349459 03/01/11 18:47 a-onl End_of_Run_2118
  • 349458 03/01/11 18:23 a-onl Start_Run_2118,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349457 03/01/11 18:22 a-onl End_of_Run_2117
  • 349452 03/01/11 17:58 a-onl Start_Run_2117,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349451 03/01/11 17:55 a-onl End_of_Run_2116
  • 349445 03/01/11 17:25 auto_snap Radiation and Ozone snapshot
  • 349444 03/01/11 17:25 auto_snap Miscellaneous snapshot
  • 349443 03/01/11 17:25 auto_snap Hall A Tools snapshot
  • 349442 03/01/11 17:25 auto_snap Compton Polarimeter snapshot
  • 349441 03/01/11 17:25 auto_snap BigBox Magnet snapshot
  • 349440 03/01/11 17:25 auto_snap Beam Position and Current snapshot
  • 349439 03/01/11 17:23 a-onl Start_Run_2116,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349438 03/01/11 17:22 a-onl End_of_Run_2115
  • 349437 03/01/11 17:20 a-onl Start_Run_2115,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.528,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=out,hand_lead=on
  • 349436 03/01/11 17:13 a-onl End_of_Run_2114
  • 349435 03/01/11 17:11 lselvy Deadtime Update
  • 349434 03/01/11 17:09 a-onl Start_Run_2114,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DAQ DT Testing,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349433 03/01/11 17:08 a-onl End_of_Run_2113
  • 349432 03/01/11 17:07 a-onl Start_Run_2113,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DAQ DT Testing,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349431 03/01/11 17:06 a-onl End_of_Run_2112
  • 349430 03/01/11 17:05 a-onl Start_Run_2112,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DAQ DT Testing,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349429 03/01/11 17:03 a-onl End_of_Run_2111
  • 349428 03/01/11 17:03 a-onl Start_Run_2111,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DAQ DT Testing,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349427 03/01/11 17:01 a-onl End_of_Run_2110
  • 349426 03/01/11 17:00 a-onl Start_Run_2110,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DAQ DT Testing,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349425 03/01/11 16:57 a-onl End_of_Run_2109
  • 349424 03/01/11 16:56 a-onl Start_Run_2109,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DAQ DT Testing,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349423 03/01/11 16:55 a-onl End_of_Run_2108
  • 349422 03/01/11 16:54 whit target screens
  • 349421 03/01/11 16:53 a-onl Start_Run_2108,Run_type=Other,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=12.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=DAQ DT Testing,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349420 03/01/11 16:42 efuchey Beam recovery procedure
  • 349419 03/01/11 16:40 a-onl End_of_Run_2106
  • 349418 03/01/11 16:39 efuchey Spot ++
  • 349417 03/01/11 16:26 jmammei settings for Al and Be0 runs
  • 349416 03/01/11 16:22 Whit target
  • 349415 03/01/11 16:16 efuchey Harp scan
  • 349414 03/01/11 16:02 Saha Shift Summary
  • 349413 03/01/11 16:01 adaq Beam Time Accounting
  • 349412 03/01/11 15:56 jmammei Run Plan for overnight Tuesday
  • 349411 03/01/11 15:42 X. Zheng Target changed haCryo_Supply_F alarm limits
  • 349410 03/01/11 15:35 lselvy Reverted CRL code for ROC 1 and 2 back
  • 349408 03/01/11 15:18 lselvy New Configuration Changes
  • 349407 03/01/11 15:16 a-onl End_of_Run_2105
  • 349406 03/01/11 15:13 a-onl Start_Run_2105,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=comsics,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349404 03/01/11 15:08 a-onl End_of_Run_2104
  • 349403 03/01/11 15:07 lselvy New Configuration
  • 349402 03/01/11 15:01 jmammei MCC requested more time for beam studies
  • 349401 03/01/11 15:01 X. Zheng Target fans back to 48Hz (20%)
  • 349400 03/01/11 14:50 X. Zheng Target temperature back to normal
  • 349399 03/01/11 14:37 V. Sulkosky Changed T1 delay to 32 ns
  • 349398 03/01/11 14:35 V. Sulkosky Changed widths on T1 and T3
  • 349397 03/01/11 14:28 a-onl Start_Run_2104,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=comsics ,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349396 03/01/11 13:57 a-onl End_of_Run_2102
  • 349395 03/01/11 13:54 a-onl End_of_Run_2102
  • 349394 03/01/11 13:52 a-onl Start_Run_2102,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=comsics DAQ tests,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349393 03/01/11 13:51 a-onl End_of_Run_2101
  • 349392 03/01/11 13:50 a-onl Start_Run_2101,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=comsics DAQ tests,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349391 03/01/11 13:47 a-onl End_of_Run_2100
  • 349390 03/01/11 13:44 a-onl Start_Run_2100,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=0,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=comsics DAQ tests ,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349389 03/01/11 13:42 a-onl End_of_Run_2099
  • 349388 03/01/11 13:36 lselvy Major revision to crl code for Rocs 1 and 2
  • 349387 03/01/11 13:29 a-onl Start_Run_2099,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=comsics DAQ tests Saw scalers disabled,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349386 03/01/11 13:24 lselvy ROC 1 adn 2 CRL code
  • 349385 03/01/11 13:18 a-onl End_of_Run_2098
  • 349384 03/01/11 13:17 a-onl Start_Run_2098,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=comsics DAQ tests Saw scalers disabled,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349383 03/01/11 13:16 a-onl End_of_Run_2097
  • 349382 03/01/11 13:15 lselvy Pedestal issues with ROC 1 and 2
  • 349381 03/01/11 13:14 a-onl Start_Run_2097,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=comsics DAQ tests,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349380 03/01/11 13:10 V. Sulkosky Buffered vs. Unbuffered
  • 349379 03/01/11 13:02 a-onl End_of_Run_2095
  • 349378 03/01/11 12:57 a-onl End_of_Run_2095
  • 349377 03/01/11 12:55 a-onl Start_Run_2095,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=comsics DAQ tests,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349376 03/01/11 12:53 a-onl End_of_Run_2094
  • 349375 03/01/11 12:52 X. Zheng cryo back to normal
  • 349374 03/01/11 12:51 a-onl Start_Run_2094,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=comsics DAQ tests,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349373 03/01/11 12:48 a-onl End_of_Run_2093
  • 349372 03/01/11 12:45 a-onl Start_Run_2093,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=comsics DAQ tests,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349371 03/01/11 12:44 a-onl End_of_Run_2092
  • 349370 03/01/11 12:41 a-onl Start_Run_2092,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=comsics DAQ tests,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349369 03/01/11 12:39 a-onl End_of_Run_2091
  • 349368 03/01/11 12:38 X. Zheng Lost both HRS magnets
  • 349367 03/01/11 12:36 a-onl Start_Run_2091,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=comsics DAQ tests,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349366 03/01/11 12:35 lselvy One further change to the crl for ROC 1 and 2
  • 349365 03/01/11 12:34 X. Zheng Target is warming up due to ESR problems
  • 349364 03/01/11 12:27 a-onl End_of_Run_2088
  • 349363 03/01/11 12:22 a-onl End_of_Run_2088
  • 349362 03/01/11 12:22 lselvy Another change to the CRL for ROCs 1 and 2
  • 349361 03/01/11 12:09 a-onl End_of_Run_2088
  • 349360 03/01/11 12:08 a-onl Start_Run_2088,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=comsics DAQ tests,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349359 03/01/11 12:04 a-onl End_of_Run_2086
  • 349358 03/01/11 12:02 Higinbotham Left HRS to 12.511 degrees
  • 349357 03/01/11 12:00 V. Sulkosky Busy Signals Gathered at weldment
  • 349356 03/01/11 11:59 lselvy Change in ROC 1 and 2 crl
  • 349355 03/01/11 11:56 a-onl End_of_Run_2086
  • 349354 03/01/11 11:49 X. Zheng Left HRS momentum changed to 0.528 GeV/c
  • 349353 03/01/11 11:13 A. Camsonne Left arm FE BUSY sent to cable spare/busy
  • 349352 03/01/11 10:59 a-onl Start_Run_2086,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=comsics DAQ tests,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349351 03/01/11 10:53 A. Camsonne Right arm fastbus busy sent to cable brown
  • 349350 03/01/11 10:52 a-onl End_of_Run_2085
  • 349349 03/01/11 10:45 A. Camsonne Uncommented scaler initialization and clear in go for sync scaler in ROC2
  • 349348 03/01/11 10:45 a-onl Start_Run_2085,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=comsics DAQ tests,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349347 03/01/11 10:43 a-onl End_of_Run_2084
  • 349346 03/01/11 10:36 lselvy Dec data class temporary change
  • 349345 03/01/11 10:34 a-onl Start_Run_2084,Run_type=Other,target_type=No Target, Empty,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=comsics DAQ tests,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349344 03/01/11 10:03 lselvy Moving software to Buffered mode
  • 349343 03/01/11 09:25 auto_snap Radiation and Ozone snapshot
  • 349342 03/01/11 09:25 auto_snap Miscellaneous snapshot
  • 349341 03/01/11 09:25 auto_snap Hall A Tools snapshot
  • 349340 03/01/11 09:25 auto_snap Compton Polarimeter snapshot
  • 349339 03/01/11 09:25 auto_snap BigBox Magnet snapshot
  • 349338 03/01/11 09:25 auto_snap Beam Position and Current snapshot
  • 349337 03/01/11 09:01 a-onl End_of_Run_2083
  • 349336 03/01/11 08:33 a-onl Start_Run_2083,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349335 03/01/11 08:32 a-onl End_of_Run_2082
  • 349334 03/01/11 08:12 adaq Beam Time Accounting
  • 349333 03/01/11 08:04 danez Owl Shift Summary
  • 349332 03/01/11 08:01 a-onl Start_Run_2082,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349331 03/01/11 08:00 a-onl End_of_Run_2081
  • 349330 03/01/11 07:30 a-onl Start_Run_2081,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349329 03/01/11 07:29 a-onl End_of_Run_2080
  • 349328 03/01/11 07:11 danez Tools Screenshot
  • 349327 03/01/11 07:08 danez Spot++ Screenshots
  • 349326 03/01/11 07:01 a-onl Start_Run_2080,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349325 03/01/11 06:59 a-onl End_of_Run_2079
  • 349324 03/01/11 06:49 danez Target Screenshots
  • 349323 03/01/11 06:30 a-onl Start_Run_2079,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349322 03/01/11 06:29 a-onl End_of_Run_2078
  • 349321 03/01/11 06:02 danez Moved to 4cm LH2
  • 349320 03/01/11 05:58 a-onl Start_Run_2078,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349319 03/01/11 05:49 a-onl End_of_Run_2077
  • 349318 03/01/11 05:44 a-onl Start_Run_2077,Run_type=Production N -> Delta,target_type=BeO,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349317 03/01/11 05:39 a-onl End_of_Run_2076
  • 349316 03/01/11 04:38 a-onl Start_Run_2076,Run_type=Production N -> Delta,target_type=4cm dummy,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349315 03/01/11 04:30 danez Moved Target to 4cm Al dummy
  • 349314 03/01/11 04:29 a-onl End_of_Run_2075
  • 349313 03/01/11 04:26 danez Target Screenshots
  • 349312 03/01/11 04:03 a-onl Start_Run_2075,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349311 03/01/11 04:02 a-onl End_of_Run_2074
  • 349310 03/01/11 03:33 a-onl Start_Run_2074,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349309 03/01/11 03:32 a-onl End_of_Run_2073
  • 349308 03/01/11 03:00 a-onl Start_Run_2073,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349307 03/01/11 03:00 a-onl End_of_Run_2072
  • 349306 03/01/11 02:30 a-onl Start_Run_2072,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349305 03/01/11 02:29 a-onl End_of_Run_2071
  • 349304 03/01/11 02:15 danez Target Screenshots
  • 349303 03/01/11 02:08 a-onl Start_Run_2071,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349302 03/01/11 02:06 a-onl End_of_Run_2070
  • 349301 03/01/11 02:06 a-onl Start_Run_2070,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349300 03/01/11 02:05 rshneor coicidence time at 15 mA and 30 mA
  • 349299 03/01/11 01:55 a-onl End_of_Run_2069
  • 349298 03/01/11 01:44 yez RHRS: Multi-peaks issue resolved after replacing the logic unit
  • 349297 03/01/11 01:37 a-onl Start_Run_2069,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=30,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349296 03/01/11 01:35 a-onl End_of_Run_2068
  • 349295 03/01/11 01:33 a-onl Start_Run_2068,Run_type=Production N -> Delta,target_type=LH2 4cm cell,beam_e0=1.16,beam_raster=on,beam_current=15,leftarm_p=0.552,leftarm_theta=25.5,leftarm_sieve=out,rightarm_p=0.805,rightarm_theta=12.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=97,hand_angle=97,bbite_sieve=sieve,hand_lead=on
  • 349294 03/01/11 01:25 auto_snap Radiation and Ozone snapshot
  • 349293 03/01/11 01:25 auto_snap Miscellaneous snapshot
  • 349292 03/01/11 01:24 auto_snap Hall A Tools snapshot
  • 349291 03/01/11 01:24 auto_snap Compton Polarimeter snapshot
  • 349290 03/01/11 01:24 auto_snap BigBox Magnet snapshot
  • 349289 03/01/11 01:24 auto_snap Beam Position and Current snapshot
  • 349288 03/01/11 01:14 yez RHRS: Replaced a 758 Logic unit for S2m odd bars and put S2m back in the T1 trigger
  • 349287 03/01/11 00:18 a-onl End_of_Run_2067
  • 349286 03/01/11 00:16 danez Tools Screenshot
  • 349285 03/01/11 00:14 danez Target Screenshots
  • 349284 03/01/11 00:07 adaq Beam Time Accounting
  • 349283 03/01/11 00:07 norum Shift summary

  • Earlier months at hallaweb.jlab.org/adaq (archive).  The HALOG was generated by version HA1.48
    If you have any comments or problems, please contact:
    Robert Michaels
    rom@jlab.org
    (757) 269-7410